### v1

# TimeBench: A Comprehensive Evaluation of Temporal Reasoning Abilities in Large Language Models

Abstract
Understanding time is a pivotal aspect of human cognition, crucial in the broader framework of grasping the intricacies of the world. Previous studies typically focus on specific aspects of time, lacking a comprehensive temporal reasoning benchmark. To address this issue, we propose TimeBench, a comprehensive hierarchical temporal reasoning benchmark that covers a broad spectrum of temporal reasoning phenomena, which provides a thorough evaluation for investigating the temporal reasoning capabilities of large language models. We conduct extensive experiments on popular LLMs, such as GPT-4, LLaMA2, and Mistral, incorporating chain-of-thought prompting. Our experimental results indicate a significant performance gap between the state-of-the-art LLMs and humans, highlighting that there is still a considerable distance to cover in temporal reasoning. We aspire for TimeBench to serve as a comprehensive benchmark, fostering research in temporal reasoning for LLMs. Our resource is available at https://github.com/zchuz/TimeBench
1 Introduction
Time flies over us, but leaves its shadow behind. Time continually moves forward, threading through the fabric of people’s lives. Understanding time is a crucial part of human comprehension of the world. Imagine witnessing the blossoming of flowers, and you’ll connect it to the arrival of spring. The ponder within it encompasses the intricate interplay of world knowledge, causality, and event temporal relationships. Temporal reasoning, in contrast to reasoning of a singular nature, comes with inherent complexity, encompassing implicit arithmetic, logical implications, and world knowledge. It is a form of integrated reasoning built upon foundational reasoning like mathematical and logical reasoning (Cobbe et al., 2021; Mishra et al., 2022; Yu et al., 2020). Recently, large language models (LLM) have demonstrated remarkable performance in complex reasoning (Hendrycks et al., 2021; Srivastava et al., 2022; Brown et al., 2020; Chowdhery et al., 2023; OpenAI, 2023; Touvron et al., 2023), but their performance in intricate temporal reasoning still lacks a definitive conclusion. The research community currently lacks a comprehensive benchmark, which makes it challenging to quantify the temporal reasoning capabilities of LLMs.
Figure 1: A brief overview of human and LLMs’ performance on TimeBench. Human scores are annotated.
Recent benchmarks for temporal reasoning typically focus only on a few aspects, such as temporal commonsense or temporal question answering (Zhou et al., 2019; Chen et al., 2021; Dhingra et al., 2022; Wang and Zhao, 2023). Due to the inherent complexity of temporal reasoning, it is challenging to accurately measure models’ temporal reasoning capabilities based on limited aspects.
To address this issue, we propose TimeBench, a comprehensive and hierarchical temporal reasoning benchmark, which is aligned with intrinsic real-world scenarios. Specifically, drawing inspiration from the human cognitive process of transitioning from abstraction and concreteness to integration (Barsalou et al., 2018), we categorize temporal reasoning into three levels: symbolic temporal reasoning, commonsense temporal reasoning, and event temporal reasoning. These levels respectively represent understanding abstract time expression, grasping concrete world knowledge, and integrating and applying this knowledge in real-world scenarios. It comprises 10 tasks with 16 sub-tasks, covering a broad spectrum of temporal reasoning phenomena. Besides, prior work typically featured only a singular task form, overly simplistic and insufficient in capturing the model’s performance in real-world scenarios. In contrast to them, we incorporate four distinct task forms, offering a more realistic simulation of challenges models encounter in real-world scenarios.
To quantify the temporal reasoning capabilities of contemporary LLMs, we extensively assess widely-used LLMs, including closed-source models such as ChatGPT (Ouyang et al., 2022) and GPT-4 (OpenAI, 2023), as well as open-source like LLaMA2 (Touvron et al., 2023), Vicuna-1.5 (Chiang et al., 2023), Mistral (Jiang et al., 2023), Baichuan2 (Yang et al., 2023), ChatGLM3 Zeng et al. (2023) and FLAN-T5 (Shen et al., 2023). We conduct experiments under zero-shot and few-shot settings, combining commonly used reasoning techniques, chain-of-thought prompting (Kojima et al., 2022; Wei et al., 2022; Chu et al., 2023a). The experimental results suggest that GPT-4 continues to outperform other models, showcasing strong temporal reasoning capabilities, as shown in Figure 1. Nevertheless, there is still a considerable gap with humans. On the contrary, open-source models generally exhibit subpar performance, lagging significantly behind GPT-3.5 and showing notable shortcomings in commonsense reasoning. These findings indicate that there is still significant room for improvement in models’ temporal reasoning capabilities. In addition, we also observe that, unlike tasks such as mathematical or logical reasoning, chain-of-thought prompting does not consistently enhance model performance; in some cases, it might even impair performance.
We summarize our contributions as follows: (1) We introduce TimeBench, a comprehensive and hierarchical temporal reasoning benchmark. (2) We conduct extensive experiments, revealing a significant gap between even sota LLM and humans, indicating substantial research opportunities in this field. (3) By conducting a thorough analysis, we uncover the dilemmas that models face in temporal reasoning and identify potential solutions.
Figure 2: Distribution of category, task and subtask in TimeBench. Degree of arc indicates the ratio of datas.
2 TimeBench Benchmark
2.1 Benchmark Design Principal
TimeBench focuses on a comprehensive evaluation of the temporal reasoning capabilities of large language models in challenging and complex scenarios. To achieve this goal, we summarize the difficulties and challenges faced in temporal reasoning, categorize it into three levels, and meticulously design tasks that better align with complex real-world scenarios.
Just as the human cognitive process unfolds from foundational cognition and conceptual understanding to practical reasoning, we delineate temporal reasoning into three hierarchical levels. Specifically, TimeBench categorizes temporal reasoning into symbolic, commonsense and event temporal reasoning, covering 10 datasets with a total of 16 subtasks. (1) Symbolic Temporal Reasoning focuses on comprehension of fundamental abstract temporal expressions. (2) Temporal Commonsense Reasoning emphasizes the mastery of temporal principles, concepts and world knowledge. (3) Event Temporal Reasoning concentrates on modeling the temporal relationships of events within authentic scenarios. Furthermore, to better align with real-world scenarios, we employ diverse question formats.
2.2 Difficulties and Challenges
We outline the necessary capabilities and challenges encountered from a human cognition perspective during temporal reasoning, and language models confront similar obstacles. Detailed information on TimeBench and challenges involved in each task is presented in Table 6.
TimeX Understanding
Time expressions (TimeX) denote words or phrases that convey information about time and represent the simplest and most basic units of expressing time, such as in August 2008, which is fundamental in comprehending time.
Temporal Commonsense
assesses the understanding of temporal aspects of world knowledge, including event order, event duration, event typical time, and event frequency, which is crucial for language models to comprehend everyday scenarios.
Event-Time Relations
evaluates the grounding ability to establish correspondence between events and time, enabling the model to comprehend the development and changes of events as they unfold over time.
Event-Event Relations
not only involve event-time grounding but also introduce multi-hop relative connections. Models with this capability can better handle temporal reasoning in complex scenarios involving multiple events.
Implicit Temporal Reasoning
involves going beyond the surface of texts, engaging in deeper reasoning such as drawing upon temporal commonsense, identifying implicit temporal factors and discerning hidden temporal relationships among events. Implicit temporal reasoning is pivotal in complex real-world scenarios where events and time are intricately interwoven.
Date Arith
Q: What is the time 2 year and 4 month before Mar, 1755
A: Nov, 1752
TimeX NLI
Premise: On 28th May 1967, I graduated.
Hypothesis: Before 23rd October 1920, I graduated.
A: Contradiction
Table 1: Examples of symbolic temporal reasoning
2.3 Symbolic Temporal Reasoning
To evaluate the language model’s comprehension of abstract time expressions, we utilize two symbolic reasoning tasks stripped of semantic content: TimeX arithmetic and TimeX inference. Table 1 shows examples of symbolic temporal reasoning.
TimeX Arithmetic (Tan et al., 2023)
assesses the model’s grasp of abstract date calculation. When provided with a date, the model needs to accurately calculate the date a certain amount of time before or after the given date.
TimeX NLI (Thukral et al., 2021)
focuses on the logical entailment relationships among abstract TimeX, including three aspects: order (s1), duration (s2), and duration with unit conversion (s3).
MCTACO
C: Ransome looks after her as well as for young Fern Simon , who has declared her love for him.
Q: How often do Ransome and Fern talk?
O: each century, once a day, once a century, every night
TimeDial
Dialog: … Person1: Do you go to work by train every day  Person2: Yes . I commute <MASK> a week by train…
O: five days, 25 days, a minute, six days
SituatedGen
Keywords: axis, one day, one month, Earth, Moon
A: Earth rotates on its axis once in one day. It takes one month for the Moon to rotate on its axis.
Table 2: Examples of commonsense temporal reasoning.
2.4 Commonsense Temporal Reasoning
We measure the model’s mastery of temporal common and world knowledge, along with its capacity for reasoning based on these insights. Table 2 presents examples of temporal commonsense reasoning in QA and generation forms.
MCTACO (Zhou et al., 2019)
evaluates diverse commonsense knowledge from different aspects of events, including duration, frequency, order, stationary and typical event time.
DurationQA (Virgo et al., 2022)
focuses specifically on temporal commonsense reasoning in the spectrum of event duration.
TimeDial (Qin et al., 2021)
considers temporal commonsense reasoning in dialogue scenarios and involves various aspects of commonsense associated with duration, order, and world knowledge.
SituatedGen (Zhang and Wan, 2023)
considers generative commonsense reasoning in a constrained text generation scenario. Given a set of contrasting keywords, the model needs to choose appropriate keywords for each sentence and generate a pair of contrasting sentences that satisfy temporal commonsense.
TimeQA
C: … He worked in Utrecht for the firm of P Smits & de Wolf from 1864 to 1867 and then returned to Widnes …
Q: Where did Ludwig Mond work between Mar 1866 and Sep 1866?
A: Utrecht
MenatQA
C: … After the French evacuated Egypt in 1801, Hurshid Pasha was named governor of Egypt in 1804. Muhammad Ali had himself named governor of Egypt in May 1805 …
Q: Which position did Hurshid Pasha hold from 1804 to 1806, if Hurshid Pasha tepped down as the governor of Egypt in 1808?
A: governor of Egypt
TempReason
C: … Peter Corke works for Queensland University of Technology from Jan, 2010 to Dec, 2022. Peter Corke works for Commonwealth Scientific from Jan, 1984 to Jan, 2009. …
Q: Which employer did Peter Corke work for before Queensland University of Technology?
A: Commonwealth Scientific
Table 3: Examples of event temporal reasoning. C=context, Q=question, A=answer
2.5 Event Temporal Reasoning
Event temporal reasoning assesses the model’s understanding of relationships between events and time in real-world scenarios, as well as its ability to reasoning under certain temporal or event constraints. Examples are shown in Table 3.
TimeQA (Chen et al., 2021)
requires the model to answer time-sensitive questions based on context containing numerous time-involved facts. It is categorized into explicit reasoning and implicit reasoning based on time indicators (before, in, etc.).
MenatQA (Wei et al., 2023)
introduces temporal factors to elicit implicit temporal reasoning, including time scope change, disruption of facts, and counterfactual questions, which provides a more in-depth assessment of implicit reasoning ability on event-time relations.
TempReason (Tan et al., 2023)
removes irrelevant context and focuses on implicit temporal reasoning within structured facts, investigating the model’s capability boundaries. It involves event-time reasoning and event-event reasoning.
TRACIE Zhou et al. (2021)
evaluates the model’s comprehension of temporal order between implicit events. The model needs to identify events implied in the context and then determine their chronological order.
2.6 Task Format and Evaluation Metrics
Real-world scenarios are diverse in task forms. To investigate the model’s authentic performance, we select diverse task forms, covering a wide spectrum of temporal reasoning phenomena.
Multi-choice & Multi-answer Questions
MCQ tasks require models to select the only correct answer from the options. However, this task form has shortcuts and may not truly reflect the model’s capability. To address this, we use the MCMAQ form, where the model needs to select all possible correct answers from options. In our task, each question includes four options, with two of them being correct. Task: MCTACO, DurationQA, TimeDial.
Natural Language Inference
is the task of determining the logical relationship between two pieces of text. Specifically, given a premise and a hypothesis, the model needs to determine whether the hypothesis can be inferred from the premise and output entailment, contradiction, or neutral. Our tasks focus on the entailment in temporal domains. Tasks: TimeX-NLI, TRACIE.
Free-form Reading Comprehension
requires models to answer questions based on the provided context, and the answer is free-form without pre-defined restrictions. Tasks: TimeQA, MenatQA, TempReason, Date Calculation.
Constrained Text Generation
refers to the task of generating text under certain constraints. Our task is keyword-constrained text generation, where the model takes keywords as input and outputs sentences that include those keywords. Task: SituatedGen.
Metrics
We adopt the evaluation metrics from previous work. Accuracy is used for NLI and date calculation tasks. MAMCQ tasks are measured using option-based EM and F1. FRC tasks (w/o date calc.) are evaluated with token-based EM and F1. For CTG tasks, we use a composite generation metrics, with more details in Appendix A.2.
3 Methodology
We perform evaluations using the prompt-based approach, including standard I-O prompting and chain-of-thought prompting. Experiments are conducted under both zero-shot and few-shot settings.
Standard I-O Prompting
We formulate specific instructions for each task. In the zero-shot setting, models follow the instructions to answer questions. In the few-shot setting, models are provided with several question-answer pairs as demonstrations and emulate those instances to answer questions.
$\mathrm{prompt_{zs}^{io}}=\{\mathrm{INST}\}\{\mathrm{Q}\}$ (1)
$\mathrm{prompt_{fs}^{io}}=\{\mathrm{INST}\}\{\mathrm{Q_{1}}\}\{\mathrm{A_{1}}\}..\{\mathrm{Q}\}$ (2)
Chain-of-Thought Prompting
We use identical instructions to standard input-output prompting. In the zero-shot chain-of-thought scenario, following Zeroshot CoT (Kojima et al., 2022), we add a reasoning trigger Let’s think step by step after questions to perform chain-of-thought reasoning. In the few-shot chain-of-thought scenario, we manually craft rationales for each task as demonstrations to guide the model to reason step-by-step. The instructions and demonstrations can be found in the Appendix A.4.
$\mathrm{prompt_{zs}^{cot}}=\{\mathrm{INST}\}\{{\mathrm{Q}}\}\{{\mathrm{TRIG}}\}$ (3)
$\mathrm{prompt_{fs}^{cot}}=\{\mathrm{INST}\}\{{\mathrm{Q_{1}}}\}\{{\mathrm{R_{1}}}\}\{{\mathrm{A_{1}}}\}..\{{\mathrm{Q}}\}$ (4)
4 Experimental Setup
4.1 Models
We evaluate several popular LLMs, including both open-source and closed-source models, with parameter sizes ranging from 6B to 175B [^1] Since OPENAI does not disclose ChatGPT-3.5/4’s size publicly, we use the size of GPT-3, which is 175 billion.. Unless otherwise stated, we use aligned models in the following experiments. By default, we use 13B model for LLaMA2/Vicuna-1.5/Baichuan2 and 11B model for FLAN-T5.
ChatGPT-3.5/GPT-4 (Ouyang et al., 2022; OpenAI, 2023)
ChatGPT is a chat model aligned through SFT and RLHF based on GPT-3 (Brown et al., 2020). GPT-4 is an upgraded version of ChatGPT with enhanced reasoning capabilities, making it the most powerful LLM. Unless otherwise stated, ChatGPT refers to gpt-3.5-turbo-0613 and GPT-4 refers to gpt-4-0613.
Llama2/Vicuna-1.5 (Touvron et al., 2023; Chiang et al., 2023)
LLaMA2 is an open foundation model trained on 2T tokens with efficient grouped-query attention Ainslie et al. (2023). LLaMA2-chat is the official aligned model with SFT and RLHF, and Vicuna-1.5 is an unofficial aligned model with SFT only.
Baichuan2 (Yang et al., 2023)
is an open foundation model pre-trained on 2.6T tokens and outperforms LLaMA2 on several benchmarks. Baichuan2-chat is the official aligned model with SFT and RLHF.
Mistral (Jiang et al., 2023)
is a 7B open foundation model incorporating efficient grouped-query attention (Ainslie et al., 2023) and sliding windows attention (Beltagy et al., 2020). It achieves the strongest performance among models of its size, even surpassing LLaMA2-13B. Mistral-instruct is the officially aligned model with SFT only.
ChatGLM3 (Zeng et al., 2023)
is an open-source bilingual model for Chinese and English, exhibiting competitive performance in models with sizes under 10B.
FLAN-T5 (Shen et al., 2023)
is an open-source instruction-following model built on top of T5 (Raffel et al., 2020) through instruction fine-tuning.
4.2 Implementation Details
We access closed-source models through OpenAI API. For open-source models, we deploy them locally and then access them through FastAPI. We set the temperature to 0 for greedy search in all experiments. In zero-shot-cot, we use Let’s think step by step as the reasoning trigger. To improve answer extraction accuracy, we use answer trigger Therefore, the answer is to prompt models to summarize the answer. In few-shot scenarios, we manually create 2-4 instances, varying with datasets. For more details about demonstrations and instructions, please refer to Appendix A.4.
Method Symbolic Commonsense Event Temporal Overall
TimeXNLI Arith DurationQA McTACO TimeDial SitGen TimeQA MenatQA TempReason TRACIE
s1 s2 s3 Acc EM F1 EM F1 EM F1 Norm E-EM I-EM S-EM O-EM C-EM L2-EM L3-EM Acc Sym. Comm. Event Overall
Human 98 96 92 100.0 64 80.8 75.8 87.1 97.8 - 100 89 87 82 84 76 96 94 82.5 96.5 84.3 88.0 88.5
GPT-4 78.6 76.0 50.7 98.0 35.0 59.2 61.2 80.0 72.0 91.1 59.3 48.9 40.4 44.4 49.0 22.0 91.0 94.0 64.8 75.8 65.4 59.3 63.5
  + CoT 80.0 76.0 60.0 92.0 35.0 58.1 67.0 82.6 65.0 89.3 - 50.0 33.0 43.4 53.0 20.0 93.0 93.0 58.0 77.0 66.2 58.3 63.2
  + FS 85.3 73.3 53.3 100.0 51.0 64.8 77.5 88.3 85.0 94.6 88.6 59.2 40.0 59.6 48.0 25.3 86.0 94.8 62.8 78.0 78.6 63.0 69.5
  + FS/CoT 92.0 84.0 64.0 100.0 42.0 55.1 68.0 72.3 79.0 93.4 - 48.0 44.4 48.5 44.0 22.0 91.0 93.0 66.4 85.0 68.3 60.8 66.5
GPT-3.5 45.4 67.6 31.2 97.0 19.2 50.5 34.1 68.5 39.2 69.1 62.3 60.5 29.5 36.5 37.5 21.0 73.6 61.8 57.4 60.3 49.0 49.8 51.2
  + CoT 33.6 64.8 33.6 71.0 12.4 23.2 28.1 45.1 34.6 67.0 - 52.5 29.0 35.8 38.5 24.0 32.0 54.2 52.0 50.8 35.1 43.5 42.6
  + FS 52.0 68.4 31.6 63.6 42.8 67.7 43.5 71.2 47.8 76.4 79.1 53.8 37.9 37.8 43.5 16.0 77.7 70.0 55.0 53.9 61.2 52.1 54.8
  + FS/CoT 51.6 71.8 36.6 84.4 20.8 41.2 21.4 38.1 48.3 71.1 - 56.5 37.5 38.1 37.5 33.0 86.2 68.0 50.2 61.1 40.2 54.0 51.8
Baichuan2 41.0 61.0 37.0 12.5 4.0 52.0 18.5 63.4 15.0 57.7 52.2 45.0 29.0 31.0 34.0 30.0 40.0 45.0 49.0 37.9 37.5 42.9 40.7
  + CoT 40.0 57.0 31.0 10.0 3.3 44.6 20.0 61.9 13.0 58.1 - 36.0 36.0 39.0 27.0 29.0 46.0 46.0 55.0 34.5 33.5 43.8 39.8
  + FS 43.0 59.0 40.0 42.5 24.7 62.1 27.5 70.2 18.0 58.9 63.7 47.0 35.0 37.0 31.0 19.0 73.0 48.0 48.0 46.1 46.4 47.3 46.9
  + FS/CoT 45.0 54.0 48.0 47.0 10.7 44.4 27.0 68.8 15.0 55.0 - 43.0 27.0 38.0 34.0 33.0 72.8 43.0 44.0 48.5 36.8 46.9 44.7
LLaMA2 30.0 49.0 34.0 22.5 4.0 38.5 8.5 40.6 10.0 35.4 57.9 46.0 21.0 28.0 23.0 18.0 43.0 55.0 49.0 33.9 27.8 40.5 36.1
  + CoT 36.0 50.0 38.0 6.0 7.3 39.1 14.0 51.7 10.0 36.9 - 45.0 30.0 20.0 18.0 21.0 43.0 56.0 47.0 32.5 26.5 40.7 36.0
  + FS 43.0 57.0 60.0 20.5 9.0 46.8 8.0 66.5 15.0 62.3 40.2 24.0 17.0 11.0 5.0 22.0 54.0 50.0 47.0 45.1 35.4 32.6 35.3
  + FS/CoT 37.0 55.0 50.0 33.0 12.0 49.5 11.0 45.6 8.0 44.5 - 35.0 21.0 34.0 23.0 7.0 72.0 54.0 50.0 43.8 28.4 40.9 38.4
Vicuna-1.5 35.0 50.0 36.0 15.0 8.0 39.2 21.5 59.1 7.0 34.2 51.8 43.0 29.0 38.0 22.0 17.0 14.0 13.0 46.0 34.0 31.5 34.2 33.4
  + CoT 42.0 51.0 37.0 3.0 1.3 29.8 11.5 50.0 7.0 33.7 - 44.0 31.0 16.0 25.0 13.0 31.0 29.0 51.0 33.3 22.2 35.2 31.8
  + FS 48.0 57.0 38.0 30.5 7.3 33.6 27.5 57.0 13.0 40.3 39.0 45.0 23.0 38.0 26.0 18.0 51.0 28.0 56.0 43.4 31.1 38.5 37.3
  + FS/CoT 38.0 59.0 39.0 39.5 10.7 37.4 14.0 45.8 12.0 41.6 - 47.0 27.0 39.0 31.0 26.0 71.0 53.0 52.0 43.9 26.9 46.3 41.3
FLAN-T5 53.0 63.0 43.0 0.0 4.0 52.0 14.0 65.0 13.0 47.7 49.5 56.0 24.0 31.0 48.0 20.0 84.0 78.0 64.0 39.8 35.0 51.5 45.3
  + CoT 56.0 66.0 45.0 0.0 4.7 49.7 14.5 63.4 13.0 42.7 - 57.6 23.9 39.0 46.0 28.0 73.0 57.0 55.0 41.8 31.3 49.5 43.9
  + FS 53.0 65.0 43.0 3.5 4.0 50.2 15.5 65.8 9.0 35.0 60.2 55.0 23.0 31.0 47.0 21.0 82.0 78.0 65.0 41.1 34.2 51.3 45.1
  + FS/CoT 54.0 68.0 46.0 3.5 4.0 50.7 13.0 64.0 11.0 43.7 - 54.0 19.0 34.0 45.0 20.0 89.0 72.0 66.0 42.9 31.1 50.7 44.7
Mistral 47.0 50.0 43.0 26.5 11.3 49.8 15.0 58.8 6.0 23.2 58.3 7.0 5.0 4.0 7.0 4.0 2.0 1.0 51.0 41.6 31.8 18.0 25.3
  + CoT 38.0 56.0 35.0 16.5 13.3 36.6 14.5 49.3 8.0 19.3 - 13.0 11.0 8.0 14.0 12.0 5.0 4.0 61.0 36.4 23.5 21.2 24.2
  + FS 51.0 57.0 35.0 18.0 12.0 55.8 25.5 71.0 13.0 52.9 63.0 24.0 12.0 4.0 7.0 7.0 23.0 23.0 57.0 40.3 41.9 25.6 32.2
  + FS/CoT 29.0 62.0 32.0 28.5 9.3 46.6 14.5 49.1 8.0 34.4 - 13.0 11.0 9.0 9.0 9.0 68.0 42.0 50.0 37.9 27.0 31.8 31.6
${}^{\dagger}$ 38.0 48.0 33.0 42.5 20.7 54.8 42.5 73.0 11.0 45.7 64.9 50.0 40.0 42.0 31.0 13.0 68.0 50.0 54.0 40.4 44.6 47.6 45.7
  + CoT 50.0 56.0 34.0 47.0 29.3 62.0 22.5 69.3 12.0 43.8 - 46.0 39.0 39.0 34.0 37.0 73.0 46.0 60.0 46.8 39.8 50.9 47.6
${}^{\dagger}$ 50.0 54.0 30.0 29.5 19.3 53.3 26.5 66.0 20.0 55.6 64.8 48.0 34.0 41.0 38.0 34.0 68.0 49.0 58.0 40.9 43.7 50.0 46.9
  + CoT 40.0 61.0 37.0 52.0 25.3 59.3 26.0 68.8 11.0 40.8 - 46.0 37.0 49.0 34.0 38.0 70.0 55.0 58.0 47.5 38.5 52.5 48.4
${}^{\dagger}$ 48.0 53.0 38.0 41.0 34.0 61.8 42.5 76.2 35.0 61.8 58.3 43.0 30.0 37.0 38.0 37.0 68.0 64.0 53.0 45.0 52.8 51.0 50.6
  + CoT 57.0 63.0 35.0 54.0 30.0 61.8 42.0 45.7 29.0 57.3 - 51.0 30.0 48.0 37.0 24.0 60.0 58.0 57.0 52.3 44.3 49.6 48.7
Table 4: Main results of TimeBench. Aligned models are under the zero-shot setting by default. Methods with † are base models without alignment, under few-shot setting, thus incomparable with other methods. Some F1 scores are omitted, please refer to the appendix for details. Best results in each group are bold, top-3 results are underlined.
5 Experimental Results
5.1 Main Results
We present the experimental results [^2] Please refer to Appendix A.3 for full results. of the aligned models in Table 4, covering the most popular LLMs. In most tasks, GPT-4 achieved the best performance, ranking first in 16 out of 19 evaluation metrics, surpassing the second-place GPT-3.5 by 14.7%. Nevertheless, it still exhibits a significant gap from human performance, with a 19% disparity. Among all open-source alignment models, Baichuan2 has achieved the best overall performance, reaching 85% performance of ChatGPT.
Results on Symbolic Temporal Reasoning
GPT-4 demonstrates strong symbolic reasoning performance, far surpassing ChatGPT and other models. While open-source models show poorer performance, with only Mistral-base achieving an overall score exceeding 50. In TimeX-NLI tasks, models other than GPT-4 perform poorly, barely surpassing a random baseline (33.3 for s1/s3, 50.0 for s2). This indicates that LLMs still have significant shortcomings in understanding abstract TimeX.
Results on Commonsense Temporal Reasoning
In temporal commonsense reasoning, GPT-4 still maintains a remarkable proximity to human-level performance. Among the open-source aligned models, Baichuan2 attains the best performance, securing the top three positions in four out of seven metrics. When extending to all open-source models, Mistral-base significantly outperforms others, surpassing the best open-source model by 6.4. Besides, we observe that chain-of-thought prompting has a negative impact on almost all models in commonsense reasoning. Furthermore, aligned models appear to exhibit varying degrees of performance degradation compared to the base model, with LLaMA2 and Mistral showing the most significant differences.
Results on Event Temporal Reasoning
GPT-4 remains the top performer in event temporal reasoning, though it is no longer markedly ahead of other models, and there is a notable disparity from human performance. Among all open-source models, LLaMA2-base achieved the best results, trailing GPT-3.5 by just 1.5%. Overall, open-source models show competitive performance in event temporal reasoning, narrowing the gap with closed-source models.
In TRACIE, which involves implied events, most models only marginally surpass the random baseline (50.0), indicating persistent challenges in understanding implicit events with current models. Additionally, there is a large performance gap comparing context-based reasoning (TimeQA, MenatQA) with fact-based reasoning (TempReason), indicating a substantial potential for improvement in event-time facts grounding in unstructured text. Furthermore, within fact-based reasoning, models exhibit a noticeable decline in event-event reasoning compared to event-time reasoning, indicating that complex multi-hop relations among events increased the difficulty of reasoning.
Figure 3: Performance difference between base and aligned models. Baichuan2 and LLaMA2 are aligned with SFT and RLHF. Vicuna, Mistral and ChatGLM3 are aligned with only SFT.
6 Analysis and Discussion
6.1 Alignment Degradation
Alignment Causes Performance Degradation
We notice that some aligned models exhibited performances inconsistent with their expectations, such as LLaMA2 and Mistral. We speculate that this is the negative impact brought about by the alignment phase. To investigate this, we conduct experiments on their corresponding base models in the few-shot setting, as shown in Table 4 (grey).
We visualize the performance differences between the base and alignment models in the few-shot scenario, as shown in Figure 3. Baichuan2-13B shows a slight decrease in performance after alignment by 0.7%, while LLaMA2-13B, Vicuna-1.5-13B and Mistral-7B exhibit catastrophic performance drops, decreasing 10.0%, 7.1% and 18.4%, respectively.
What leads to the degradation?
Combining experimental results with model output cases, we attribute this phenomenon to two reasons:
  • The training during the alignment phase has to some extent disrupted the original language modeling capability, causing the model to deviate from strictly adhering to demonstrations during in-context learning.
  • Excessive alignment of models towards conversation and harmlessness has led to a decrease in its usability, resulting in generating instruction-irrelevant content or refusing to follow instructions.
Figure 4: Performance gap with and without CoT prompts. The results are averaged from GPT-4, GPT-3.5, Baichuan2, LLaMA2/Vicuna-1.5 and Mistral.
6.2 Chain-of-Thought in Temporal Reasoning
Previous research has found that chain-of-thought prompting (Wei et al., 2022; Kojima et al., 2022; Chu et al., 2023a) enables models to think step by step, facilitating the resolution of complex reasoning tasks such as mathematical and logical reasoning. How about applying cot prompting to temporal reasoning, will it bring consistent improvement? To investigate this, we select several popular models and analyze their performance affected when using chain-of-thought prompts across various task categories.
Chain-of-Thought inconsistently boosts, or even impairs performance
As illustrated in Figure 4, introducing zero-shot chain-of-thought prompting results in an overall performance decline of 6%, with even a 16% degradation in commonsense reasoning. In the few-shot scenario, chain-of-thought prompts also fail to yield comprehensive improvements, showing varied effects depending on the task type, resulting in a 13% improvement in symbolic reasoning but an 18% decrease in commonsense reasoning. In event temporal reasoning, cot prompting in both scenarios only resulted in minor impacts, with -1% and 1.7%, respectively. In summary, zero-shot chain-of-thought consistently hinders performance, whereas few-shot chain-of-thought demonstrates performance variations dependent on the nature of the task.
Figure 5: $\Delta\text{Score}$ between the chain-of-thought prompting and direct I-O prompting. Top: zero-shot setting, Bottom: few-shot setting, Left: variation in each task, Right: averaged variation in the symbolic, commonsense, event, and overall tasks.
Impact of Chain-of-Thought across tasks
In order to explore the impact of CoT on various tasks thoroughly, we delve into the performance changes of each model across specific tasks within each category, as illustrated in Figure  5. In the zero-shot setting, open-source models achieve a slight improvement in event temporal reasoning with chain-of-thought prompting, while in other cases, they face performance degradation. While in the few-shot setting, almost all models exhibit significant improvement in symbolic temporal reasoning, with a concurrent prevalent decline in commonsense temporal reasoning.
Differing from the previous uniform trends observed within the same task category, event temporal reasoning exhibits both increases and decreases simultaneously. We attribute this to the nature of the datasets; both MenatQA and TempReason involve implicit reasoning, encompassing temporal factors and event-event relationships, requiring multi-hop reasoning for resolution. This aligns well with the step-by-step problem-solving nature of cot prompting, resulting in performance improvement. On the other hand, TimeQA only necessitates implicit reasoning for a subset of questions, making cot prompts less impactful. This is reflected in the marginal performance change with and without chain-of-thought prompting.
6.3 Impact of Model Size
We investigate the relationship between the performance and parameter size of base models with decoder architecture ranging from 6B to 13B. With the scaling up of the parameter size for the same model, there is a notable performance enhancement. When the parameter size expands from 7B to 13B, LLaMA2 and Baichuan2 show performance improvements of 14.2% and 11.1%, respectively. Under the same 13B parameter size, LLaMA2 slightly outperforms Baichuan2 by 1.6%. With around 7B parameter size, there is no significant difference in performance among LLaMA2, Baichuan2, ChatGLM3, while Mistral exhibits extremely strong performance, even surpassing LLaMA2-13B by as much as 4.5% with only 53% parameters.
6.4 Dive into Each Categories
Model Order Duration Freq. Stationarity Typical Avg.
GPT-4 76.4 $\downarrow$ 92.8 $\uparrow$ 83.3 $\uparrow$ 71.4 $\downarrow$ 54.5 $\downarrow$ 77.5
GPT-3.5 50.5 $\uparrow$ 39.8 $\downarrow$ 55.2 $\uparrow$ 48.4 $\uparrow$ 28.7 $\downarrow$ 43.5
${}^{\dagger}$ 40.5 $\downarrow$ 51.8 $\uparrow$ 43.7 $\uparrow$ 46.2 $\uparrow$ 29.8 $\downarrow$ 42.5
${}^{\dagger}$ 18.9 $\downarrow$ 37.0 $\uparrow$ 27.1 $\uparrow$ 7.6 $\downarrow$ 25.5 $\downarrow$ 26.5
${}^{\dagger}$ 27.0 $\downarrow$ 44.4 $\uparrow$ 58.3 $\uparrow$ 38.5 $\downarrow$ 38.3 $\downarrow$ 42.5
Table 5: Results in each temporal commonsense aspect under few-shot setting. Models with †  are base models. Red $\downarrow$ and Green $\uparrow$ represent the performance is lower or higher than its average.
LLM underperforms in (multi-hop) symbolic reasoning
Except for GPT-4, the performance of other models in symbol temporal reasoning is unsatisfactory. Our experimental results reveal a notable decline in models’ performance on the duration-conversion task (s3) compared to other atomic tasks (s1, s2, arith). Delving into the reasons, task s3 necessitates a two-step reasoning process, initially unifying time units and subsequently engaging in numerical comparison. In contrast, other atomic tasks can be completed with a single inference. To sum up, LLMs show unsatisfactory performance in symbol temporal reasoning, with a more pronounced decline when confronted with multi-step reasoning.
LLM’s mastery of temporal commonsense knowledge differs across aspects.
We analyze models’ performance across various common-sense aspects, as shown in Table 5. We use the model’s average performance in common-sense tasks as a baseline, if the model outperforms its own average in a specific aspect, it suggests greater proficiency and vice versa. The findings show that models generally demonstrate good knowledge of event duration and frequency, while their grasp of event order and typical events is comparatively weaker.
LLM demonstrates poor implicit temporal reasoning abilities
Comparing explicit and implicit context-based reasoning (TimeQA-explicit v.s. others), we note a marked decrease in the model’s performance in implicit reasoning. Additionally, on TRACIE with numerous implied events, most models exceed only a random baseline (50.0). Even GPT-4 achieves a mere 66.4% accuracy. To summarize, LLM performs inadequately when dealing with different types of implicit reasoning, constraining its reasoning abilities in complex scenarios.
LLM is a good factual reasoner rather than factual extractor
When humans engage in context-based temporal reasoning, it typically involves two steps: initially extracting time-fact pairs from the context and subsequently conducting fact-based reasoning. TempReason provides extracted facts for conducting fact-based reasoning. By comparing the model’s performance in context-based reasoning and fact-based reasoning, we can identify the bottlenecks in LLM’s event temporal reasoning. Our findings reveal that LLM excels in TempReason, signifying its strong capability in fact-based reasoning. However, LLM’s performance in context-based reasoning was significantly weaker than in the former. We attribute this gap to the model’s deficiency in factual extraction capabilities. This gap highlights limitations in context-based temporal reasoning.
7 Related Work
7.1 Temporal Reasoning
Numerous research efforts address diverse challenges in temporal reasoning. Early research mainly relies on TimeML (Pustejovsky et al., 2003), focusing TimeX extraction and temporal relation extraction (Verhagen et al., 2007, 2010; UzZaman et al., 2013; Llorens et al., 2015; Miller et al., 2015; Mathur et al., 2021; Vashishtha et al., 2019). The advent of pre-trained language models (PLMs) has brought about commonsense reasoning as a tool to explore the world knowledge in models (Zhou et al., 2019; Qin et al., 2021; Dhingra et al., 2022). Recently, much attention has shifted towards event temporal reasoning (Chen et al., 2021; Tan et al., 2023; Zhu et al., 2023; Son and Oh, 2023; Chu et al., 2023b). Besides, Wang and Zhao (2023) introduces a unified form for assessing the temporal understanding capability of language models.
Distinguished from other datasets and benchmarks, TimeBench is multispectral, and closely aligned with intricate real-world scenarios, offering a comprehensive and hierarchical evaluation of LLM’s temporal reasoning abilities.
7.2 Large-scaled Language Model
In recent years, there has been rapid progress in the research of large-scale language models (LLM) (Zhao et al., 2023). They exhibit outstanding performance across a multitude of tasks without the need for fine-tuning (Brown et al., 2020; Kojima et al., 2022). Furthermore, they have achieved astonishing results in complex reasoning tasks, such as mathematical reasoning (Cobbe et al., 2021; Mishra et al., 2022) and logical reasoning (Yu et al., 2020; Liu et al., 2023). Moreover, some studies suggest that the chain-of-thought prompting can further enhance the model’s capabilities in complex reasoning scenario (Wei et al., 2022; Kojima et al., 2022; Chu et al., 2023a).
8 Conclusion
Time is a crucial dimension for understanding the world, and temporal reasoning inherently possesses complexity. The lack of a comprehensive benchmark makes it challenging to quantify LLMs’ temporal reasoning capabilities. In this work, we present TimeBench, a comprehensive and hierarchical benchmark for LLM temporal reasoning, tailored to mirror temporal reasoning in complex real-world scenarios. We conduct extensive experiments on state-of-the-art LLMs to investigate their temporal reasoning capabilities. Our findings indicate a substantial gap between state-of-the-art LLM and human performance, emphasizing the need for further research in this area. Moreover, we provide a meticulous analysis and discussion, outlining the current challenges that models face and suggesting potential directions for improvement.
Appendix A Appendix
A.1 Benchmark Details
We present the subtasks for each dataset in TimeBench, including data quantities, task formats, challenges faced by each task, and the capabilities required by models, as shown in Table 6.
A.2 Metrics for CTG
Following SituatedGen (Zhang and Wan, 2023), we use BLEU-4 (Papineni et al., 2002), METEOR (Banerjee and Lavie, 2005), ROUGE-L (Lin, 2004), CIDEr(Vedantam et al., 2015), and MATCH (Zhang and Wan, 2023) scores to metric the results of Constrained Text Generation. [^3] We utilize pycocoevalcap package to calucate BLEU-4, METEOR, ROUGE-L, CIDEr.
The overall score is calculated as the sum of the above scores. We set the weight of CIDEr to $1/10$ for balance when summing.
$S=\textrm{BLEU-4}+\textrm{METEOR}+\textrm{ROUGE-L}\\ +\textrm{CIDER}/10+\textrm{MATCH}$
As the overall score $S$ does not represent a percentile, we proceeded to normalize the models’ scores to align with humans’ relative performance levels.
A.3 Full Results
The overall score is derived from the average of all corresponding metrics. For brevity, we omit some F1 scores in the main tables in the main text. Please refer to Table 7 for the full experimental results, and detailed metrics for SituatedGen can be found in Table 8.
A.4 Prompts
The prompt instructions are showcased in Figure 6. The demonstrations can be found from Figure 7 to Figure 15.
Dataset Format # Challenges
Symbolic
TimeX Arith FRC 4,000 TimeX Arithmetic
TimeX NLI NLI 6,965 TimeX Causality
   - Order - 2,213 order
   - Duration - 2,332 duration
   - Conversion - 2,420 duration + time unit conversion
Commonsense
MCTACO MCMAQ 852 Temporal Commonsense
TimeDial MCMAQ 1,446 Temporal Commonsense
DurationQA MCMAQ 687 Event Duration
SituatedGen CTG 115 Temporal Commonsense
Event
TimeQA FRC 1,000 Context-based Reasoning
   - Explicit - 500 explicit, event-time reasoning
   - Implicit - 500 implicit, event-time reasoning
MenatQA FRC 1,599 Implicit, Context-based Reasoning
   - Order - 400 event-time reasoning
   - Scope - 400 event-time reasoning
   - Counterfactual - 400 event-time reasoning
TempReason FRC 1,876 Implicit, Fact-based Reasoning
   - l2 (e2t) - 839 event-time reasoning
   - l3 (e2e) - 1,037 event-event reasoning
TRACIE NLI 500 Implicit, Implied Event-Event Reasoning
In total 19,000
Table 6: The task formats and challenges in TimeBench.
Table 7: Full results of TimeBench. Aligned models are under zero-shot setting by default. Methods with † are base models without alignment, under few-shot setting, thus incomparable with other methods. We consider human performance as 100 points and normalize models’ results accordingly.
Method Symbolic Commonsense Event Overall
TimeXNLI Date Arith DurationQA McTACO TimeDial SitGen TimeQA MenatQA TempReason TRACIE
s1 s2 s3 Acc EM F1 EM F1 EM F1 Norm E-EM E-F1 H-EM H-F1 S-EM S-F1 O-EM O-F1 C-EM C-F1 L2-EM L2-F1 L3-EM L3-F1 Acc Sym. Comm. Event Overall
Human 98 96 92 100 64 80.8 75.8 87.1 97.8 - 100 89 93.3 87 91.1 82 85.6 84 87.3 76 79.9 96 97.1 94 95.3 82.5 96.5 84.3 88.0 88.5
GPT-4 78.6 76.0 50.7 98.0 35.0 59.2 61.2 80.0 72.0 91.1 59.3 48.9 60.6 40.4 46.5 44.4 57.0 49.0 57.0 22.0 23.1 91.0 95.3 94.0 95.0 64.8 75.8 65.4 59.3 63.5
  + CoT 80.0 76.0 60.0 92.0 35.0 58.1 67.0 82.6 65.0 89.3 - 50.0 61.3 33.0 41.2 43.4 54.6 53.0 59.6 20.0 22.6 93.0 97.0 93.0 94.5 58.0 77.0 66.2 58.3 63.2
  + FS 85.3 73.3 53.3 100 51.0 64.8 77.5 88.3 85.0 94.6 88.6 59.2 73.7 40.0 51.0 59.6 72.4 48.0 54.8 25.3 28.7 86.0 92.4 94.8 95.9 62.8 78.0 78.6 63.0 69.5
  + FS/CoT 92.0 84.0 64.0 100.0 42.0 55.1 68.0 72.3 79.0 93.4 - 48.0 66.9 44.4 52.8 48.5 65.3 44.0 52.6 22.0 25.9 91.0 96.9 93.0 94.6 66.4 85.0 68.3 60.8 66.5
GPT-3.5 45.4 67.6 31.2 97.0 19.2 50.5 34.1 68.5 39.2 69.1 62.3 60.5 70.8 29.5 35.4 36.5 40.9 37.5 43.9 21.0 22.9 73.6 81.2 61.8 73.8 57.4 60.3 49.0 49.8 51.2
  + CoT 33.6 64.8 33.6 71.0 12.4 23.2 28.1 45.1 34.6 67.0 - 52.5 64.4 29.0 35.1 35.8 39.7 38.5 42.9 24.0 26.3 32.0 57.6 54.2 68.0 52.0 50.8 35.1 43.5 42.6
  + FS 52.0 68.4 31.6 63.6 42.8 67.7 43.5 71.2 47.8 76.4 79.1 53.8 66.1 37.9 48.4 37.8 43.2 43.5 51.6 16.0 17.9 77.7 84.7 70.0 78.0 55.0 53.9 61.2 52.1 54.8
  + FS/CoT 51.6 71.8 36.6 84.4 20.8 41.2 21.4 38.1 48.3 71.1 - 56.5 68.0 37.5 47.0 38.1 42.5 37.5 41.7 33.0 37.8 86.2 89.9 68.0 76.6 50.2 61.1 40.2 54.0 51.8
Baichuan2 41.0 61.0 37.0 12.5 4.0 52.0 18.5 63.4 15.0 57.7 52.2 45.0 55.4 29.0 34.6 31.0 48.8 34.0 44.3 30.0 39.5 40.0 57.4 45.0 61.4 49.0 37.9 37.5 42.9 40.7
  + CoT 40.0 57.0 31.0 10.0 3.3 44.6 20.0 61.9 13.0 58.1 - 36.0 41.5 36.0 40.9 39.0 52.0 27.0 38.5 29.0 43.2 46.0 62.8 46.0 64.3 55.0 34.5 33.5 43.8 39.8
  + FS 43.0 59.0 40.0 42.5 24.7 62.1 27.5 70.2 18.0 58.9 63.7 47.0 60.7 35.0 45.7 37.0 51.9 31.0 41.5 19.0 31.8 73.0 81.1 48.0 59.4 48.0 46.1 46.4 47.3 46.9
  + FS/CoT 45.0 54.0 48.0 47.0 10.7 44.4 27.0 68.8 15.0 55.0 - 43.0 57.8 27.0 36.7 38.0 49.8 34.0 40.7 33.0 43.0 72.8 80.4 43.0 60.2 44.0 48.5 36.8 46.9 44.7
LLaMA2 30.0 49.0 34.0 22.5 4.0 38.5 8.5 40.6 10.0 35.4 57.9 46.0 61.9 21.0 30.5 28.0 46.1 23.0 36.1 18.0 26.9 43.0 53.1 55.0 69.4 49.0 33.9 27.8 40.5 36.1
  + CoT 36.0 50.0 38.0 6.0 7.3 39.1 14.0 51.7 10.0 36.9 - 45.0 58.7 30.0 38.9 20.0 40.9 18.0 32.5 21.0 33.6 43.0 58.0 56.0 68.4 47.0 32.5 26.5 40.7 36.0
  + FS 43.0 57.0 60.0 20.5 9.0 46.8 8.0 66.5 15.0 62.3 40.2 24.0 34.2 17.0 18.4 11.0 25.9 5.0 14.6 22.0 33.3 54.0 68.1 50.0 64.8 47.0 45.1 35.4 32.6 35.3
  + FS/CoT 37.0 55.0 50.0 33.0 12.0 49.5 11.0 45.6 8.0 44.5 - 35.0 46.0 21.0 25.4 34.0 46.7 23.0 36.5 7.0 16.5 72.0 80.8 54.0 66.2 50.0 43.8 28.4 40.9 38.4
Vicuna-1.5 35.0 50.0 36.0 15.0 8.0 39.2 21.5 59.1 7.0 34.2 51.8 43.0 60.4 29.0 37.0 38.0 46.8 22.0 37.4 17.0 23.2 14.0 42.1 13.0 43.6 46.0 34.0 31.5 34.2 33.4
  + CoT 42.0 51.0 37.0 3.0 1.3 29.8 11.5 50.0 7.0 33.7 - 44.0 56.9 31.0 36.4 16.0 38.2 25.0 37.7 13.0 20.4 31.0 49.0 29.0 49.1 51.0 33.3 22.2 35.2 31.8
  + FS 48.0 57.0 38.0 30.5 7.3 33.6 27.5 57.0 13.0 40.3 39.0 45.0 58.3 23.0 25.9 38.0 42.6 26.0 41.4 18.0 20.1 51.0 61.8 28.0 42.6 56.0 43.4 31.1 38.5 37.3
  + FS/CoT 38.0 59.0 39.0 39.5 10.7 37.4 14.0 45.8 12.0 41.6 - 47.0 59.5 27.0 30.7 39.0 48.1 31.0 35.9 26.0 31.2 71.0 77.5 53.0 65.5 52.0 43.9 26.9 46.3 41.3
FLAN-T5 53.0 63.0 43.0 0.0 4.0 52.0 14.0 65.0 13.0 47.7 49.5 56.0 61.7 24.0 26.8 31.0 33.6 48.0 52.2 20.0 21.8 84.0 87.9 78.0 83.9 64.0 39.8 35.0 51.5 45.3
  + CoT 56.0 66.0 45.0 0.0 4.7 49.7 14.5 63.4 13.0 42.7 - 57.6 64.4 23.9 28.2 39.0 41.6 46.0 50.2 28.0 30.6 73.0 79.5 57.0 68.9 55.0 41.8 31.3 49.5 43.9
  + FS 53.0 65.0 43.0 3.5 4.0 50.2 15.5 65.8 9.0 35.0 60.2 55.0 64.3 23.0 25.0 31.0 33.6 47.0 50.6 21.0 22.4 82.0 87.0 78.0 84.5 65.0 41.1 34.2 51.3 45.1
  + FS/CoT 54.0 68.0 46.0 3.5 4.0 50.7 13.0 64.0 11.0 43.7 - 54.0 59.4 19.0 21.2 34.0 36.6 45.0 49.2 20.0 21.7 89.0 93.8 72.0 79.7 66.0 42.9 31.1 50.7 44.7
Mistral 47.0 50.0 43.0 26.5 11.3 49.8 15.0 58.8 6.0 23.2 58.3 7.0 28.2 5.0 21.4 4.0 24.3 7.0 22.3 4.0 21.7 2.0 39.6 1.0 31.6 51.0 41.6 31.8 18.0 25.3
  + CoT 38.0 56.0 35.0 16.5 13.3 36.6 14.5 49.3 8.0 19.3 - 13.0 31.3 11.0 22.4 8.0 21.1 14.0 24.9 12.0 25.6 5.0 34.0 4.0 31.2 61.0 36.4 23.5 21.2 24.2
  + FS 51.0 57.0 35.0 18.0 12.0 55.8 25.5 71.0 13.0 52.9 63.0 24.0 43.5 12.0 23.9 4.0 21.3 7.0 21.2 7.0 23.0 23.0 48.9 23.0 44.9 57.0 40.3 41.9 25.6 32.2
  + FS/CoT 29.0 62.0 32.0 28.5 9.3 46.6 14.5 49.1 8.0 34.4 - 13.0 33.7 11.0 25.3 9.0 26.4 9.0 24.3 9.0 20.3 68.0 78.6 42.0 57.4 50.0 37.9 27.0 31.8 31.6
ChatGLM3 38.0 50.0 34.0 2.0 3.0 34.1 7.0 43.6 14.0 56.7 38.9 20.0 41.2 14.0 31.7 25.0 33.8 17.0 26.0 24.0 32.2 42.0 57.0 30.0 54.0 50.0 31.0 28.2 33.2 31.5
  + CoT 27.0 49.0 37.0 0.0 1.0 24.8 3.0 37.1 10.0 44.8 - 24.0 41.7 10.0 25.4 27.0 34.6 22.0 28.1 35.0 41.2 28.0 44.5 27.0 52.0 48.0 28.3 20.1 32.6 28.9
  + FS 37.0 52.0 30.0 0.0 10.0 53.0 9.0 52.9 19.0 54.7 50.1 11.0 19.6 13.0 17.0 25.0 35.0 24.0 30.4 20.0 25.6 27.0 46.5 33.0 53.3 54.0 29.8 35.5 29.0 30.9
  + FS/CoT 32.0 66.0 31.0 0.0 4.0 34.8 8.0 43.6 11.0 44.0 - 25.0 43.3 19.0 27.5 23.0 30.2 19.0 25.7 36.0 43.0 50.0 62.8 34.0 56.0 48.0 32.3 24.2 36.2 32.7
Baichuan2-7B 35.0 50.0 37.0 4.5 4.0 47.9 10.5 55.3 15.0 54.3 42.0 26.0 41.5 20.0 34.7 20.0 35.1 19.0 31.2 6.0 20.4 22.0 43.4 29.0 47.7 55.0 31.6 32.7 30.1 31.0
  + CoT 38.0 43.0 32.0 1.0 5.3 37.9 13.0 58.0 15.0 44.2 - 41.0 53.5 28.0 38.8 29.0 39.9 23.0 33.2 18.0 29.3 21.0 41.2 29.0 47.2 54.0 28.5 28.9 35.1 32.5
  + FS 40.0 50.0 36.0 20.0 28.7 59.4 26.5 66.9 17.0 53.0 49.8 45.0 60.7 30.0 42.1 27.0 37.8 23.0 35.7 10.0 20.4 40.0 57.4 37.0 53.0 51.0 36.5 43.0 38.0 39.1
  + FS/CoT 41.0 50.0 36.0 23.5 13.0 45.7 17.5 58.1 7.0 39.2 - 36.0 51.2 29.0 43.0 42.0 52.5 25.0 39.3 20.0 31.0 57.0 70.1 39.0 60.2 49.0 37.6 30.1 43.0 39.0
LLaMA2-7B 39.0 53.0 30.0 13.0 2.7 39.3 4.0 41.0 1.0 6.3 24.5 37.0 49.0 14.0 29.0 7.0 26.8 8.0 21.1 9.0 16.0 48.0 63.9 32.0 47.9 49.0 33.8 17.0 30.5 27.4
  + CoT 44.0 50.0 33.0 5.0 2.7 35.0 4.5 40.0 1.0 1.7 - 27.0 49.9 17.0 31.6 11.0 31.4 10.0 24.5 7.0 17.8 44.0 56.9 32.0 48.1 46.0 33.0 14.1 30.3 26.8
  + FS 44.0 60.0 34.0 11.0 4.0 62.8 8.0 64.7 8.0 40.0 30.5 36.0 50.8 20.0 29.4 5.0 22.3 6.0 18.0 6.0 17.9 12.0 36.3 23.0 44.3 53.0 37.3 31.1 25.3 28.7
  + FS/CoT 38.0 51.0 36.0 14.5 11.0 42.8 25.0 65.6 13.0 53.4 - 36.0 53.5 21.0 34.1 1.0 13.6 3.0 11.2 5.0 14.0 22.0 46.7 21.0 42.3 51.0 34.9 35.1 25.0 29.0
Vicuna-1.5-7B 37.0 58.0 43.0 5.0 1.3 40.4 9.5 52.5 6.0 32.0 47.8 35.0 47.1 11.0 18.5 20.0 35.7 15.0 25.7 12.0 17.3 14.0 33.0 14.0 46.8 54.0 35.8 27.1 26.6 28.1
  + CoT 36.0 50.0 36.0 1.5 1.3 39.4 8.5 49.2 9.0 36.2 - 30.0 40.9 14.0 24.6 16.0 26.2 14.0 28.5 12.0 25.0 9.0 27.7 7.0 40.3 54.0 30.9 23.9 24.6 25.5
  + FS 43.0 57.0 37.0 8.5 3.3 44.6 5.5 42.1 7.0 36.8 67.0 24.0 31.9 12.0 14.9 16.0 21.8 20.0 27.5 17.0 22.2 13.0 34.3 6.0 32.2 54.0 36.4 29.5 23.1 26.9
  + FS/CoT 35.0 54.0 35.0 8.0 2.7 37.2 10.0 47.5 10.0 41.3 - 31.0 39.9 13.0 16.6 15.0 26.7 15.0 23.8 16.0 23.1 55.0 66.3 32.0 48.1 43.0 33.0 24.8 31.0 29.8
${}^{\dagger}$ 38.0 48.0 33.0 42.5 20.7 54.8 42.5 73.0 11.0 45.7 64.9 50.0 59.4 40.0 54.2 42.0 52.7 31.0 38.0 13.0 21.4 68.0 77.3 50.0 63.5 54.0 40.4 44.6 47.6 45.7
  + CoT 50.0 56.0 34.0 47.0 29.3 62.0 22.5 69.3 12.0 43.8 - 46.0 58.2 39.0 49.6 39.0 49.8 34.0 40.1 37.0 45.6 73.0 81.3 46.0 65.6 60.0 46.8 39.8 50.9 47.6
${}^{\dagger}$ 50.0 54.0 30.0 29.5 19.3 53.3 26.5 66.0 20.0 55.6 64.8 48.0 59.3 34.0 48.6 41.0 49.6 38.0 43.4 34.0 37.5 68.0 78.7 49.0 62.7 58.0 40.9 43.7 50.0 46.9
  + CoT 40.0 61.0 37.0 52.0 25.3 59.3 26.0 68.8 11.0 40.8 - 46.0 59.4 37.0 49.1 49.0 58.4 34.0 43.8 38.0 44.1 70.0 78.0 55.0 68.2 58.0 47.5 38.5 52.5 48.4
${}^{\dagger}$ 27.0 66.0 41.0 32.5 28.0 59.8 34.5 69.4 5.0 34.3 59.8 40.0 53.8 35.0 50.1 41.0 49.6 33.0 38.5 18.0 22.9 49.0 65.9 34.0 51.0 55.0 41.6 41.5 42.4 42.1
  + CoT 30.0 56.0 34.0 34.0 23.3 57.0 33.0 69.5 12.0 44.5 - 41.0 51.2 31.0 40.7 38.0 46.4 26.0 32.6 41.7 46.3 46.0 61.5 43.8 64.1 53.0 38.5 39.9 44.2 42.3
${}^{\dagger}$ 26.0 50.0 30.0 20.0 19.3 54.5 20.0 59.6 15.0 45.2 62.4 44.0 54.4 30.0 45.3 42.0 49.8 34.0 41.9 30.0 35.8 50.0 64.0 36.0 53.3 49.0 31.5 39.4 44.0 40.8
  + CoT 37.0 52.0 36.0 25.5 21.3 56.9 26.5 67.0 16.0 41.9 - 32.0 45.6 27.0 36.1 41.0 50.9 30.0 38.0 51.0 57.3 45.0 59.7 37.0 57.7 50.0 37.6 38.3 43.9 41.5
${}^{\dagger}$ 48.0 53.0 38.0 41.0 34.0 61.8 42.5 76.2 35.0 61.8 58.3 43.0 55.9 30.0 45.3 37.0 49.4 38.0 47.8 37.0 45.5 68.0 76.7 64.0 74.8 53.0 45.0 52.8 51.0 50.6
  + CoT 57.0 63.0 35.0 54.0 30.0 61.8 42.0 45.7 29.0 57.3 - 51.0 60.4 30.0 46.2 48.0 57.2 37.0 47.9 24.0 33.2 60.0 65.9 58.0 67.9 57.0 52.3 44.3 49.6 48.7
${}^{\dagger}$ 48.0 70.0 32.0 35.0 3.3 51.8 13.5 62.6 11.0 55.0 61.6 50.0 57.2 24.0 26.3 30.0 35.4 38.0 41.5 22.0 22.5 67.0 76.4 35 55.9 58.0 46.3 37.0 42.6 41.6
  + CoT 47.0 68.0 32.0 46.0 8.7 53.9 15.5 64.3 13.0 56.5 45.0 52.5 23.0 24.5 30.0 35.0 37.0 40.2 22.0 22.5 72.0 79.4 42 60.3 54.0 48.3 35.3 42.6 41.8
Method BLEU-4 METEOR ROUGE-L CIDEr MATCH Overall Norm
Human 39.9 40.4 56.3 397 98.1 274.4 100.0
GPT4 8.23 31.27 28.84 38.45 90.41 162.59 59.25
  + FS 28.64 38.99 55.69 298.64 90.11 243.29 88.66
GPT3.5 13.38 30.12 35.91 125.41 78.76 170.70 62.21
  + FS 27.24 33.77 51.18 282.75 76.54 217.01 79.08
Baichuan2 8.33 25.86 30.07 82.63 70.63 143.15 52.17
  + FS 15.79 30.23 40.96 169.14 71.01 174.91 63.74
LLaMA2 4.66 21.43 20.80 17.72 61.62 110.28 40.19
  + FS 15.15 27.49 37.55 138.13 64.94 158.93 57.92
Vicuna1.5 7.73 26.35 29.15 69.16 71.91 142.06 51.77
  + FS 6.85 18.66 25.99 92.96 46.19 106.99 38.99
FLANT5 16.20 24.43 29.38 95.17 56.38 135.91 49.53
  + FS 12.88 30.38 36.27 92.20 76.44 165.19 60.20
Mistral 5.82 22.89 24.19 44.03 63.74 121.03 44.11
  + FS 18.96 29.02 43.15 185.61 63.24 172.93 63.02
ChatGLM3 6.56 21.11 21.96 41.48 53.02 106.80 38.92
  + FS 10.53 24.17 33.44 124.50 56.94 137.53 50.12
Baichuan2-7b 5.17 21.99 23.73 44.80 59.85 115.22 41.99
  + FS 15.06 23.45 32.29 137.94 52.04 136.64 49.79
LLaMA2-7b 2.77 13.46 14.69 14.34 34.83 67.18 24.48
  + FS 6.90 15.82 21.77 52.99 33.81 83.60 30.47
Vicuna1.5-7b 6.29 24.34 26.91 46.90 68.84 131.07 47.77
  + FS 20.71 30.19 45.20 203.20 67.58 184.00 67.05
${}^{\dagger}$ 17.86 29.75 44.28 198.83 66.35 178.12 64.91
${}^{\dagger}$ 17.54 29.44 45.21 200.14 65.64 177.84 64.81
${}^{\dagger}$ 15.30 27.54 41.80 171.59 62.40 164.20 59.84
${}^{\dagger}$ 17.49 28.33 45.24 202.08 59.98 171.25 62.41
${}^{\dagger}$ 14.54 27.39 41.72 168.89 59.42 159.96 58.30
${}^{\dagger}$ 17.11 29.35 40.74 156.49 66.18 169.02 61.60
Table 8: Full results of SituatedGen. Aligned models are under zero-shot setting by default. The top-3 results are bold. Methods with † are base models without alignment, under few-shot setting, thus incomparable with other methods. We consider human performance as 100 points and normalize models’ results accordingly.
DurationQA, MCTACO
Answer the following question, select all the possible correct options, and each question has at least one correct option.
Context: {}
Question: {}
Options: {}
Answer:
TimeDial
There is a two-person dialogue with several options.
Choose all appropriate options to substitute the <mask> in the dialogue, and each question has at least one correct option.
Dialogue: {}
Options: {}
Answer:
TRACIE
Read the following story and hypothesis, determine whether the hypothesis can be inferred from the story.
You need to understand the implicit temporal relationships between events to make judgments.
Story: {}
Hypothesis: {}
Options: A. Entailment B. Contradiction
Answer:
SituatedGen
Generate a pair of contrastive sentences with the given set of keywords.
Keywords: {}
Date Arithmetic
Question: {}? Answer:
TimeQA
I will give you a question with context.
You need to answer my question based on the context.
If you can infer the answer from the context, then output your answer. Otherwise, if there is no answer, output [unanswerable].
Context: {}
Question: {}
Answer:
TempReason
I will give you a question with context.
You need to answer my question based on the context.
Context: {}
Question: {}
Answer:
MenatQA
Get answers for the question based on the contxt, where answers derived from substrings in the context or categorized as [unanswerable].
Context: {}
Question: {}
Answer:
TimeX-NLI
Read the following statements about time and determine if the hypothesis can be inferred from the premise.
Premise: {}
Hypothesis: {}
Options: A. Entailment B. Contradiction C. Neutral
Answer:
Figure 6: Zeroshot instructions and input formats.
CoT Demonstration of TimeX-NLI (3-shot, order)
Answer the following question, select all the possible correct options, and each question has at least one correct option.
Premise: On Wednesday, they got married.
Hypothesis: Before Friday, they got married.
Options: A. Entailment B. Contradiction C. Neutral
Answer: Wednesday is before Friday. As a result, we can infer that if something happens on Wednesday, it definitely happens before Friday. Therefore, the answer is A. Entailment.
Premise: We went to Disneyland on Monday.
Hypothesis: We went to Disneyland after Wednesday.
Options: A. Entailment B. Contradiction C. Neutral
Answer: Monday is before Wednesday. As a result, We can infer that if something happens on Monday, it definitely can not happen after Wednesday. Therefore, the answer is B. Contradiction.
Premise: The failing company issued major layoffs after Tuesday.
Hypothesis: The failing company issued major layoffs after Thursday.
Options: A. Entailment B. Contradiction C. Neutral
Answer: Tuesday is before Thursday. If something happened after Tuesday, we cannot be certain whether it occurred after Thursday. Therefore, the answer is C. Neutral.
Figure 7: Chain-of-Thought demonstrations of TimeX-NLI (s1-order).
CoT Demonstration of Date Arithmetic (4-shot)
Question: What is the time 4 year and 1 month after Apr, 2000?
Answer: First, 4 years after 2000 is 2004. Next, 1 month after April is May. Therefore, 4 year and 1 month after Apr, 2000 is May, 2004.
Question: What is the time 3 year and 4 month before Jun, 1840?
Answer: First, subtracting 3 years from 1840 gives 1837. Next, subtracting 4 months from June gives February. Therefore, 3 year and 4 month before Jun, 1840 is Feb, 1837.
Question: What is the time 7 year and 11 month after Feb, 1819?
Answer: First, 7 years after 1819 is 1826. Next, 11 months after February is January of the next year. Therefore, 7 years and 11 months after Feb, 1819 is Jan, 1827.
Question: What is the time 6 year and 9 month before Jan, 1234?
Answer: First, subtracting 6 years from 1234 gives 1228. Next, subtracting 9 months from January gives April of the previous year. Therefore, 6 year and 9 month before Jan, 1234 is Apr, 1227.
Figure 8: Chain-of-Thought demonstrations of Date Arithmetic.
CoT Demonstration of TRACIE (4-shot)
Read the following story and hypothesis, determine whether the hypothesis can be inferred from the story.
You need to understand the implicit temporal relationships between events to make judgments
Story: Joe was a police officer. Joe was patrolling the streets of the city in his cruiser. S̈uddenly, Joe was alerted to a crime happening near him by dispatch.J̈oe responded to the scene and found a bank robber fleeing on foot. Joe arrested the criminal and was promoted.
Hypothesis: Joe put on his police uniform. starts after Joe arrest the criminal
Options: A. Entailment B. Contradiction
Answer: From the story we know Joe was patrolling. In the work state, Joe has already put on the police uniform. So we can infer that Joe put on his police uniform before arresting the criminal. This conflicts with hypothesis. Therefore, the answer is B. Contradiction.
Figure 9: Chain-of-Thought demonstrations of TRACIE.
CoT Demonstration of DurationQA (4-shot)
Answer the following question, select all the possible correct options, and each question has at least one correct option.
Context: actually i have an project on it so please give me as much as you have information about migratory birds in punjab
Question: How long did it take for them to have information about migratory birds in punjab?
Options: A. several months B. 12 weeks C. a few minutes D. almost instantly
Answer: This is a conversation scenario. In the conversation, providing relevant information about migratory birds in punjab to him is in real-time and takes very little time. Therefore, the answer is C. a few minutes, D. almost instantly.
Context: Hope she stops laying eggs because she will get really skinny !
Question: How long did it take for her to lay eggs?
Options: A. 1 week B. 22 hours C. 2 years D. 4 years
Answer: According to commonsense knowledge, the time it takes for birds to lay eggs typically varies from one day to several days. Therefore, the answer is A. 1 week, B. 22 hours.
Figure 10: Chain-of-Thought demonstrations of DurationQA.
CoT Demonstration of MCTACO (4-shot)
Answer the following question, select all the possible correct options, and each question has at least one correct option.
Context: She ordered the tastiest kind of each vegetable and the prettiest kind of each flower.
Question: How often does she order vegetables and flowers?
Options: A. once a second B. three days a week C. every 10 centuries D. once a week
Answer: According to commonsense knowledge, ordering vegetables and flowers typically happens on a regular basis, usually every few days. Therefore, the answer is B. three days a week, D. once a week.
Context: Wallace, 38, called Gastonia home from the age of 8 until she graduated from Hunter Huss High School in 1983.
Question: When did Wallace wake up for high school?
Options: A. at 6 am B. at 1 am C. 7:00 AM D. at 6 pm
Answer: According to commonsense knowledge, waking up for high school typically happens in the morning, usually between 6 AM and 8 AM. Therefore, the answer is A. at 6 am, C. 7:00 AM.
Figure 11: Chain-of-Thought demonstrations of MCTACO.
CoT Demonstration of TimeDial (4-shot)
There is a two-person dialogue with several options.
Choose all appropriate options to substitute the <mask> in the dialogue, and each question has at least one correct option.
Dialogue:
A:What schools have you attended ?
B: I finished Young Primary School in 1998 , and entered Xi ’ an Middle School that same September . I graduated from there in <MASK> , and that September I entered Wuhan University , where I’m studying now .
A: How do you think the education you have received will contribute to your work in this company ?
B: I think I have a good understanding of fundamentals in the areas your company deals with , and I can go on from here to build up the specific skills and knowledge I need to do my job well .
A: Your graduation thesis was on Medical Application of Laser , right ? What were your conclusions ?
B: Yes . I did some work on that , and I found out some really interesting things about the conductivity of liquid helium . I was sure I had a great discovery until my teacher told me the same discovery already made twenty years ago . I think the most important thing , I learnt though , was the importance of keeping good records .
Options: A. 1998 B. July of 2004 C. March of 2003 D. twenty years ago
Answer: Based on the dialogue, B entered middle school in Sep 1998. According to commonsense knowledge, it usually takes around 6 years from entering middle school to graduating from high school (and entering university). Adding 6 years to 1998 would be 2004, so the answer should be around the year 2004. Therefore, the answer is B. July of 2004, C. March of 2003.
Figure 12: Chain-of-Thought demonstrations of TimeDial.
CoT Demonstration of TimeQA, MenatQA (2-shot, implicit)
I will give you a question with context.
You need to answer my question based on the context.
If you can infer the answer from the context, then output your answer. Otherwise, if there is no answer, output [unanswerable]
Context: Theo-Ben Gurirab Theo-Ben Gurirab ( 23 January 1938 2̆013 14 July 2018 ) was a Namibian politician who served in various senior government positions . He served as the second Prime Minister of Namibia from 28 August 2002 to 20 March 2005 , following the demotion and subsequent resignation of Hage Geingob . Previously he was the countrys first Minister of Foreign Affairs from 1990 to 2002 , and was President of the United Nations General Assembly from 1999 to 2000 . He was Speaker of the National Assembly of Namibia from 2005 to 2015 , when he was replaced by Peter Katjavivi . Gurirab ultimately resigned from politics in 2015 . Death . Gurirab died at a Windhoek hospital on 14 July 2018 of natural causes . He is buried at Heroes Acre .
Question: Theo-Ben Gurirab took which position after Jan 2007?
Answer: Based on the context, we can summarize the following facts: Theo-Ben Gurirab served as second Prime Minister of Namibia from August 2002 to March 2005. Prior to that, he was the countrys first Minister of Foreign Affairs from 1990 to 2002 and and was President of the United Nations General Assembly from 1999 to 2000. From 2005 to 2015, he held the position of Speaker of the National Assembly of Namibia. He resigned from politics in 2015 and passed away in July 2018. According to the aforementioned facts, he took the position of Speaker of the National Assembly of Namibia in January 2007. Therefore, the answer is Speaker of the National Assembly of Namibia.
Figure 13: Chain-of-Thought demonstrations of TimeQA, MenatQA, implicit reasoning.
CoT Demonstration of TempReason (4-shot, event-time)
I will give you a question with context.
You need to answer my question based on the context.
Context (facts): Gian Piero Gasperini is the head coach of Atalanta B.C. from Jun, 2016 to Dec, 2022.
Edoardo Reja is the head coach of Atalanta B.C. from Mar, 2015 to Jun, 2016.
Stefano Colantuono is the head coach of Atalanta B.C. from Jun, 2010 to Mar, 2015.
Bortolo Mutti is the head coach of Atalanta B.C. from Jan, 2010 to Jun, 2010.
Emiliano Mondonico is the head coach of Atalanta B.C. from Jul, 1987 to Jun, 1990.
Marcello Lippi is the head coach of Atalanta B.C. from Jul, 1992 to Jun, 1993.
Angelo Gregucci is the head coach of Atalanta B.C. from Jul, 2009 to Sep, 2009.
Luigi Delneri is the head coach of Atalanta B.C. from Jul, 2007 to Jun, 2009.
Ottavio Bianchi is the head coach of Atalanta B.C. from Jul, 1981 to Jun, 1983.
Antonio Conte is the head coach of Atalanta B.C. from Sep, 2009 to Jan, 2010.
Nedo Sonetti is the head coach of Atalanta B.C. from Jul, 1983 to Jun, 1987.
Valter Bonacina is the head coach of Atalanta B.C. from Jan, 2010 to Jan, 2010.
Question: Who was the head coach of the team Atalanta B.C. in Feb, 2016?
Answer: According to the context, Edoardo Reja was the head coach of Atalanta B.C. from Mar, 2015 to Jun, 2016. In Feb 2016, the head coach of the team Atalanta B.C. is Edoardo Reja. Therefore, the answer is Edoardo Reja.
Figure 14: Chain-of-Thought demonstrations of TempReason, event-time reasoning.
CoT Demonstration of TempReason (4-shot, event-event)
I will give you a question with context.
You need to answer my question based on the context.
Context (facts): Nicholas Macpherson holds the position of Member of the House of Lords from Oct, 2016 to Dec, 2022.
Nicholas Macpherson holds the position of Principal Private Secretary to the Chancellor of the Exchequer from Jan, 1993 to Jan, 1997.
Nicholas Macpherson holds the position of Permanent Secretary to the Treasury from Aug, 2005 to Jan, 2016.
Question: Which position did Nicholas Macpherson hold before Member of the House of Lords?
Answer: According to the context, Nicholas Macpherson holds the position of Permanent Secretary to the Treasury from Aug, 2005 to Jan, 2016. Afterthat, Nicholas Macpherson holds the position of Member of the House of Lords from Oct, 2016 to Dec, 2022. Nicholas Macpherson hold the position of Permanent Secretary to the Treasury before Member of the House of Lords. Therefore, the answer is Permanent Secretary to the Treasury."
Figure 15: Chain-of-Thought demonstrations of TempReason, event-event reasoning.

### v2

@@
 Abstract
-Understanding time is a pivotal aspect of human cognition, crucial in the broader framework of grasping the intricacies of the world.
+Grasping the concept of time is a fundamental facet of human cognition, indispensable for truly comprehending the intricacies of the world.
 Previous studies typically focus on specific aspects of time, lacking a comprehensive temporal reasoning benchmark.
-To address this issue, we propose TimeBench, a comprehensive hierarchical temporal reasoning benchmark that covers a broad spectrum of temporal reasoning phenomena, which provides a thorough evaluation for investigating the temporal reasoning capabilities of large language models.
-We conduct extensive experiments on popular LLMs, such as GPT-4, LLaMA2, and Mistral, incorporating chain-of-thought prompting.
+To address this, we propose TimeBench, a comprehensive hierarchical temporal reasoning benchmark that covers a broad spectrum of temporal reasoning phenomena.
+TimeBench provides a thorough evaluation for investigating the temporal reasoning capabilities of large language models.
+We conduct extensive experiments on GPT-4, LLaMA2, and other popular LLMs under various settings.
 Our experimental results indicate a significant performance gap between the state-of-the-art LLMs and humans, highlighting that there is still a considerable distance to cover in temporal reasoning.
-We aspire for TimeBench to serve as a comprehensive benchmark, fostering research in temporal reasoning for LLMs.
-Our resource is available at https://github.com/zchuz/TimeBench
+Besides, LLMs exhibit capability discrepancies across different reasoning categories.
+Furthermore, we thoroughly analyze the impact of multiple aspects on temporal reasoning and emphasize the associated challenges.
+We aspire for TimeBench to serve as a comprehensive benchmark, fostering research in temporal reasoning111 Data is available at: GitHub.
+[…]
+[…]
+TimeBench: A Comprehensive Evaluation of Temporal Reasoning Abilities in Large Language Models
+Zheng Chu1, Jingchang Chen1, Qianglong Chen2, Weijiang Yu3, Haotian Wang1, Ming Liu1,4thanks:  Corresponding Author., Bing Qin1,4 1Harbin Institute of Technology, Harbin, China 2Zhejiang University   3Sun Yat-sen University   4Peng Cheng Laboratory {zchu,jcchen,mliu,qinb}@ir.hit.edu.cn {chenqianglong.ai, wanght1998, weijiangyu8}@gmail.com
 1 Introduction
 Time flies over us, but leaves its shadow behind.
-Time continually moves forward, threading through the fabric of people’s lives.
 Understanding time is a crucial part of human comprehension of the world.
-Imagine witnessing the blossoming of flowers, and you’ll connect it to the arrival of spring.
+Envision the blossoming of flowers, and you’ll associate it with the arrival of spring.
 The ponder within it encompasses the intricate interplay of world knowledge, causality, and event temporal relationships.
@@
 It is a form of integrated reasoning built upon foundational reasoning like mathematical and logical reasoning (Cobbe et al., 2021; Mishra et al., 2022; Yu et al., 2020).
-Recently, large language models (LLM) have demonstrated remarkable performance in complex reasoning (Hendrycks et al., 2021; Srivastava et al., 2022; Brown et al., 2020; Chowdhery et al., 2023; OpenAI, 2023; Touvron et al., 2023), but their performance in intricate temporal reasoning still lacks a definitive conclusion.
-The research community currently lacks a comprehensive benchmark, which makes it challenging to quantify the temporal reasoning capabilities of LLMs.
+Recently, large language models (LLMs) have demonstrated remarkable performance in complex reasoning (Hendrycks et al., 2021; Srivastava et al., 2022; Brown et al., 2020; Chowdhery et al., 2023; OpenAI, 2023; Touvron et al., 2023), but their performance in temporal reasoning has not yet been extensively explored.
 Figure 1: A brief overview of human and LLMs’ performance on TimeBench. Human scores are annotated.
-Recent benchmarks for temporal reasoning typically focus only on a few aspects, such as temporal commonsense or temporal question answering (Zhou et al., 2019; Chen et al., 2021; Dhingra et al., 2022; Wang and Zhao, 2023).
+Recent research for temporal reasoning typically focuses only on a few aspects, such as temporal commonsense or temporal question answering (Zhou et al., 2019; Chen et al., 2021; Dhingra et al., 2022; Wang and Zhao, 2023).
 Due to the inherent complexity of temporal reasoning, it is challenging to accurately measure models’ temporal reasoning capabilities based on limited aspects.
-To address this issue, we propose TimeBench, a comprehensive and hierarchical temporal reasoning benchmark, which is aligned with intrinsic real-world scenarios.
+To address this issue, we propose TimeBench, a comprehensive and hierarchical temporal reasoning benchmark.
 Specifically, drawing inspiration from the human cognitive process of transitioning from abstraction and concreteness to integration (Barsalou et al., 2018), we categorize temporal reasoning into three levels: symbolic temporal reasoning, commonsense temporal reasoning, and event temporal reasoning.
 These levels respectively represent understanding abstract time expression, grasping concrete world knowledge, and integrating and applying this knowledge in real-world scenarios.
-It comprises 10 tasks with 16 sub-tasks, covering a broad spectrum of temporal reasoning phenomena.
-Besides, prior work typically featured only a singular task form, overly simplistic and insufficient in capturing the model’s performance in real-world scenarios.
-In contrast to them, we incorporate four distinct task forms, offering a more realistic simulation of challenges models encounter in real-world scenarios.
-To quantify the temporal reasoning capabilities of contemporary LLMs, we extensively assess widely-used LLMs, including closed-source models such as ChatGPT (Ouyang et al., 2022) and GPT-4 (OpenAI, 2023), as well as open-source like LLaMA2 (Touvron et al., 2023), Vicuna-1.5 (Chiang et al., 2023), Mistral (Jiang et al., 2023), Baichuan2 (Yang et al., 2023), ChatGLM3 Zeng et al.
-(2023) and FLAN-T5 (Shen et al., 2023).
-We conduct experiments under zero-shot and few-shot settings, combining commonly used reasoning techniques, chain-of-thought prompting (Kojima et al., 2022; Wei et al., 2022; Chu et al., 2023a).
-The experimental results suggest that GPT-4 continues to outperform other models, showcasing strong temporal reasoning capabilities, as shown in Figure 1.
-Nevertheless, there is still a considerable gap with humans.
-On the contrary, open-source models generally exhibit subpar performance, lagging significantly behind GPT-3.5 and showing notable shortcomings in commonsense reasoning.
+TimeBench comprises 10 tasks with 16 sub-tasks, covering a broad spectrum of temporal reasoning phenomena.
+Besides, prior work typically features only a single task form, too simplistic to capture the model’s performance.
+In contrast, we incorporate four distinct task forms, offering a more realistic simulation of challenges.
+To quantify the temporal reasoning capabilities of contemporary LLMs, we extensively assess widely-used LLMs, including proprietary models such as ChatGPT (Ouyang et al., 2022) and GPT-4 (OpenAI, 2023), as well as open-source like LLaMA2 (Touvron et al., 2023), Vicuna-1.5 (Chiang et al., 2023), Mistral (Jiang et al., 2023), Baichuan2 (Yang et al., 2023a), ChatGLM3 Zeng et al.
+(2023) and FLAN-T5 (Chung et al., 2022).
+We conduct experiments under zero-shot and few-shot settings, combining commonly used reasoning techniques, chain-of-thought prompting (Kojima et al., 2022; Wei et al., 2022).
+The experimental results suggest that GPT-4 outperforms other models, showcasing strong temporal reasoning capabilities, as shown in Figure 1.
+Nevertheless, there is still a considerable gap between the strongest models and humans.
+On the contrary, open-source models show inferior performance in temporal reasoning, attributed to shortcomings in abstract time understanding, temporal relations modeling, and a lack of temporal commonsense.
+In addition, we also observe that chain-of-thought prompting does not yield a consistent improvement in performance.
 These findings indicate that there is still significant room for improvement in models’ temporal reasoning capabilities.
-In addition, we also observe that, unlike tasks such as mathematical or logical reasoning, chain-of-thought prompting does not consistently enhance model performance; in some cases, it might even impair performance.
-We summarize our contributions as follows: (1) We introduce TimeBench, a comprehensive and hierarchical temporal reasoning benchmark.
-(2) We conduct extensive experiments, revealing a significant gap between even sota LLM and humans, indicating substantial research opportunities in this field.
-(3) By conducting a thorough analysis, we uncover the dilemmas that models face in temporal reasoning and identify potential solutions.
-Figure 2: Distribution of category, task and subtask in TimeBench. Degree of arc indicates the ratio of datas.
+Moreover, we have conducted a thorough analysis of the deficiencies and obstacles faced by models in temporal reasoning.
+We aspire for temporal reasoning to garner increased attention within the research community. Our contributions can be summarized as follows:
+  • We introduce TimeBench, a comprehensive and hierarchical benchmark to quantify the temporal reasoning abilities of LLMs.
+  • We conduct extensive experiments with several LLMs, revealing a significant gap between even SOTA LLM and humans, indicating substantial research opportunities in this field.
+  • By conducting a thorough analysis, we reveal the dilemmas that LLMs face in temporal reasoning and identify potential solutions.
 2 TimeBench Benchmark
@@
 TimeBench focuses on a comprehensive evaluation of the temporal reasoning capabilities of large language models in challenging and complex scenarios.
-To achieve this goal, we summarize the difficulties and challenges faced in temporal reasoning, categorize it into three levels, and meticulously design tasks that better align with complex real-world scenarios.
+To achieve this goal, we summarize the difficulties and challenges faced in temporal reasoning, categorize them into three levels, and integrate diverse task formats to better align with the intricate nature of temporal reasoning.
 Just as the human cognitive process unfolds from foundational cognition and conceptual understanding to practical reasoning, we delineate temporal reasoning into three hierarchical levels.
 Specifically, TimeBench categorizes temporal reasoning into symbolic, commonsense and event temporal reasoning, covering 10 datasets with a total of 16 subtasks.
-(1) Symbolic Temporal Reasoning focuses on comprehension of fundamental abstract temporal expressions.
+(1) Symbolic Temporal Reasoning focuses on the comprehension of fundamental abstract temporal expressions.
 (2) Temporal Commonsense Reasoning emphasizes the mastery of temporal principles, concepts and world knowledge.
-(3) Event Temporal Reasoning concentrates on modeling the temporal relationships of events within authentic scenarios.
-Furthermore, to better align with real-world scenarios, we employ diverse question formats.
+(3) Event Temporal Reasoning concentrates on modeling the temporal relationships between events and times within authentic scenarios.
 2.2 Difficulties and Challenges
-We outline the necessary capabilities and challenges encountered from a human cognition perspective during temporal reasoning, and language models confront similar obstacles.
-Detailed information on TimeBench and challenges involved in each task is presented in Table 6.
-TimeX Understanding
-Time expressions (TimeX) denote words or phrases that convey information about time and represent the simplest and most basic units of expressing time, such as in August 2008, which is fundamental in comprehending time.
+We delineate the essential competencies and the challenges that arise from a human cognitive standpoint in the realm of temporal reasoning, and language models confront similar challenges.
+We present the dataset statistics, task formats, and the associated challenges in Table 7.
+Time Expression Understanding
+Time expressions (TimeX) denote words or phrases that convey information about time and represent the simplest and most basic units of expressing time, such as in April 2000, after 2008.
+Grasping time expressions is the most foundational step in understanding temporal elements within the textual modality.
 Temporal Commonsense
-assesses the understanding of temporal aspects of world knowledge, including event order, event duration, event typical time, and event frequency, which is crucial for language models to comprehend everyday scenarios.
+assesses the understanding of temporal world knowledge, including event order, event duration, typical time, event frequency and stationary, which is crucial for language models to comprehend daily scenarios.
 Event-Time Relations
-evaluates the grounding ability to establish correspondence between events and time, enabling the model to comprehend the development and changes of events as they unfold over time.
+assesses the model’s grounding capability to establish temporal relationships between events and their temporal context, thereby enabling models to grasp the progression and transformations of events as they dynamically evolve through time.
 Event-Event Relations
-not only involve event-time grounding but also introduce multi-hop relative connections.
+not only involve event-time grounding but also introduce multi-hop relative connections between events.
 Models with this capability can better handle temporal reasoning in complex scenarios involving multiple events.
@@
 Implicit temporal reasoning is pivotal in complex real-world scenarios where events and time are intricately interwoven.
-Date Arith
 Q: What is the time 2 year and 4 month before Mar, 1755
 A: Nov, 1752
-TimeX NLI
+Date Arith
 Premise: On 28th May 1967, I graduated.
@@
 A: Contradiction
+TimeX NLI
 Table 1: Examples of symbolic temporal reasoning
-2.3 Symbolic Temporal Reasoning
-To evaluate the language model’s comprehension of abstract time expressions, we utilize two symbolic reasoning tasks stripped of semantic content: TimeX arithmetic and TimeX inference.
-Table 1 shows examples of symbolic temporal reasoning.
-TimeX Arithmetic (Tan et al., 2023)
-assesses the model’s grasp of abstract date calculation. When provided with a date, the model needs to accurately calculate the date a certain amount of time before or after the given date.
-TimeX NLI (Thukral et al., 2021)
-focuses on the logical entailment relationships among abstract TimeX, including three aspects: order (s1), duration (s2), and duration with unit conversion (s3).
-MCTACO
 C: Ransome looks after her as well as for young Fern Simon , who has declared her love for him.
@@
 O: each century, once a day, once a century, every night
-TimeDial
+MCTACO
 Dialog: … Person1: Do you go to work by train every day  Person2: Yes . I commute <MASK> a week by train…
 O: five days, 25 days, a minute, six days
-SituatedGen
+TimeDial
 Keywords: axis, one day, one month, Earth, Moon
 A: Earth rotates on its axis once in one day. It takes one month for the Moon to rotate on its axis.
+SituatedGen
 Table 2: Examples of commonsense temporal reasoning.
-2.4 Commonsense Temporal Reasoning
-We measure the model’s mastery of temporal common and world knowledge, along with its capacity for reasoning based on these insights.
-Table 2 presents examples of temporal commonsense reasoning in QA and generation forms.
-MCTACO (Zhou et al., 2019)
-evaluates diverse commonsense knowledge from different aspects of events, including duration, frequency, order, stationary and typical event time.
-DurationQA (Virgo et al., 2022)
-focuses specifically on temporal commonsense reasoning in the spectrum of event duration.
-TimeDial (Qin et al., 2021)
-considers temporal commonsense reasoning in dialogue scenarios and involves various aspects of commonsense associated with duration, order, and world knowledge.
-SituatedGen (Zhang and Wan, 2023)
-considers generative commonsense reasoning in a constrained text generation scenario.
-Given a set of contrasting keywords, the model needs to choose appropriate keywords for each sentence and generate a pair of contrasting sentences that satisfy temporal commonsense.
-TimeQA
-C: … He worked in Utrecht for the firm of P Smits & de Wolf from 1864 to 1867 and then returned to Widnes …
+2.3 Symbolic Temporal Reasoning
+To evaluate the language model’s comprehension of abstract time expressions, we utilize two symbolic reasoning tasks stripped of semantic content: date arithmetic and time expression inference.
+Table 1 shows examples of symbolic temporal reasoning.
+Date Arithmetic
+(Tan et al., 2023) assesses the model’s grasp of abstract date calculation.
+When provided with a date, the model needs to accurately calculate the date a certain amount of time before or after the given date.
+The smallest unit is one day.
+TimeX NLI
+(Thukral et al., 2021) focuses on the logical entailment relationships among abstract TimeX, including three aspects: order (s1), duration (s2), and duration with unit conversion (s3).
+C: … He worked in Utrecht for the firm of P Smits & de Wolf from 1864 to 1867 and then returned to …
 Q: Where did Ludwig Mond work between Mar 1866 and Sep 1866?
 A: Utrecht
-MenatQA
+TimeQA
 C: … After the French evacuated Egypt in 1801, Hurshid Pasha was named governor of Egypt in 1804. Muhammad Ali had himself named governor of Egypt in May 1805 …
@@
 A: governor of Egypt
-TempReason
+MenatQA
 C: … Peter Corke works for Queensland University of Technology from Jan, 2010 to Dec, 2022. Peter Corke works for Commonwealth Scientific from Jan, 1984 to Jan, 2009. …
@@
 A: Commonwealth Scientific
-Table 3: Examples of event temporal reasoning. C=context, Q=question, A=answer
+TempReason
+Table 3: Examples of event temporal reasoning.
+2.4 Commonsense Temporal Reasoning
+We measure the model’s mastery of temporal commonsense and world knowledge, along with its capacity for reasoning based on these insights.
+Table 2 presents examples of temporal commonsense reasoning in QA and generation forms.
+MCTACO
+(Zhou et al., 2019) evaluates diverse commonsense knowledge from different aspects of events, including duration, frequency, order, stationary and typical event time.
+DurationQA
+(Virgo et al., 2022) focuses specifically on temporal commonsense reasoning in the spectrum of event duration.
+TimeDial
+(Qin et al., 2021) considers temporal commonsense reasoning in dialogue scenarios and involves various aspects of commonsense associated with duration, order, and world knowledge.
+SituatedGen
+(Zhang and Wan, 2023) considers generative commonsense reasoning in a constrained text generation scenario.
+Given a set of contrasting keywords, the model needs to choose appropriate keywords for each sentence and generate a pair of contrasting sentences that satisfy temporal commonsense.
 2.5 Event Temporal Reasoning
@@
 Examples are shown in Table 3.
-TimeQA (Chen et al., 2021)
-requires the model to answer time-sensitive questions based on context containing numerous time-involved facts.
+TimeQA
+(Chen et al., 2021) requires the model to answer time-sensitive questions based on context containing numerous time-involved facts.
 It is categorized into explicit reasoning and implicit reasoning based on time indicators (before, in, etc.).
-MenatQA (Wei et al., 2023)
-introduces temporal factors to elicit implicit temporal reasoning, including time scope change, disruption of facts, and counterfactual questions, which provides a more in-depth assessment of implicit reasoning ability on event-time relations.
-TempReason (Tan et al., 2023)
-removes irrelevant context and focuses on implicit temporal reasoning within structured facts, investigating the model’s capability boundaries.
+MenatQA
+(Wei et al., 2023) introduces time-sensitive factors to elicit implicit temporal reasoning, including time scope change, disruption of facts, and counterfactual questions, which provides a more in-depth assessment of implicit reasoning ability on event-time relations.
+TempReason
+(Tan et al., 2023) removes irrelevant context and focuses on implicit temporal reasoning within structured facts, investigating the model’s capability boundaries.
 It involves event-time reasoning and event-event reasoning.
-TRACIE Zhou et al. (2021)
-evaluates the model’s comprehension of temporal order between implicit events. The model needs to identify events implied in the context and then determine their chronological order.
-2.6 Task Format and Evaluation Metrics
-Real-world scenarios are diverse in task forms. To investigate the model’s authentic performance, we select diverse task forms, covering a wide spectrum of temporal reasoning phenomena.
-Multi-choice & Multi-answer Questions
-MCQ tasks require models to select the only correct answer from the options.
-However, this task form has shortcuts and may not truly reflect the model’s capability.
-To address this, we use the MCMAQ form, where the model needs to select all possible correct answers from options.
-In our task, each question includes four options, with two of them being correct.
-Task: MCTACO, DurationQA, TimeDial.
+TRACIE
+(Zhou et al., 2021) evaluates the model’s comprehension of temporal order between implicit events.
+The model needs to identify events implied in the context and then determine their chronological order.
+2.6 Task Formats and Evaluation Metrics
+TimeBench is a multispectral benchmark encompassing four task types: free-form reading comprehension, natural language inference, constrained text generation, and multi-select questions.
+For detailed task types and their corresponding evaluation metrics, please refer to Appendix A.3 and A.4.
+3 Methodology
+We perform evaluations using the prompt-based approach, including standard prompting and chain-of-thought prompting. Experiments are conducted under both zero-shot and few-shot settings.
+Standard Prompting
+We formulate specific instructions for each task.
+In the zero-shot setting, models follow the instructions to answer questions.
+In the few-shot setting, models are provided with several question-answer pairs as demonstrations and emulate those instances to answer questions.
+$\mathrm{prompt_{zs}^{sp}}=\{\mathrm{INST}\}\{\mathrm{Q}\}$ (1)
+$\mathrm{prompt_{fs}^{sp}}=\{\mathrm{INST}\}\{\mathrm{Q_{1}}\}\{\mathrm{A_{1}}\}..\{\mathrm{Q}\}$ (2)
+Chain-of-Thought Prompting
+The instructions of CoT are the same as standard prompting.
+In the zero-shot setting, following Zeroshot CoT (Kojima et al., 2022), we add a reasoning trigger Let’s think step by step after questions to perform chain-of-thought reasoning.
+In the few-shot setting, we manually annotate CoT demonstrations for each task to guide the step-by-step reasoning.
+Prompts can be found in Appendix B.3.
+$\mathrm{prompt_{zs}^{cot}}=\{\mathrm{INST}\}\{{\mathrm{Q}}\}\{{\mathrm{TRIG}}\}$ (3)
+$\mathrm{prompt_{fs}^{cot}}=\{\mathrm{INST}\}\{{\mathrm{Q_{1}}}\}\{{\mathrm{R_{1}}}\}\{{\mathrm{A_{1}}}\}..\{{\mathrm{Q}}\}$ (4)
+4 Experimental Setup
+Method Symbolic Commonsense Event Temporal Overall
+TimeXNLI Arith DQA McT. TiD. SitGen TimeQA MenatQA TempR TRACIE Sym. Comm. Event Avg.
+s1 s2 s3 Exp. Imp. Sco. Ord. Ctf. L2 L3
+Human 98.0 96.0 92.0 100.0 80.8 87.1 97.8 100.0 93.3 91.1 85.6 87.3 79.9 97.1 95.3 82.5 96.5 91.4 89.0 91.5
+GPT-4 85.3 73.3 53.3 100.0 64.8 88.3 94.6 88.6 73.7 51.0 72.4 54.8 28.7 92.4 95.9 62.8 78.0 84.1 66.5 73.7
++ FS CoT 92.0 84.0 64.0 100.0 55.1 72.3 93.4 - 66.9 52.8 65.3 52.6 25.9 96.9 94.6 66.4 85.0 73.6 65.2 72.1
+GPT-3.5 52.0 68.4 31.6 63.6 67.7 71.2 76.4 79.1 66.1 48.4 43.2 51.6 17.9 84.7 78.0 55.0 53.9 73.6 55.6 59.7
++ FS CoT 51.6 71.8 36.6 84.4 41.2 38.1 71.1 - 68.0 47.0 42.5 41.7 37.8 89.9 76.6 50.2 61.1 50.1 56.7 56.6
+[…]
++ FS CoT 52.0 73.0 39.0 79.5 62.3 79.1 61.1 - 64.3 43.0 57.7 45.2 53.1 87.5 81.6 67.0 60.9 67.5 62.4 63.0
+[…]
++ FS CoT 40.0 61.0 37.0 52.0 59.3 68.8 40.8 - 59.4 49.1 58.4 43.8 44.1 78.0 68.2 58.0 47.5 56.3 57.4 54.5
+[…]
++ FS CoT 37.0 52.0 36.0 25.5 56.9 67.0 41.9 - 45.6 36.1 50.9 38.0 57.3 59.7 57.7 50.0 37.6 55.3 49.4 47.4
+[…]
++ FS CoT 50.0 56.0 34.0 47.0 62.0 69.3 43.8 - 58.2 49.6 49.8 40.1 45.6 81.3 65.6 60.0 46.8 58.4 56.3 54.2
+[…]
++ FS CoT 30.0 56.0 34.0 34.0 57.0 69.5 44.5 - 51.2 40.7 46.4 32.6 46.3 61.5 64.1 53.0 38.5 57.0 49.5 48.1
+[…]
++ FS CoT 57.0 63.0 35.0 54.0 61.8 45.7 57.3 - 60.4 46.2 57.2 47.9 33.2 65.9 67.9 57.0 52.3 54.9 54.5 54.0
+[…]
++ FS CoT 47.0 68.0 32.0 46.0 53.9 64.3 56.5 - 52.5 24.5 35.0 40.2 22.5 79.4 60.3 54.0 48.3 58.2 46.1 49.1
+Table 4: Experimental results under few-shot settings (standard prompting by default).
+denotes the base model without alignment.
+Global top-3 results are bold.
+Figure 8 provides a horizontal comparison of the performance of all models.
+Full results in Appendix B.2.
+Method Symbolic Commonsense Event Temporal Overall
+TimeXNLI Arith DQA McT. TiD. SitGen TimeQA MenatQA TempR TRACIE Sym. Comm. Event Avg.
+s1 s2 s3 Exp. Imp. Sco. Ord. Ctf. L2 L3
+Human 98.0 96.0 92.0 100.0 80.8 87.1 97.8 100.0 93.3 91.1 85.6 87.3 79.9 97.1 95.3 82.5 96.5 91.4 89.0 91.5
+GPT-4 78.6 76.0 50.7 98.0 59.2 80.0 91.1 59.3 60.6 46.5 57.0 57.0 23.1 95.3 95.0 64.8 75.8 72.4 62.4 68.3
++ CoT 80.0 76.0 60.0 92.0 58.1 82.6 89.3 - 61.3 41.2 54.6 59.6 22.6 97.0 94.5 58.0 77.0 76.7 61.1 68.5
+GPT-3.5 45.4 67.6 31.2 97.0 50.5 68.6 69.1 62.3 70.8 35.4 40.9 43.9 22.9 81.2 73.8 57.4 60.3 62.6 53.3 57.4
++ CoT 33.6 64.8 33.6 71.0 23.2 45.1 67.0 - 64.4 35.1 39.7 42.9 26.3 57.6 68.1 52.0 50.8 45.1 48.3 48.3
+LLaMA270b 44.0 47.0 32.0 78.5 59.2 68.9 57.0 25.0 40.8 40.6 18.9 16.6 12.0 63.5 54.5 48.0 50.4 52.5 36.8 44.1
++ CoT 30.0 66.0 28.0 53.5 57.3 67.1 58.6 31.4 19.5 12.2 12.7 20.8 37.5 40.5 51.0 44.4 61.0 28.2 39.1
+LLaMA213b 30.0 49.0 34.0 22.5 38.5 40.6 35.4 57.9 61.9 30.5 46.1 36.1 26.9 53.1 69.4 49.0 33.9 43.1 46.6 42.6
++ CoT 36.0 50.0 38.0 6.0 39.2 51.7 36.9 - 58.7 38.9 40.9 32.5 33.6 58.0 68.4 47.0 32.5 42.6 47.3 42.4
+LLaMA27b 39.0 53.0 30.0 13.0 39.3 41.0 6.3 24.5 49.0 29.0 26.8 21.1 16.0 63.9 47.9 49.0 33.8 27.8 37.8 34.3
++ CoT 44.0 50.0 33.0 5.0 35.0 40.0 1.7 - 49.9 31.6 31.4 24.5 17.8 56.9 48.1 46.0 33.0 25.6 38.3 34.3
+Baichuan213b 41.0 61.0 37.0 12.5 52.0 63.4 57.7 52.2 55.4 34.6 48.8 44.3 39.5 57.4 61.4 49.0 37.9 56.3 48.8 48.0
++ CoT 40.0 57.0 31.0 10.0 44.6 61.9 58.1 - 41.5 40.9 52.0 38.5 43.2 62.8 64.3 55.0 34.5 54.9 49.8 46.7
+Baichuan27b 35.0 50.0 37.0 4.5 47.9 55.3 54.3 42.0 41.5 34.7 35.2 31.2 20.4 43.4 47.7 55.0 31.6 49.9 38.6 39.7
++ CoT 38.0 43.0 32.0 1.0 37.9 58.0 44.2 - 53.5 38.8 39.9 33.2 29.3 41.2 47.2 54.0 28.5 46.7 42.1 39.4
+Vicuna1.513b 35.0 50.0 36.0 15.0 39.2 59.1 34.2 51.8 60.4 37.0 46.8 37.4 23.2 42.1 43.6 46.0 34.0 46.1 42.1 41.1
++ CoT 42.0 51.0 37.0 3.0 29.8 50.0 33.7 - 56.9 36.4 38.2 37.7 20.4 49.0 49.1 51.0 33.3 37.8 42.3 39.0
+Vicuna1.57b 37.0 58.0 43.0 5.0 40.4 52.5 32.0 47.8 47.1 18.5 35.7 25.7 17.3 33.0 46.8 54.0 35.8 43.2 34.8 37.1
++ CoT 36.0 50.0 36.0 1.5 39.4 49.2 36.2 - 40.9 24.6 26.2 28.5 25.0 27.7 40.3 54.0 30.9 41.6 33.4 34.4
+FLANT511b 53.0 63.0 43.0 0.0 52.0 65.0 47.7 49.5 61.7 26.8 33.6 52.2 21.8 87.9 83.9 64.0 39.8 53.6 54.0 50.3
++ CoT 56.0 66.0 45.0 0.0 49.7 63.4 42.7 - 64.4 28.2 41.6 50.2 30.6 79.5 68.9 55.0 41.8 51.9 52.3 49.4
+Mistral7b 47.0 50.0 43.0 26.5 49.8 58.8 23.2 58.3 28.2 21.4 24.3 22.3 21.7 39.6 31.6 51.0 41.6 47.5 30.0 37.3
++ CoT 38.0 56.0 35.0 16.5 36.6 49.3 19.3 - 31.3 22.4 21.1 24.9 25.6 34.0 31.2 61.0 36.4 35.1 31.4 33.5
+ChatGLM36b 38.0 50.0 34.0 2.0 34.1 43.6 56.7 38.9 41.2 31.7 33.8 26.0 32.2 57.0 54.0 50.0 31.0 43.3 40.7 39.0
++ CoT 27.0 49.0 37.0 0.0 24.8 37.1 44.8 - 41.7 25.4 34.6 28.1 41.2 44.5 52.0 48.0 28.3 35.6 39.4 35.7
+Table 5: Experimental results under zero-shot settings (standart prompting by default). All models are alignment models (-chat or -instruct). Global top-3 results are bold.
+4.1 Models
+[…]
+4.2 Implementation Details
+We access proprietary models through Azure API 0613 version.
+For open-source models, we deploy them locally through FastAPI.
+We set the temperature to 0.0 for greedy decoding in all experiments.
+To improve answer extraction accuracy, we prompt models with trigger Therefore, the answer is before model outputs to deduce final answers.
+5 Experimental Results
+5.1 Few-shot Results
+[…]
+In symbolic temporal reasoning tasks, GPT-4 demonstrates exceptional performance.
+However, other models exhibit a significant decline in comparison to GPT-4.
+In commonsense temporal reasoning tasks, GPT4 lags behind humans by only 8.0%, indicating its powerful internal knowledge reservoir.
+With the model scale shrinking, its knowledge reservoir also decreases gradually, leading to a decline in performance.
+Notably, there is a significant gap of 25.2% between LLMs and humans in event temporal reasoning, which suggests that LLMs encounter major challenges in modeling intricate event-time relationships.
+5.2 Zero-shot Results
+Experimental results of alignment models under zero-shot settings are shown in Table 5.
+In zero-shot settings, GPT-4 and GPT-3.5 rank first and second, respectively, and they significantly outperform all open-source models by a large margin.
+It is noteworthy that open-source models exhibit a larger performance decline compared to proprietary models when transitioning from few-shot to zero-shot scenarios.
+GPT, Baichuan2 and LLaMA2 suffer drops of 5.6%, 14.6% and 27.2%, respectively.
+We attribute this performance decline to the quality of alignment.
+Restricted by their limited instruction-following capability, open-source models struggle to fully unleash their performance solely through instructions.
+Therefore, few-shot prompting is a better approach for stimulating their temporal reasoning abilities.
+[…]
+5.3 Chain-of-Thought in Temporal Reasoning
+Previous research has found that chain-of-thought prompting can enhance the model’s reasoning ability (Wei et al., 2022; Kojima et al., 2022).
+We aim to explore the following questions: Does CoT prompting bring consistent improvement in temporal reasoning?
+Due to the diversity of temporal reasoning, the above question has not yet been definitively answered.
+To investigate this, we select several popular LLMs and analyze their performance affected by chain-of-thought prompting.
+Chain-of-thought reasoning is not consistently effective.
+As illustrated in Figure 2, introducing zero-shot CoT prompting results in consistent declines, with an overall decrease of 7.4%.
+In the few-shot scenario, CoT prompting also fails to yield consistent improvements, varying depending on the task.
+There is a 10.8% improvement in symbolic reasoning, while a significant decline of 15.2% in commonsense reasoning.
+In event temporal reasoning, there is a slight improvement of 1.3%.
+Next, we will conduct a more detailed analysis of the impact of CoT on specific tasks.
+Figure 3: $\Delta\text{Score}$ between the chain-of-thought prompting and direct I-O prompting.
+Top: zero-shot setting, Bottom: few-shot setting, Left: variation in each task, Right: averaged variation in the symbolic, commonsense, event, and overall tasks.
+Impact of CoT prompting across tasks.
+In order to explore the impact of CoT on various tasks thoroughly, we delve into the performance changes of each model across specific tasks within each category, as illustrated in Figure 3.
+In the zero-shot setting, open-source models achieve a slight improvement in event temporal reasoning with chain-of-thought prompting, while in other cases, they face performance degradation.
+While in the few-shot setting, almost all models exhibit significant improvement in symbolic temporal reasoning, with a concurrent prevalent decline in commonsense temporal reasoning.
+We attribute this to the knowledge sensitivity inherent in commonsense reasoning, where step-by-step reasoning cannot compensate for the lack of knowledge.
+In event temporal reasoning, improvements mainly stem from datasets involving implicit multi-step reasoning (MenatQA and TempReason), indicating that CoT is more effective for multi-hop questions.
+In summary, zero-shot CoT consistently has a negative impact on temporal reasoning.
+While in few-shot scenario, CoT has a positive impact on symbolic and complex tasks, while negatively affecting knowledge-sensitive tasks.
+6 Analysis and Discussion
+6.1 Scaling Effect of Model Size
+Figure 4: Scaling effect of model size and overall temporal reasoning performance.
+The x-axis (model size) is shown in the log scale.
+Results show a log-linearity between parameter size and performance.
+We investigate how the scale of models affects temporal reasoning capabilities.
+The trend is illustrated in Figure 4.
+As the model scale increases, there is a notable improvement in performance.
+When the parameter size expands from 7B to 13B, LLaMA2 and Baichuan2 show improvements of 13.0% and 10.5%, respectively.
+Furthermore, when LLaMA scales up to 70B, the trend of performance improvement continues without stopping.
+The overall improvement follows a log-linear relationship with scale.
+There are no significant performance differences among LLaMA2, Baichuan2, and ChatGLM3 under similar parameter specifications, while Mistral demonstrates impressive prowess, outperforming all other 13B models with nearly half the number of parameters.
+6.2 Challenges in Temporal Reasoning
+Model Order Duration Freq. Stationarity Typical Avg.
+GPT-4 76.4 $\downarrow$ 92.8 $\uparrow$ 83.3 $\uparrow$ 71.4 $\downarrow$ 54.5 $\downarrow$ 77.5
+GPT-3.5 50.5 $\uparrow$ 39.8 $\downarrow$ 55.2 $\uparrow$ 48.4 $\uparrow$ 28.7 $\downarrow$ 43.5
+[…]
+[…]
+[…]
+Table 6: Results in each temporal commonsense aspect under few-shot setting.
+Models with †  are base models.
+Red $\downarrow$ and Green $\uparrow$ represent the performance is lower or higher than its average performance.
+Metric is EM.
+LLMs underperform in (multi-hop) symbolic reasoning
+[…]
+Mastery of commonsense knowledge varies in LLMs
+We analyze models’ performance across various commonsense aspects, as shown in Table 6.
+We regard the model’s average performance in commonsense reasoning tasks as the baseline.
+If the model outperforms the baseline in a specific aspect, it suggests greater proficiency in this type of knowledge, and vice versa.
+The findings indicate that LLMs generally demonstrate good knowledge of event duration and frequency.
+However, their comprehension of event order and typical events is relatively weaker The uneven mastery of commonsense knowledge significantly affects the model’s reasoning performance, especially when dealing with complex questions that involve multiple types of knowledge.
+Retrieval-augmented reasoning presents a promising avenue for mitigating the model’s knowledge scarcity.
+LLMs exhibit poor implicit temporal reasoning capabilities.
+When comparing explicit and implicit event temporal reasoning, specifically TimeQA-explicit versus others, we observe a significant performance decrease in implicit reasoning.
+Additionally, on TRACIE with numerous implied events, most models only surpass a random baseline (50.0).
+Even GPT-4, despite its advanced capabilities, achieves only a 66.4% accuracy, suggesting that the LLM struggles with modeling implicit temporal relationships.
+We consider it helpful to explicitly model the temporal relationships between events and time expressions, for instance constructing timelines or temporal graphs.
+LLMs are good factual reasoners rather than factual extractors
+When humans engage in temporal reasoning, it generally involves two steps: first, extracting time-fact pairs from the context, and then performing fact-based reasoning.
+TempReason provides extracted facts for conducting fact-based reasoning.
+By comparing the model’s performance in context-based (TimeQA) against fact-based (TempReason) reasoning, we identify the bottleneck in event temporal reasoning.
+LLMs excel in TempReason, which signifies their strong capability in fact-based reasoning.
+However, their performance in context-based reasoning is significantly weaker compared to their performance in fact-based reasoning.
+This implies that errors could arise during the extraction of time-sensitive facts from the context.
+We attribute this performance gap to the model’s deficiency in factual extraction capabilities Thus, we consider LLMs to be strong factual reasoners rather than factual extractors in event temporal reasoning.
+Figure 5: Performance difference between base and alignment models under few-shot setting. Baichuan2 and LLaMA2 are aligned with SFT and RLHF. Vicuna, Mistral and ChatGLM3 are aligned with only SFT.
+6.3 Alignment Impairs Temporal Reasoning
+In the experiments mentioned earlier (Table 5), we observe a sharp decline in zero-shot performance of alignment models.
+To investigate whether alignment is the cause of the decline in temporal reasoning, we conducted experiments on alignment models under few-shot settings.
+Figure 5 illustrates the overall performance decline after alignment.
+With the exception of Baichuan2, all other models are severely impaired, experiencing a significant drop of up to 22%.
+Through manual analysis of error cases, we have summarized two reasons: (1) Alignment reduces the model’s usability, causing it to tend towards refusal to answer when confronted with knowledge-sensitive questions.
+(2) Alignment damages the model’s in-context learning capability, resulting in situations where the model deviates from the demonstrations.
+Furthermore, we believe that the lack of temporal reasoning-related training data in alignment exacerbates this issue, leading to disparities between different reasoning capabilities, such as mathematical and temporal reasoning.
+Figure 6: Error analysis for Symbolic, Commonsense, and Event Temporal. We select 100 test samples from each subtask for GPT-4, GPT-3.5 and LLaMa2-base70b.
+6.4 Error Analysis
+We manually analyze 100 predictions by GPT-4, GPT-3.5 and LLaMa2-base70b from each subtask. The visualization of errors is shown in Figure  6.
+Symbolic Reasoning
+We categorize symbolic reasoning errors into five groups: (a) Expression: The model provides an incorrect time calculation expression.
+(b) Computation: The model provides the correct time calculation expression, but there is a calculation error.
+(c) Conversion: The model has an error in the conversion of time units.
+(d) Comparison: The model has an error when comparing two time-expressions (or intervals).
+(e) Combination: The model encounters errors in the combination of multiple above operations.
+LLMs exhibit numerous computation, conversion, and comparison errors, which suggests a substantial deficiency in their understanding of fundamental temporal expressions.
+Additionally, a higher frequency of errors is observed in combination questions, highlighting that multi-step reasoning continues to be a significant challenge for current models
+Commonsense Reasoning
+We categorize the errors of commonsense reasoning into two groups: (a) No Answer: The model fails to provide a final answer.
+(b) Reasoning Error: The model encounters reasoning errors, which can be subdivided into five types of knowledge-related errors.
+We observe that GPT series models have a higher No Answer rate, while LLaMA is always able to provide answers.
+This discrepancy can be attributed to two factors: firstly, the models may lack the necessary commonsense knowledge to formulate an answer; secondly, the preference alignment mechanism may prompt the model to abstain from answering when confronted with questions outside its knowledge scope.
+Integration of retrieval can alleviate the problem of knowledge scarcity to a certain degree.
+Event Temporal Reasoning
+We categorize the errors of event temporal reasoning into four groups: (a) No Answer: The model is unable to find the answer in the context.
+(b) Reasoning Error: The model encounters reasoning errors.
+(c) Hallucination: The model’s prediction does not exist in the context, known as hallucination reasoning.
+(d) Metric: The model’s prediction is correct, but the metric is limited by the evaluation criteria.
+It can be observed that, except for reasoning errors, failures to provide answers account for approximately 30%, indicating that models still have flaws in grounding temporal facts from context.
+Additionally, models occasionally experience hallucination phenomena, leading to erroneous reasoning.
+7 Related Work
+7.1 Temporal Reasoning
+There are numerous efforts addressing diverse challenges in temporal reasoning.
+Early research mainly relies on TimeML (Pustejovsky et al., 2003), focusing TimeX extraction and temporal relation extraction (Verhagen et al., 2007, 2010; UzZaman et al., 2013; Llorens et al., 2015; Miller et al., 2015; Mathur et al., 2021; Vashishtha et al., 2019).
+The advent of pre-trained language models (PLMs) has brought about commonsense reasoning as a tool to explore the world knowledge in models (Zhou et al., 2019; Qin et al., 2021; Dhingra et al., 2022).
+Recently, much attention has shifted towards event temporal reasoning (Chen et al., 2021; Tan et al., 2023; Wei et al., 2023).
+Han et al.
+(2021); Yang et al.
+(2023b); Son and Oh (2023); Chen et al.
+(2023) continuously pre-trains LLMs on time-aware data to elicit temporal reasoning, and Zhu et al.
+(2023); Su et al.
+(2023); Chu et al.
+(2023) explicitly represent temporal relationships using temporal graphs and timelines.
+Additionally, some works extend beyond text, evaluating temporal reasoning in structured tables and video domains (Gupta et al., 2023; Ko et al., 2023).
+Some concurrent studies also analyze the temporal reasoning abilities of LLMs.
+Jain et al.
+(2023); Qiu et al.
+(2023) focus on temporal commonsense and Wang and Zhao (2023) introduces a unified form for accessing the overall abilities.
+Distinguished from other works, TimeBench is multispectral, offering a comprehensive evaluation of LLM’s temporal reasoning abilities.
+7.2 Large Language Models
+In recent years, there has been rapid progress in the research of large language models (LLM) (Zhao et al., 2023).
+They exhibit outstanding performance across a multitude of tasks without the need for fine-tuning (Brown et al., 2020; Kojima et al., 2022).
+Furthermore, they have achieved astonishing results in complex reasoning tasks, such as mathematical reasoning (Cobbe et al., 2021; Mishra et al., 2022) and logical reasoning (Yu et al., 2020; Liu et al., 2023).
+Moreover, some studies suggest that the chain-of-thought prompting can further enhance the model’s capabilities in complex reasoning scenarios (Wei et al., 2022; Kojima et al., 2022; Chu et al., 2024; Zhang et al., 2023).
+8 Conclusion
+Temporal reasoning entails inherent diversity and complexity.
+The lack of a comprehensive benchmark makes it challenging to quantify LLMs’ temporal reasoning capabilities.
+In this work, we present TimeBench, a comprehensive and hierarchical benchmark for LLM temporal reasoning, tailored to mirror temporal reasoning in complex scenarios.
+We conduct extensive experiments on state-of-the-art LLMs to investigate their temporal reasoning capabilities.
+Our findings indicate a substantial gap between state-of-the-art LLMs and human performance, emphasizing the need for further research in this area.
+Moreover, we provide a meticulous analysis and discussion, outlining the current challenges that models face and suggesting potential directions for improvement.
+Limitations
+TimeBench is a comprehensive benchmark to quantify the temporal reasoning capabilities of LLMs.
+While we have taken various factors into account, there are a few limitations.
+Firstly, our evaluation only applied prompt-based method under zero-shot and few-shot setting, lacking evaluations specifically tailored for models fine-tuned on the temporal domain.
+Secondly, the instructions and demonstrations were manually crafted, which may potentially lead to discrepancies in prompts interpretation among different LLMs.
+Thirdly, the dataset constituting the benchmark includes data from past years and a portion sourced from Wikipedia, which may contaminate the training corpus of LLMs.
+Acknowledgements
+The research in this article is supported by the National Key Research and Development Project (2021YFF0901602), the National Science Foundation of China (U22B2059, 62276083), and Shenzhen Foundational Research Funding (JCYJ20200109113441941), Major Key Project of PCL (PCL2021A06).
+Ming Liu is the corresponding author.
+Appendix A TimeBench Details
+TimeBench features 3 major categories, 10 tasks and 15 subtasks, each with distinct challenges, totaling 19,000 instances. Detailed statistics are available in Figure 7 and Table 7.
+A.1 Benchmark Construction
+TimeX Arithmetic
+(Tan et al., 2023) TimeX Arithmetic data is derived from the l1: time-time reasoning data in TempReason.
+We retain 4,000 instances, where time expressions are calculated with a minimum unit of one day.
+TimeX NLI
+(Thukral et al., 2021) The original data of TimeXNLI is in NLI format, including three sub-tasks, Temp-Order, Temp-Duration, and Cross-Unit Duration, including 6,140, 3,540, and 15,840 instances respectively.
+We conduct a random sampling of 2,213, 2,332 and 2,429 entries, resulting in a combined total of 6,965 instances.
+MCTACO
+(Zhou et al., 2019) The original MCTACO dataset consists of yes/no questions, containing 1,332 questions with 9,442 options.
+To guarantee that the questions are presented in a 4-way multi-select style, we initially remove questions that have less than four options.
+Subsequently, to ensure that each question has at least one correct option, we filter out questions where all options are labeled as "no".
+For each remaining question, we randomly sample four options, striving to maintain a balance between correct and incorrect options.
+In most cases, a question is accompanied by 2 correct and 2 incorrect options.
+A minority of questions have an option distribution of 1-3 or 3-1.
+After the aforementioned filtering process, we obtain 852 pieces of data in a 4-way multi-select format.
+DurationQA
+(Virgo et al., 2022) The original DurationQA has the same format as MCTACO, which consists of 694 questions with 4,868 options.
+Following the identical filtration procedure as MCTACO, we finally obtained a collection of 687 questions in a 4-way multi-select format.
+TimeDial
+(Qin et al., 2021) consists of 4-way multi-select instances in a two-person dialogue scenario.
+We leave the original data unaltered and simply randomize the sequence of options, yielding 1,446 pieces of 4-way multi-select instances.
+SituateGen
+(Zhang and Wan, 2023) SituatedGen includes 1,220 test cases, which span across two distinct reasoning domains: time and geography.
+We manually screen the original test data and retain those with clear time features for temporal reasoning evaluation, resulting in 115 instances.
+TimeQA
+(Chen et al., 2021) The original data of TimeQA includes two splits, Easy and Hard, with each question containing 20 Wikipedia paragraphs.
+The excessively long context may exceed the model’s maximum length limit and incur significant inference overhead.
+Therefore, we have reduced the context of the original data.
+For the paragraphs in the original data, we refer to those containing the answer as relevant paragraphs, and the rest as irrelevant paragraphs.
+For each question, we keep the first paragraph, all relevant paragraphs, and one random irrelevant paragraph as distractor.
+This ensures that most questions have at least three paragraphs.
+After that, we sample 500 pieces of data from those where the context length is less than 650 tokens.
+For both Easy and Hard splits, we apply the aforementioned filtration, resulting in 500 questions each, totaling 1,000 instances.
+TempReason
+(Tan et al., 2023) TempReason dataset contains 5,397 entries for l2 (event-time reasoning) and 4,426 entries for l3 (event-event reasoning).
+In the original dataset, each question corresponds with a text context and extracted facts.
+Similar to TimeQA, we apply a filter based on context length.
+We preserve questions with a context length between 300 and 600 tokens, yielding 839 and 1,037 instances, respectively.
+Notably, every remaining question is applicable to either context-based reasoning or fact-based reasoning.
+MenatQA
+(Wei et al., 2023) MenatQA consists of 999 data entries, formatted similarly to TimeQA, where each question is accompanied by several corresponding paragraphs.
+Following the paper’s proposed method, we modify the original data by incorporating the three time-sensitive factors: scope, order, and counterfactual.
+Subsequently, for each factor, we randomly sample 400 instances, resulting in a total of 1,200 data points.
+TRACIE
+(Zhou et al., 2021) The original TRACIE dataset consists of yes/no type questions, containing 4,248 test instances. We randomly sample 500 instances from the iid split in the test set.
+A.2 Human Performance Evaluation
+Unless otherwise stated, the results of human evaluation are derived from original dataset papers.
+Please refer to the corresponding paper for human evaluation details.
+TimeXNLI, Date Arith, and MCTACO are manually evaluated by three authors from the TimeBench team.
+Within each subtask, we randomly sample 50 instances, and the average of the performances by three human evaluators is considered the final human performance.
+A.3 Task Formats
+TimeBench is a multispectral benchmark, which features four different task formats.
+Multi-Select Questions
+Previous work utilizes the Multiple Choice (MC) form, which requires models to select the only correct answer from the options.
+However, this task form has shortcuts and may not truly reflect the model’s abilities.
+To address this, we employ the Multi-Select (M-S) task form, where the model needs to select all possible correct answers from the options provided.
+In our task, each question presents four options, with at most two of them being correct.
 Natural Language Inference
@@
 Our tasks focus on the entailment in temporal domains.
-Tasks: TimeX-NLI, TRACIE.
 Free-form Reading Comprehension
-requires models to answer questions based on the provided context, and the answer is free-form without pre-defined restrictions. Tasks: TimeQA, MenatQA, TempReason, Date Calculation.
+requires models to answer questions based on the provided context, and the ground truth answer is free-form without pre-defined format restrictions.
 Constrained Text Generation
 refers to the task of generating text under certain constraints.
-Our task is keyword-constrained text generation, where the model takes keywords as input and outputs sentences that include those keywords.
-Task: SituatedGen.
-Metrics
-We adopt the evaluation metrics from previous work.
-Accuracy is used for NLI and date calculation tasks.
-MAMCQ tasks are measured using option-based EM and F1.
-FRC tasks (w/o date calc.) are evaluated with token-based EM and F1.
-For CTG tasks, we use a composite generation metrics, with more details in Appendix A.2.
-3 Methodology
-We perform evaluations using the prompt-based approach, including standard I-O prompting and chain-of-thought prompting. Experiments are conducted under both zero-shot and few-shot settings.
-Standard I-O Prompting
-We formulate specific instructions for each task.
-In the zero-shot setting, models follow the instructions to answer questions.
-In the few-shot setting, models are provided with several question-answer pairs as demonstrations and emulate those instances to answer questions.
-$\mathrm{prompt_{zs}^{io}}=\{\mathrm{INST}\}\{\mathrm{Q}\}$ (1)
-$\mathrm{prompt_{fs}^{io}}=\{\mathrm{INST}\}\{\mathrm{Q_{1}}\}\{\mathrm{A_{1}}\}..\{\mathrm{Q}\}$ (2)
-Chain-of-Thought Prompting
-We use identical instructions to standard input-output prompting.
-In the zero-shot chain-of-thought scenario, following Zeroshot CoT (Kojima et al., 2022), we add a reasoning trigger Let’s think step by step after questions to perform chain-of-thought reasoning.
-In the few-shot chain-of-thought scenario, we manually craft rationales for each task as demonstrations to guide the model to reason step-by-step.
-The instructions and demonstrations can be found in the Appendix A.4.
-$\mathrm{prompt_{zs}^{cot}}=\{\mathrm{INST}\}\{{\mathrm{Q}}\}\{{\mathrm{TRIG}}\}$ (3)
-$\mathrm{prompt_{fs}^{cot}}=\{\mathrm{INST}\}\{{\mathrm{Q_{1}}}\}\{{\mathrm{R_{1}}}\}\{{\mathrm{A_{1}}}\}..\{{\mathrm{Q}}\}$ (4)
-4 Experimental Setup
-4.1 Models
-We evaluate several popular LLMs, including both open-source and closed-source models, with parameter sizes ranging from 6B to 175B [^1] Since OPENAI does not disclose ChatGPT-3.5/4’s size publicly, we use the size of GPT-3, which is 175 billion..
-Unless otherwise stated, we use aligned models in the following experiments.
-By default, we use 13B model for LLaMA2/Vicuna-1.5/Baichuan2 and 11B model for FLAN-T5.
-ChatGPT-3.5/GPT-4 (Ouyang et al., 2022; OpenAI, 2023)
-ChatGPT is a chat model aligned through SFT and RLHF based on GPT-3 (Brown et al., 2020).
+The task is keyword-constrained text generation, where the model takes keywords as input and outputs sentences that include those keywords.
+A.4 Evaluation Metrics
+Accuracy is used for NLI and date arithmetic tasks.
+M-S tasks are evaluated using option-level EM and F1.
+FRC tasks (excluding date arithmetic) are assessed with token-level EM and F1.
+For CTG task, we take the average of multiple generation metrics, which are outlined as follows.
+Metrics for SituatedGen
+Following SituatedGen (Zhang and Wan, 2023), we use BLEU-4 (Papineni et al., 2002), METEOR (Banerjee and Lavie, 2005), ROUGE-L (Lin, 2004), CIDEr (Vedantam et al., 2015), and MATCH (Zhang and Wan, 2023) scores to metric the results of CTG.
+[^3] We utilize pycocoevalcap package to calucate BLEU-4, METEOR, ROUGE-L, CIDEr.
+The overall score is calculated as the sum of the above scores. We set the weight of CIDEr to $1/10$ for balancing when summation.
+$S=\textrm{BLEU-4}+\textrm{METEOR}+\textrm{ROUGE-L}\\ +\textrm{CIDER}/10+\textrm{MATCH}$
+As the overall score $S$ does not represent a percentile, we proceed to normalize the models’ scores to align with humans’ relative performance levels.
+Figure 7: The quantity and proportion of data for each task and its respective subtasks within TimeBench.
+Appendix B Supplemental Materials
+B.1 Models
+ChatGPT-3.5/GPT-4
+(Ouyang et al., 2022; OpenAI, 2023) ChatGPT is a chat model aligned through SFT and RLHF based on GPT-3 (Brown et al., 2020).
 GPT-4 is an upgraded version of ChatGPT with enhanced reasoning capabilities, making it the most powerful LLM.
 Unless otherwise stated, ChatGPT refers to gpt-3.5-turbo-0613 and GPT-4 refers to gpt-4-0613.
-Llama2/Vicuna-1.5 (Touvron et al., 2023; Chiang et al., 2023)
-LLaMA2 is an open foundation model trained on 2T tokens with efficient grouped-query attention Ainslie et al.
+Llama2/Vicuna-1.5
+(Touvron et al., 2023; Chiang et al., 2023) LLaMA2 is an open foundation model trained on 2T tokens with efficient grouped-query attention Ainslie et al.
 (2023).
-LLaMA2-chat is the official aligned model with SFT and RLHF, and Vicuna-1.5 is an unofficial aligned model with SFT only.
-Baichuan2 (Yang et al., 2023)
-is an open foundation model pre-trained on 2.6T tokens and outperforms LLaMA2 on several benchmarks. Baichuan2-chat is the official aligned model with SFT and RLHF.
-Mistral (Jiang et al., 2023)
-is a 7B open foundation model incorporating efficient grouped-query attention (Ainslie et al., 2023) and sliding windows attention (Beltagy et al., 2020).
+LLaMA2-chat is the official aligned model with SFT and RLHF, and Vicuna-1.5 is aligned with SFT only by the community [^4] https://lmsys.org/.
+Baichuan2
+(Yang et al., 2023a) is an open foundation model pre-trained on 2.6T tokens, which is competitive with LLaMA2. Baichuan2-chat is the official aligned model with SFT and RLHF.
+Mistral
+(Jiang et al., 2023) is a 7B open foundation model incorporating efficient grouped-query attention (Ainslie et al., 2023) and sliding windows attention (Beltagy et al., 2020).
 It achieves the strongest performance among models of its size, even surpassing LLaMA2-13B.
 Mistral-instruct is the officially aligned model with SFT only.
-ChatGLM3 (Zeng et al., 2023)
-is an open-source bilingual model for Chinese and English, exhibiting competitive performance in models with sizes under 10B.
-FLAN-T5 (Shen et al., 2023)
-is an open-source instruction-following model built on top of T5 (Raffel et al., 2020) through instruction fine-tuning.
-4.2 Implementation Details
-We access closed-source models through OpenAI API.
-For open-source models, we deploy them locally and then access them through FastAPI.
-We set the temperature to 0 for greedy search in all experiments.
-In zero-shot-cot, we use Let’s think step by step as the reasoning trigger.
-To improve answer extraction accuracy, we use answer trigger Therefore, the answer is to prompt models to summarize the answer.
-In few-shot scenarios, we manually create 2-4 instances, varying with datasets.
-For more details about demonstrations and instructions, please refer to Appendix A.4.
-Method Symbolic Commonsense Event Temporal Overall
-TimeXNLI Arith DurationQA McTACO TimeDial SitGen TimeQA MenatQA TempReason TRACIE
-s1 s2 s3 Acc EM F1 EM F1 EM F1 Norm E-EM I-EM S-EM O-EM C-EM L2-EM L3-EM Acc Sym. Comm. Event Overall
-Human 98 96 92 100.0 64 80.8 75.8 87.1 97.8 - 100 89 87 82 84 76 96 94 82.5 96.5 84.3 88.0 88.5
-GPT-4 78.6 76.0 50.7 98.0 35.0 59.2 61.2 80.0 72.0 91.1 59.3 48.9 40.4 44.4 49.0 22.0 91.0 94.0 64.8 75.8 65.4 59.3 63.5
-  + CoT 80.0 76.0 60.0 92.0 35.0 58.1 67.0 82.6 65.0 89.3 - 50.0 33.0 43.4 53.0 20.0 93.0 93.0 58.0 77.0 66.2 58.3 63.2
-  + FS 85.3 73.3 53.3 100.0 51.0 64.8 77.5 88.3 85.0 94.6 88.6 59.2 40.0 59.6 48.0 25.3 86.0 94.8 62.8 78.0 78.6 63.0 69.5
-  + FS/CoT 92.0 84.0 64.0 100.0 42.0 55.1 68.0 72.3 79.0 93.4 - 48.0 44.4 48.5 44.0 22.0 91.0 93.0 66.4 85.0 68.3 60.8 66.5
-GPT-3.5 45.4 67.6 31.2 97.0 19.2 50.5 34.1 68.5 39.2 69.1 62.3 60.5 29.5 36.5 37.5 21.0 73.6 61.8 57.4 60.3 49.0 49.8 51.2
-  + CoT 33.6 64.8 33.6 71.0 12.4 23.2 28.1 45.1 34.6 67.0 - 52.5 29.0 35.8 38.5 24.0 32.0 54.2 52.0 50.8 35.1 43.5 42.6
-  + FS 52.0 68.4 31.6 63.6 42.8 67.7 43.5 71.2 47.8 76.4 79.1 53.8 37.9 37.8 43.5 16.0 77.7 70.0 55.0 53.9 61.2 52.1 54.8
-  + FS/CoT 51.6 71.8 36.6 84.4 20.8 41.2 21.4 38.1 48.3 71.1 - 56.5 37.5 38.1 37.5 33.0 86.2 68.0 50.2 61.1 40.2 54.0 51.8
-Baichuan2 41.0 61.0 37.0 12.5 4.0 52.0 18.5 63.4 15.0 57.7 52.2 45.0 29.0 31.0 34.0 30.0 40.0 45.0 49.0 37.9 37.5 42.9 40.7
-  + CoT 40.0 57.0 31.0 10.0 3.3 44.6 20.0 61.9 13.0 58.1 - 36.0 36.0 39.0 27.0 29.0 46.0 46.0 55.0 34.5 33.5 43.8 39.8
-  + FS 43.0 59.0 40.0 42.5 24.7 62.1 27.5 70.2 18.0 58.9 63.7 47.0 35.0 37.0 31.0 19.0 73.0 48.0 48.0 46.1 46.4 47.3 46.9
-  + FS/CoT 45.0 54.0 48.0 47.0 10.7 44.4 27.0 68.8 15.0 55.0 - 43.0 27.0 38.0 34.0 33.0 72.8 43.0 44.0 48.5 36.8 46.9 44.7
-LLaMA2 30.0 49.0 34.0 22.5 4.0 38.5 8.5 40.6 10.0 35.4 57.9 46.0 21.0 28.0 23.0 18.0 43.0 55.0 49.0 33.9 27.8 40.5 36.1
-  + CoT 36.0 50.0 38.0 6.0 7.3 39.1 14.0 51.7 10.0 36.9 - 45.0 30.0 20.0 18.0 21.0 43.0 56.0 47.0 32.5 26.5 40.7 36.0
-  + FS 43.0 57.0 60.0 20.5 9.0 46.8 8.0 66.5 15.0 62.3 40.2 24.0 17.0 11.0 5.0 22.0 54.0 50.0 47.0 45.1 35.4 32.6 35.3
-  + FS/CoT 37.0 55.0 50.0 33.0 12.0 49.5 11.0 45.6 8.0 44.5 - 35.0 21.0 34.0 23.0 7.0 72.0 54.0 50.0 43.8 28.4 40.9 38.4
-Vicuna-1.5 35.0 50.0 36.0 15.0 8.0 39.2 21.5 59.1 7.0 34.2 51.8 43.0 29.0 38.0 22.0 17.0 14.0 13.0 46.0 34.0 31.5 34.2 33.4
-  + CoT 42.0 51.0 37.0 3.0 1.3 29.8 11.5 50.0 7.0 33.7 - 44.0 31.0 16.0 25.0 13.0 31.0 29.0 51.0 33.3 22.2 35.2 31.8
-  + FS 48.0 57.0 38.0 30.5 7.3 33.6 27.5 57.0 13.0 40.3 39.0 45.0 23.0 38.0 26.0 18.0 51.0 28.0 56.0 43.4 31.1 38.5 37.3
-  + FS/CoT 38.0 59.0 39.0 39.5 10.7 37.4 14.0 45.8 12.0 41.6 - 47.0 27.0 39.0 31.0 26.0 71.0 53.0 52.0 43.9 26.9 46.3 41.3
-FLAN-T5 53.0 63.0 43.0 0.0 4.0 52.0 14.0 65.0 13.0 47.7 49.5 56.0 24.0 31.0 48.0 20.0 84.0 78.0 64.0 39.8 35.0 51.5 45.3
-  + CoT 56.0 66.0 45.0 0.0 4.7 49.7 14.5 63.4 13.0 42.7 - 57.6 23.9 39.0 46.0 28.0 73.0 57.0 55.0 41.8 31.3 49.5 43.9
-  + FS 53.0 65.0 43.0 3.5 4.0 50.2 15.5 65.8 9.0 35.0 60.2 55.0 23.0 31.0 47.0 21.0 82.0 78.0 65.0 41.1 34.2 51.3 45.1
-  + FS/CoT 54.0 68.0 46.0 3.5 4.0 50.7 13.0 64.0 11.0 43.7 - 54.0 19.0 34.0 45.0 20.0 89.0 72.0 66.0 42.9 31.1 50.7 44.7
-Mistral 47.0 50.0 43.0 26.5 11.3 49.8 15.0 58.8 6.0 23.2 58.3 7.0 5.0 4.0 7.0 4.0 2.0 1.0 51.0 41.6 31.8 18.0 25.3
-  + CoT 38.0 56.0 35.0 16.5 13.3 36.6 14.5 49.3 8.0 19.3 - 13.0 11.0 8.0 14.0 12.0 5.0 4.0 61.0 36.4 23.5 21.2 24.2
-  + FS 51.0 57.0 35.0 18.0 12.0 55.8 25.5 71.0 13.0 52.9 63.0 24.0 12.0 4.0 7.0 7.0 23.0 23.0 57.0 40.3 41.9 25.6 32.2
-  + FS/CoT 29.0 62.0 32.0 28.5 9.3 46.6 14.5 49.1 8.0 34.4 - 13.0 11.0 9.0 9.0 9.0 68.0 42.0 50.0 37.9 27.0 31.8 31.6
-${}^{\dagger}$ 38.0 48.0 33.0 42.5 20.7 54.8 42.5 73.0 11.0 45.7 64.9 50.0 40.0 42.0 31.0 13.0 68.0 50.0 54.0 40.4 44.6 47.6 45.7
-  + CoT 50.0 56.0 34.0 47.0 29.3 62.0 22.5 69.3 12.0 43.8 - 46.0 39.0 39.0 34.0 37.0 73.0 46.0 60.0 46.8 39.8 50.9 47.6
-${}^{\dagger}$ 50.0 54.0 30.0 29.5 19.3 53.3 26.5 66.0 20.0 55.6 64.8 48.0 34.0 41.0 38.0 34.0 68.0 49.0 58.0 40.9 43.7 50.0 46.9
-  + CoT 40.0 61.0 37.0 52.0 25.3 59.3 26.0 68.8 11.0 40.8 - 46.0 37.0 49.0 34.0 38.0 70.0 55.0 58.0 47.5 38.5 52.5 48.4
-${}^{\dagger}$ 48.0 53.0 38.0 41.0 34.0 61.8 42.5 76.2 35.0 61.8 58.3 43.0 30.0 37.0 38.0 37.0 68.0 64.0 53.0 45.0 52.8 51.0 50.6
-  + CoT 57.0 63.0 35.0 54.0 30.0 61.8 42.0 45.7 29.0 57.3 - 51.0 30.0 48.0 37.0 24.0 60.0 58.0 57.0 52.3 44.3 49.6 48.7
-Table 4: Main results of TimeBench.
-Aligned models are under the zero-shot setting by default.
-Methods with † are base models without alignment, under few-shot setting, thus incomparable with other methods.
-Some F1 scores are omitted, please refer to the appendix for details.
-Best results in each group are bold, top-3 results are underlined.
-5 Experimental Results
-5.1 Main Results
-We present the experimental results [^2] Please refer to Appendix A.3 for full results.
-of the aligned models in Table 4, covering the most popular LLMs.
-In most tasks, GPT-4 achieved the best performance, ranking first in 16 out of 19 evaluation metrics, surpassing the second-place GPT-3.5 by 14.7%.
-Nevertheless, it still exhibits a significant gap from human performance, with a 19% disparity.
-Among all open-source alignment models, Baichuan2 has achieved the best overall performance, reaching 85% performance of ChatGPT.
-Results on Symbolic Temporal Reasoning
-GPT-4 demonstrates strong symbolic reasoning performance, far surpassing ChatGPT and other models.
-While open-source models show poorer performance, with only Mistral-base achieving an overall score exceeding 50.
-In TimeX-NLI tasks, models other than GPT-4 perform poorly, barely surpassing a random baseline (33.3 for s1/s3, 50.0 for s2).
-This indicates that LLMs still have significant shortcomings in understanding abstract TimeX.
-Results on Commonsense Temporal Reasoning
-In temporal commonsense reasoning, GPT-4 still maintains a remarkable proximity to human-level performance.
-Among the open-source aligned models, Baichuan2 attains the best performance, securing the top three positions in four out of seven metrics.
-When extending to all open-source models, Mistral-base significantly outperforms others, surpassing the best open-source model by 6.4.
-Besides, we observe that chain-of-thought prompting has a negative impact on almost all models in commonsense reasoning.
-Furthermore, aligned models appear to exhibit varying degrees of performance degradation compared to the base model, with LLaMA2 and Mistral showing the most significant differences.
-Results on Event Temporal Reasoning
-GPT-4 remains the top performer in event temporal reasoning, though it is no longer markedly ahead of other models, and there is a notable disparity from human performance.
-Among all open-source models, LLaMA2-base achieved the best results, trailing GPT-3.5 by just 1.5%.
-Overall, open-source models show competitive performance in event temporal reasoning, narrowing the gap with closed-source models.
-In TRACIE, which involves implied events, most models only marginally surpass the random baseline (50.0), indicating persistent challenges in understanding implicit events with current models.
-Additionally, there is a large performance gap comparing context-based reasoning (TimeQA, MenatQA) with fact-based reasoning (TempReason), indicating a substantial potential for improvement in event-time facts grounding in unstructured text.
-Furthermore, within fact-based reasoning, models exhibit a noticeable decline in event-event reasoning compared to event-time reasoning, indicating that complex multi-hop relations among events increased the difficulty of reasoning.
-Figure 3: Performance difference between base and aligned models. Baichuan2 and LLaMA2 are aligned with SFT and RLHF. Vicuna, Mistral and ChatGLM3 are aligned with only SFT.
-6 Analysis and Discussion
-6.1 Alignment Degradation
-Alignment Causes Performance Degradation
-We notice that some aligned models exhibited performances inconsistent with their expectations, such as LLaMA2 and Mistral.
-We speculate that this is the negative impact brought about by the alignment phase.
-To investigate this, we conduct experiments on their corresponding base models in the few-shot setting, as shown in Table 4 (grey).
-We visualize the performance differences between the base and alignment models in the few-shot scenario, as shown in Figure 3.
-Baichuan2-13B shows a slight decrease in performance after alignment by 0.7%, while LLaMA2-13B, Vicuna-1.5-13B and Mistral-7B exhibit catastrophic performance drops, decreasing 10.0%, 7.1% and 18.4%, respectively.
-What leads to the degradation?
-Combining experimental results with model output cases, we attribute this phenomenon to two reasons:
-  • The training during the alignment phase has to some extent disrupted the original language modeling capability, causing the model to deviate from strictly adhering to demonstrations during in-context learning.
-  • Excessive alignment of models towards conversation and harmlessness has led to a decrease in its usability, resulting in generating instruction-irrelevant content or refusing to follow instructions.
-Figure 4: Performance gap with and without CoT prompts. The results are averaged from GPT-4, GPT-3.5, Baichuan2, LLaMA2/Vicuna-1.5 and Mistral.
-6.2 Chain-of-Thought in Temporal Reasoning
-Previous research has found that chain-of-thought prompting (Wei et al., 2022; Kojima et al., 2022; Chu et al., 2023a) enables models to think step by step, facilitating the resolution of complex reasoning tasks such as mathematical and logical reasoning.
-How about applying cot prompting to temporal reasoning, will it bring consistent improvement?
-To investigate this, we select several popular models and analyze their performance affected when using chain-of-thought prompts across various task categories.
-Chain-of-Thought inconsistently boosts, or even impairs performance
-As illustrated in Figure 4, introducing zero-shot chain-of-thought prompting results in an overall performance decline of 6%, with even a 16% degradation in commonsense reasoning.
-In the few-shot scenario, chain-of-thought prompts also fail to yield comprehensive improvements, showing varied effects depending on the task type, resulting in a 13% improvement in symbolic reasoning but an 18% decrease in commonsense reasoning.
-In event temporal reasoning, cot prompting in both scenarios only resulted in minor impacts, with -1% and 1.7%, respectively.
-In summary, zero-shot chain-of-thought consistently hinders performance, whereas few-shot chain-of-thought demonstrates performance variations dependent on the nature of the task.
-Figure 5: $\Delta\text{Score}$ between the chain-of-thought prompting and direct I-O prompting.
-Top: zero-shot setting, Bottom: few-shot setting, Left: variation in each task, Right: averaged variation in the symbolic, commonsense, event, and overall tasks.
-Impact of Chain-of-Thought across tasks
-In order to explore the impact of CoT on various tasks thoroughly, we delve into the performance changes of each model across specific tasks within each category, as illustrated in Figure  5.
-In the zero-shot setting, open-source models achieve a slight improvement in event temporal reasoning with chain-of-thought prompting, while in other cases, they face performance degradation.
-While in the few-shot setting, almost all models exhibit significant improvement in symbolic temporal reasoning, with a concurrent prevalent decline in commonsense temporal reasoning.
-Differing from the previous uniform trends observed within the same task category, event temporal reasoning exhibits both increases and decreases simultaneously.
-We attribute this to the nature of the datasets; both MenatQA and TempReason involve implicit reasoning, encompassing temporal factors and event-event relationships, requiring multi-hop reasoning for resolution.
-This aligns well with the step-by-step problem-solving nature of cot prompting, resulting in performance improvement.
-On the other hand, TimeQA only necessitates implicit reasoning for a subset of questions, making cot prompts less impactful.
-This is reflected in the marginal performance change with and without chain-of-thought prompting.
-6.3 Impact of Model Size
-We investigate the relationship between the performance and parameter size of base models with decoder architecture ranging from 6B to 13B.
-With the scaling up of the parameter size for the same model, there is a notable performance enhancement.
-When the parameter size expands from 7B to 13B, LLaMA2 and Baichuan2 show performance improvements of 14.2% and 11.1%, respectively.
-Under the same 13B parameter size, LLaMA2 slightly outperforms Baichuan2 by 1.6%.
-With around 7B parameter size, there is no significant difference in performance among LLaMA2, Baichuan2, ChatGLM3, while Mistral exhibits extremely strong performance, even surpassing LLaMA2-13B by as much as 4.5% with only 53% parameters.
-6.4 Dive into Each Categories
-Model Order Duration Freq. Stationarity Typical Avg.
-GPT-4 76.4 $\downarrow$ 92.8 $\uparrow$ 83.3 $\uparrow$ 71.4 $\downarrow$ 54.5 $\downarrow$ 77.5
-GPT-3.5 50.5 $\uparrow$ 39.8 $\downarrow$ 55.2 $\uparrow$ 48.4 $\uparrow$ 28.7 $\downarrow$ 43.5
-${}^{\dagger}$ 40.5 $\downarrow$ 51.8 $\uparrow$ 43.7 $\uparrow$ 46.2 $\uparrow$ 29.8 $\downarrow$ 42.5
-${}^{\dagger}$ 18.9 $\downarrow$ 37.0 $\uparrow$ 27.1 $\uparrow$ 7.6 $\downarrow$ 25.5 $\downarrow$ 26.5
-${}^{\dagger}$ 27.0 $\downarrow$ 44.4 $\uparrow$ 58.3 $\uparrow$ 38.5 $\downarrow$ 38.3 $\downarrow$ 42.5
-Table 5: Results in each temporal commonsense aspect under few-shot setting.
-Models with †  are base models.
-Red $\downarrow$ and Green $\uparrow$ represent the performance is lower or higher than its average.
-LLM underperforms in (multi-hop) symbolic reasoning
-Except for GPT-4, the performance of other models in symbol temporal reasoning is unsatisfactory.
-Our experimental results reveal a notable decline in models’ performance on the duration-conversion task (s3) compared to other atomic tasks (s1, s2, arith).
-Delving into the reasons, task s3 necessitates a two-step reasoning process, initially unifying time units and subsequently engaging in numerical comparison.
-In contrast, other atomic tasks can be completed with a single inference.
-To sum up, LLMs show unsatisfactory performance in symbol temporal reasoning, with a more pronounced decline when confronted with multi-step reasoning.
-LLM’s mastery of temporal commonsense knowledge differs across aspects.
-We analyze models’ performance across various common-sense aspects, as shown in Table 5.
-We use the model’s average performance in common-sense tasks as a baseline, if the model outperforms its own average in a specific aspect, it suggests greater proficiency and vice versa.
-The findings show that models generally demonstrate good knowledge of event duration and frequency, while their grasp of event order and typical events is comparatively weaker.
-LLM demonstrates poor implicit temporal reasoning abilities
-Comparing explicit and implicit context-based reasoning (TimeQA-explicit v.s.
-others), we note a marked decrease in the model’s performance in implicit reasoning.
-Additionally, on TRACIE with numerous implied events, most models exceed only a random baseline (50.0).
-Even GPT-4 achieves a mere 66.4% accuracy.
-To summarize, LLM performs inadequately when dealing with different types of implicit reasoning, constraining its reasoning abilities in complex scenarios.
-LLM is a good factual reasoner rather than factual extractor
-When humans engage in context-based temporal reasoning, it typically involves two steps: initially extracting time-fact pairs from the context and subsequently conducting fact-based reasoning.
-TempReason provides extracted facts for conducting fact-based reasoning.
-By comparing the model’s performance in context-based reasoning and fact-based reasoning, we can identify the bottlenecks in LLM’s event temporal reasoning.
-Our findings reveal that LLM excels in TempReason, signifying its strong capability in fact-based reasoning.
-However, LLM’s performance in context-based reasoning was significantly weaker than in the former.
-We attribute this gap to the model’s deficiency in factual extraction capabilities.
-This gap highlights limitations in context-based temporal reasoning.
-7 Related Work
-7.1 Temporal Reasoning
-Numerous research efforts address diverse challenges in temporal reasoning.
-Early research mainly relies on TimeML (Pustejovsky et al., 2003), focusing TimeX extraction and temporal relation extraction (Verhagen et al., 2007, 2010; UzZaman et al., 2013; Llorens et al., 2015; Miller et al., 2015; Mathur et al., 2021; Vashishtha et al., 2019).
-The advent of pre-trained language models (PLMs) has brought about commonsense reasoning as a tool to explore the world knowledge in models (Zhou et al., 2019; Qin et al., 2021; Dhingra et al., 2022).
-Recently, much attention has shifted towards event temporal reasoning (Chen et al., 2021; Tan et al., 2023; Zhu et al., 2023; Son and Oh, 2023; Chu et al., 2023b).
-Besides, Wang and Zhao (2023) introduces a unified form for assessing the temporal understanding capability of language models.
-Distinguished from other datasets and benchmarks, TimeBench is multispectral, and closely aligned with intricate real-world scenarios, offering a comprehensive and hierarchical evaluation of LLM’s temporal reasoning abilities.
-7.2 Large-scaled Language Model
-In recent years, there has been rapid progress in the research of large-scale language models (LLM) (Zhao et al., 2023).
-They exhibit outstanding performance across a multitude of tasks without the need for fine-tuning (Brown et al., 2020; Kojima et al., 2022).
-Furthermore, they have achieved astonishing results in complex reasoning tasks, such as mathematical reasoning (Cobbe et al., 2021; Mishra et al., 2022) and logical reasoning (Yu et al., 2020; Liu et al., 2023).
-Moreover, some studies suggest that the chain-of-thought prompting can further enhance the model’s capabilities in complex reasoning scenario (Wei et al., 2022; Kojima et al., 2022; Chu et al., 2023a).
-8 Conclusion
-Time is a crucial dimension for understanding the world, and temporal reasoning inherently possesses complexity.
-The lack of a comprehensive benchmark makes it challenging to quantify LLMs’ temporal reasoning capabilities.
-In this work, we present TimeBench, a comprehensive and hierarchical benchmark for LLM temporal reasoning, tailored to mirror temporal reasoning in complex real-world scenarios.
-We conduct extensive experiments on state-of-the-art LLMs to investigate their temporal reasoning capabilities.
-Our findings indicate a substantial gap between state-of-the-art LLM and human performance, emphasizing the need for further research in this area.
-Moreover, we provide a meticulous analysis and discussion, outlining the current challenges that models face and suggesting potential directions for improvement.
-Appendix A Appendix
-A.1 Benchmark Details
-We present the subtasks for each dataset in TimeBench, including data quantities, task formats, challenges faced by each task, and the capabilities required by models, as shown in Table 6.
-A.2 Metrics for CTG
-Following SituatedGen (Zhang and Wan, 2023), we use BLEU-4 (Papineni et al., 2002), METEOR (Banerjee and Lavie, 2005), ROUGE-L (Lin, 2004), CIDEr(Vedantam et al., 2015), and MATCH (Zhang and Wan, 2023) scores to metric the results of Constrained Text Generation.
-[^3] We utilize pycocoevalcap package to calucate BLEU-4, METEOR, ROUGE-L, CIDEr.
-The overall score is calculated as the sum of the above scores. We set the weight of CIDEr to $1/10$ for balance when summing.
-$S=\textrm{BLEU-4}+\textrm{METEOR}+\textrm{ROUGE-L}\\ +\textrm{CIDER}/10+\textrm{MATCH}$
-As the overall score $S$ does not represent a percentile, we proceeded to normalize the models’ scores to align with humans’ relative performance levels.
-A.3 Full Results
+ChatGLM3
+(Zeng et al., 2023) is an open-source bilingual LLM for Chinese and English, exhibiting competitive performance under 10B.
+FLAN-T5
+(Chung et al., 2022) is an open-source instruction model built on top of T5 (Raffel et al., 2020) through instruction fine-tuning.
+B.2 Full Results
 The overall score is derived from the average of all corresponding metrics.
-For brevity, we omit some F1 scores in the main tables in the main text.
-Please refer to Table 7 for the full experimental results, and detailed metrics for SituatedGen can be found in Table 8.
-A.4 Prompts
-The prompt instructions are showcased in Figure 6. The demonstrations can be found from Figure 7 to Figure 15.
+For brevity, we omit some F1 scores in tables in the main text.
+Please refer to Table LABEL:tab:full_results for the full experimental results.
+The full results of SituatedGen can be found in Table 8.
+B.3 Prompts
+The prompt formats are showcased in Figure 9. The demonstrations can be found from Figure 10 to 18.
+Figure 8: Performance comparison between state-of-the-art LLMs. Up: GPT-4/3.5 and alignment models under zero-shot setting. Down: GPT-4/3.5 and base models under few-shot setting.
 Dataset Format # Challenges
@@
 TimeX NLI NLI 6,965 TimeX Causality
-   - Order - 2,213 order
-   - Duration - 2,332 duration
-   - Conversion - 2,420 duration + time unit conversion
+- Order - 2,213 order
+- Duration - 2,332 duration
+- Conversion - 2,420 duration + time unit conversion
 Commonsense
-MCTACO MCMAQ 852 Temporal Commonsense
-TimeDial MCMAQ 1,446 Temporal Commonsense
-DurationQA MCMAQ 687 Event Duration
+MCTACO M-S 852 Temporal Commonsense
+TimeDial M-S 1,446 Temporal Commonsense
+DurationQA M-S 687 Event Duration
 SituatedGen CTG 115 Temporal Commonsense
@@
 TimeQA FRC 1,000 Context-based Reasoning
-   - Explicit - 500 explicit, event-time reasoning
-   - Implicit - 500 implicit, event-time reasoning
+- Explicit - 500 explicit, event-time reasoning
+- Implicit - 500 implicit, event-time reasoning
 MenatQA FRC 1,599 Implicit, Context-based Reasoning
-   - Order - 400 event-time reasoning
-   - Scope - 400 event-time reasoning
-   - Counterfactual - 400 event-time reasoning
+- Order - 400 event-time reasoning
+- Scope - 400 event-time reasoning
+- Counterfactual - 400 event-time reasoning
 TempReason FRC 1,876 Implicit, Fact-based Reasoning
-   - l2 (e2t) - 839 event-time reasoning
-   - l3 (e2e) - 1,037 event-event reasoning
+- l2 (e2t) - 839 event-time reasoning
+- l3 (e2e) - 1,037 event-event reasoning
 TRACIE NLI 500 Implicit, Implied Event-Event Reasoning
 In total 19,000
-Table 6: The task formats and challenges in TimeBench.
+Table 7: The statistics, task formats and challenges in TimeBench.
+Method BLEU-4 METEOR ROUGE-L CIDEr MATCH Overall Norm
+Human 39.9 40.4 56.3 397 98.1 274.4 100.0
+GPT-4 8.23 31.27 28.84 38.45 90.41 162.59 59.25
++ FS 28.64 38.99 55.69 298.64 90.11 243.29 88.66
+GPT-3.5 13.38 30.12 35.91 125.41 78.76 170.70 62.21
++ FS 27.24 33.77 51.18 282.75 76.54 217.01 79.08
+LLaMA270b 5.15 13.62 15.83 22.07 31.79 68.60 25.00
++ FS 19.10 29.09 41.74 171.36 65.29 172.35 62.81
+LLaMA213b 4.66 21.43 20.80 17.72 61.62 110.28 40.19
++ FS 15.15 27.49 37.55 138.13 64.94 158.93 57.92
+LLaMA27b 2.77 13.46 14.69 14.34 34.83 67.18 24.48
++ FS 6.90 15.82 21.77 52.99 33.81 83.60 30.47
+Baichuan213b 8.33 25.86 30.07 82.63 70.63 143.15 52.17
++ FS 15.79 30.23 40.96 169.14 71.01 174.91 63.74
+Baichuan27b 5.17 21.99 23.73 44.80 59.85 115.22 41.99
++ FS 15.06 23.45 32.29 137.94 52.04 136.64 49.79
+Vicuna1.513b 7.73 26.35 29.15 69.16 71.91 142.06 51.77
++ FS 6.85 18.66 25.99 92.96 46.19 106.99 38.99
+Vicuna1.57b 6.29 24.34 26.91 46.90 68.84 131.07 47.77
++ FS 20.71 30.19 45.20 203.20 67.58 184.00 67.05
+FLAN-T5 16.20 24.43 29.38 95.17 56.38 135.91 49.53
++ FS 12.88 30.38 36.27 92.20 76.44 165.19 60.20
+Mistral7b 5.82 22.89 24.19 44.03 63.74 121.03 44.11
++ FS 18.96 29.02 43.15 185.61 63.24 172.93 63.02
+ChatGLM36b 6.56 21.11 21.96 41.48 53.02 106.80 38.92
++ FS 10.53 24.17 33.44 124.50 56.94 137.53 50.12
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+Table 8: Full results of SituatedGen.
+Aligned models are under zero-shot setting by default.
+The top-3 results are bold.
+Methods with † are base models without alignment, under few-shot setting.
+We consider human performance as 100 points and normalize models’ results accordingly.
 Table 9: Full results of TimeBench.
@@
 Method Symbolic Commonsense Event Overall
-TimeXNLI Date Arith DurationQA McTACO TimeDial SitGen TimeQA MenatQA TempReason TRACIE
-s1 s2 s3 Acc EM F1 EM F1 EM F1 Norm E-EM E-F1 H-EM H-F1 S-EM S-F1 O-EM O-F1 C-EM C-F1 L2-EM L2-F1 L3-EM L3-F1 Acc Sym. Comm. Event Overall
-Human 98 96 92 100 64 80.8 75.8 87.1 97.8 - 100 89 93.3 87 91.1 82 85.6 84 87.3 76 79.9 96 97.1 94 95.3 82.5 96.5 84.3 88.0 88.5
-GPT-4 78.6 76.0 50.7 98.0 35.0 59.2 61.2 80.0 72.0 91.1 59.3 48.9 60.6 40.4 46.5 44.4 57.0 49.0 57.0 22.0 23.1 91.0 95.3 94.0 95.0 64.8 75.8 65.4 59.3 63.5
-  + CoT 80.0 76.0 60.0 92.0 35.0 58.1 67.0 82.6 65.0 89.3 - 50.0 61.3 33.0 41.2 43.4 54.6 53.0 59.6 20.0 22.6 93.0 97.0 93.0 94.5 58.0 77.0 66.2 58.3 63.2
-  + FS 85.3 73.3 53.3 100 51.0 64.8 77.5 88.3 85.0 94.6 88.6 59.2 73.7 40.0 51.0 59.6 72.4 48.0 54.8 25.3 28.7 86.0 92.4 94.8 95.9 62.8 78.0 78.6 63.0 69.5
-  + FS/CoT 92.0 84.0 64.0 100.0 42.0 55.1 68.0 72.3 79.0 93.4 - 48.0 66.9 44.4 52.8 48.5 65.3 44.0 52.6 22.0 25.9 91.0 96.9 93.0 94.6 66.4 85.0 68.3 60.8 66.5
-GPT-3.5 45.4 67.6 31.2 97.0 19.2 50.5 34.1 68.5 39.2 69.1 62.3 60.5 70.8 29.5 35.4 36.5 40.9 37.5 43.9 21.0 22.9 73.6 81.2 61.8 73.8 57.4 60.3 49.0 49.8 51.2
-  + CoT 33.6 64.8 33.6 71.0 12.4 23.2 28.1 45.1 34.6 67.0 - 52.5 64.4 29.0 35.1 35.8 39.7 38.5 42.9 24.0 26.3 32.0 57.6 54.2 68.0 52.0 50.8 35.1 43.5 42.6
-  + FS 52.0 68.4 31.6 63.6 42.8 67.7 43.5 71.2 47.8 76.4 79.1 53.8 66.1 37.9 48.4 37.8 43.2 43.5 51.6 16.0 17.9 77.7 84.7 70.0 78.0 55.0 53.9 61.2 52.1 54.8
-  + FS/CoT 51.6 71.8 36.6 84.4 20.8 41.2 21.4 38.1 48.3 71.1 - 56.5 68.0 37.5 47.0 38.1 42.5 37.5 41.7 33.0 37.8 86.2 89.9 68.0 76.6 50.2 61.1 40.2 54.0 51.8
-Baichuan2 41.0 61.0 37.0 12.5 4.0 52.0 18.5 63.4 15.0 57.7 52.2 45.0 55.4 29.0 34.6 31.0 48.8 34.0 44.3 30.0 39.5 40.0 57.4 45.0 61.4 49.0 37.9 37.5 42.9 40.7
-  + CoT 40.0 57.0 31.0 10.0 3.3 44.6 20.0 61.9 13.0 58.1 - 36.0 41.5 36.0 40.9 39.0 52.0 27.0 38.5 29.0 43.2 46.0 62.8 46.0 64.3 55.0 34.5 33.5 43.8 39.8
-  + FS 43.0 59.0 40.0 42.5 24.7 62.1 27.5 70.2 18.0 58.9 63.7 47.0 60.7 35.0 45.7 37.0 51.9 31.0 41.5 19.0 31.8 73.0 81.1 48.0 59.4 48.0 46.1 46.4 47.3 46.9
-  + FS/CoT 45.0 54.0 48.0 47.0 10.7 44.4 27.0 68.8 15.0 55.0 - 43.0 57.8 27.0 36.7 38.0 49.8 34.0 40.7 33.0 43.0 72.8 80.4 43.0 60.2 44.0 48.5 36.8 46.9 44.7
-LLaMA2 30.0 49.0 34.0 22.5 4.0 38.5 8.5 40.6 10.0 35.4 57.9 46.0 61.9 21.0 30.5 28.0 46.1 23.0 36.1 18.0 26.9 43.0 53.1 55.0 69.4 49.0 33.9 27.8 40.5 36.1
-  + CoT 36.0 50.0 38.0 6.0 7.3 39.1 14.0 51.7 10.0 36.9 - 45.0 58.7 30.0 38.9 20.0 40.9 18.0 32.5 21.0 33.6 43.0 58.0 56.0 68.4 47.0 32.5 26.5 40.7 36.0
-  + FS 43.0 57.0 60.0 20.5 9.0 46.8 8.0 66.5 15.0 62.3 40.2 24.0 34.2 17.0 18.4 11.0 25.9 5.0 14.6 22.0 33.3 54.0 68.1 50.0 64.8 47.0 45.1 35.4 32.6 35.3
-  + FS/CoT 37.0 55.0 50.0 33.0 12.0 49.5 11.0 45.6 8.0 44.5 - 35.0 46.0 21.0 25.4 34.0 46.7 23.0 36.5 7.0 16.5 72.0 80.8 54.0 66.2 50.0 43.8 28.4 40.9 38.4
-Vicuna-1.5 35.0 50.0 36.0 15.0 8.0 39.2 21.5 59.1 7.0 34.2 51.8 43.0 60.4 29.0 37.0 38.0 46.8 22.0 37.4 17.0 23.2 14.0 42.1 13.0 43.6 46.0 34.0 31.5 34.2 33.4
-  + CoT 42.0 51.0 37.0 3.0 1.3 29.8 11.5 50.0 7.0 33.7 - 44.0 56.9 31.0 36.4 16.0 38.2 25.0 37.7 13.0 20.4 31.0 49.0 29.0 49.1 51.0 33.3 22.2 35.2 31.8
-  + FS 48.0 57.0 38.0 30.5 7.3 33.6 27.5 57.0 13.0 40.3 39.0 45.0 58.3 23.0 25.9 38.0 42.6 26.0 41.4 18.0 20.1 51.0 61.8 28.0 42.6 56.0 43.4 31.1 38.5 37.3
-  + FS/CoT 38.0 59.0 39.0 39.5 10.7 37.4 14.0 45.8 12.0 41.6 - 47.0 59.5 27.0 30.7 39.0 48.1 31.0 35.9 26.0 31.2 71.0 77.5 53.0 65.5 52.0 43.9 26.9 46.3 41.3
-FLAN-T5 53.0 63.0 43.0 0.0 4.0 52.0 14.0 65.0 13.0 47.7 49.5 56.0 61.7 24.0 26.8 31.0 33.6 48.0 52.2 20.0 21.8 84.0 87.9 78.0 83.9 64.0 39.8 35.0 51.5 45.3
-  + CoT 56.0 66.0 45.0 0.0 4.7 49.7 14.5 63.4 13.0 42.7 - 57.6 64.4 23.9 28.2 39.0 41.6 46.0 50.2 28.0 30.6 73.0 79.5 57.0 68.9 55.0 41.8 31.3 49.5 43.9
-  + FS 53.0 65.0 43.0 3.5 4.0 50.2 15.5 65.8 9.0 35.0 60.2 55.0 64.3 23.0 25.0 31.0 33.6 47.0 50.6 21.0 22.4 82.0 87.0 78.0 84.5 65.0 41.1 34.2 51.3 45.1
-  + FS/CoT 54.0 68.0 46.0 3.5 4.0 50.7 13.0 64.0 11.0 43.7 - 54.0 59.4 19.0 21.2 34.0 36.6 45.0 49.2 20.0 21.7 89.0 93.8 72.0 79.7 66.0 42.9 31.1 50.7 44.7
-Mistral 47.0 50.0 43.0 26.5 11.3 49.8 15.0 58.8 6.0 23.2 58.3 7.0 28.2 5.0 21.4 4.0 24.3 7.0 22.3 4.0 21.7 2.0 39.6 1.0 31.6 51.0 41.6 31.8 18.0 25.3
-  + CoT 38.0 56.0 35.0 16.5 13.3 36.6 14.5 49.3 8.0 19.3 - 13.0 31.3 11.0 22.4 8.0 21.1 14.0 24.9 12.0 25.6 5.0 34.0 4.0 31.2 61.0 36.4 23.5 21.2 24.2
-  + FS 51.0 57.0 35.0 18.0 12.0 55.8 25.5 71.0 13.0 52.9 63.0 24.0 43.5 12.0 23.9 4.0 21.3 7.0 21.2 7.0 23.0 23.0 48.9 23.0 44.9 57.0 40.3 41.9 25.6 32.2
-  + FS/CoT 29.0 62.0 32.0 28.5 9.3 46.6 14.5 49.1 8.0 34.4 - 13.0 33.7 11.0 25.3 9.0 26.4 9.0 24.3 9.0 20.3 68.0 78.6 42.0 57.4 50.0 37.9 27.0 31.8 31.6
-ChatGLM3 38.0 50.0 34.0 2.0 3.0 34.1 7.0 43.6 14.0 56.7 38.9 20.0 41.2 14.0 31.7 25.0 33.8 17.0 26.0 24.0 32.2 42.0 57.0 30.0 54.0 50.0 31.0 28.2 33.2 31.5
-  + CoT 27.0 49.0 37.0 0.0 1.0 24.8 3.0 37.1 10.0 44.8 - 24.0 41.7 10.0 25.4 27.0 34.6 22.0 28.1 35.0 41.2 28.0 44.5 27.0 52.0 48.0 28.3 20.1 32.6 28.9
-  + FS 37.0 52.0 30.0 0.0 10.0 53.0 9.0 52.9 19.0 54.7 50.1 11.0 19.6 13.0 17.0 25.0 35.0 24.0 30.4 20.0 25.6 27.0 46.5 33.0 53.3 54.0 29.8 35.5 29.0 30.9
-  + FS/CoT 32.0 66.0 31.0 0.0 4.0 34.8 8.0 43.6 11.0 44.0 - 25.0 43.3 19.0 27.5 23.0 30.2 19.0 25.7 36.0 43.0 50.0 62.8 34.0 56.0 48.0 32.3 24.2 36.2 32.7
-Baichuan2-7B 35.0 50.0 37.0 4.5 4.0 47.9 10.5 55.3 15.0 54.3 42.0 26.0 41.5 20.0 34.7 20.0 35.1 19.0 31.2 6.0 20.4 22.0 43.4 29.0 47.7 55.0 31.6 32.7 30.1 31.0
-  + CoT 38.0 43.0 32.0 1.0 5.3 37.9 13.0 58.0 15.0 44.2 - 41.0 53.5 28.0 38.8 29.0 39.9 23.0 33.2 18.0 29.3 21.0 41.2 29.0 47.2 54.0 28.5 28.9 35.1 32.5
-  + FS 40.0 50.0 36.0 20.0 28.7 59.4 26.5 66.9 17.0 53.0 49.8 45.0 60.7 30.0 42.1 27.0 37.8 23.0 35.7 10.0 20.4 40.0 57.4 37.0 53.0 51.0 36.5 43.0 38.0 39.1
-  + FS/CoT 41.0 50.0 36.0 23.5 13.0 45.7 17.5 58.1 7.0 39.2 - 36.0 51.2 29.0 43.0 42.0 52.5 25.0 39.3 20.0 31.0 57.0 70.1 39.0 60.2 49.0 37.6 30.1 43.0 39.0
-LLaMA2-7B 39.0 53.0 30.0 13.0 2.7 39.3 4.0 41.0 1.0 6.3 24.5 37.0 49.0 14.0 29.0 7.0 26.8 8.0 21.1 9.0 16.0 48.0 63.9 32.0 47.9 49.0 33.8 17.0 30.5 27.4
-  + CoT 44.0 50.0 33.0 5.0 2.7 35.0 4.5 40.0 1.0 1.7 - 27.0 49.9 17.0 31.6 11.0 31.4 10.0 24.5 7.0 17.8 44.0 56.9 32.0 48.1 46.0 33.0 14.1 30.3 26.8
-  + FS 44.0 60.0 34.0 11.0 4.0 62.8 8.0 64.7 8.0 40.0 30.5 36.0 50.8 20.0 29.4 5.0 22.3 6.0 18.0 6.0 17.9 12.0 36.3 23.0 44.3 53.0 37.3 31.1 25.3 28.7
-  + FS/CoT 38.0 51.0 36.0 14.5 11.0 42.8 25.0 65.6 13.0 53.4 - 36.0 53.5 21.0 34.1 1.0 13.6 3.0 11.2 5.0 14.0 22.0 46.7 21.0 42.3 51.0 34.9 35.1 25.0 29.0
-Vicuna-1.5-7B 37.0 58.0 43.0 5.0 1.3 40.4 9.5 52.5 6.0 32.0 47.8 35.0 47.1 11.0 18.5 20.0 35.7 15.0 25.7 12.0 17.3 14.0 33.0 14.0 46.8 54.0 35.8 27.1 26.6 28.1
-  + CoT 36.0 50.0 36.0 1.5 1.3 39.4 8.5 49.2 9.0 36.2 - 30.0 40.9 14.0 24.6 16.0 26.2 14.0 28.5 12.0 25.0 9.0 27.7 7.0 40.3 54.0 30.9 23.9 24.6 25.5
-  + FS 43.0 57.0 37.0 8.5 3.3 44.6 5.5 42.1 7.0 36.8 67.0 24.0 31.9 12.0 14.9 16.0 21.8 20.0 27.5 17.0 22.2 13.0 34.3 6.0 32.2 54.0 36.4 29.5 23.1 26.9
-  + FS/CoT 35.0 54.0 35.0 8.0 2.7 37.2 10.0 47.5 10.0 41.3 - 31.0 39.9 13.0 16.6 15.0 26.7 15.0 23.8 16.0 23.1 55.0 66.3 32.0 48.1 43.0 33.0 24.8 31.0 29.8
-${}^{\dagger}$ 38.0 48.0 33.0 42.5 20.7 54.8 42.5 73.0 11.0 45.7 64.9 50.0 59.4 40.0 54.2 42.0 52.7 31.0 38.0 13.0 21.4 68.0 77.3 50.0 63.5 54.0 40.4 44.6 47.6 45.7
-  + CoT 50.0 56.0 34.0 47.0 29.3 62.0 22.5 69.3 12.0 43.8 - 46.0 58.2 39.0 49.6 39.0 49.8 34.0 40.1 37.0 45.6 73.0 81.3 46.0 65.6 60.0 46.8 39.8 50.9 47.6
-${}^{\dagger}$ 50.0 54.0 30.0 29.5 19.3 53.3 26.5 66.0 20.0 55.6 64.8 48.0 59.3 34.0 48.6 41.0 49.6 38.0 43.4 34.0 37.5 68.0 78.7 49.0 62.7 58.0 40.9 43.7 50.0 46.9
-  + CoT 40.0 61.0 37.0 52.0 25.3 59.3 26.0 68.8 11.0 40.8 - 46.0 59.4 37.0 49.1 49.0 58.4 34.0 43.8 38.0 44.1 70.0 78.0 55.0 68.2 58.0 47.5 38.5 52.5 48.4
-${}^{\dagger}$ 27.0 66.0 41.0 32.5 28.0 59.8 34.5 69.4 5.0 34.3 59.8 40.0 53.8 35.0 50.1 41.0 49.6 33.0 38.5 18.0 22.9 49.0 65.9 34.0 51.0 55.0 41.6 41.5 42.4 42.1
-  + CoT 30.0 56.0 34.0 34.0 23.3 57.0 33.0 69.5 12.0 44.5 - 41.0 51.2 31.0 40.7 38.0 46.4 26.0 32.6 41.7 46.3 46.0 61.5 43.8 64.1 53.0 38.5 39.9 44.2 42.3
-${}^{\dagger}$ 26.0 50.0 30.0 20.0 19.3 54.5 20.0 59.6 15.0 45.2 62.4 44.0 54.4 30.0 45.3 42.0 49.8 34.0 41.9 30.0 35.8 50.0 64.0 36.0 53.3 49.0 31.5 39.4 44.0 40.8
-  + CoT 37.0 52.0 36.0 25.5 21.3 56.9 26.5 67.0 16.0 41.9 - 32.0 45.6 27.0 36.1 41.0 50.9 30.0 38.0 51.0 57.3 45.0 59.7 37.0 57.7 50.0 37.6 38.3 43.9 41.5
-${}^{\dagger}$ 48.0 53.0 38.0 41.0 34.0 61.8 42.5 76.2 35.0 61.8 58.3 43.0 55.9 30.0 45.3 37.0 49.4 38.0 47.8 37.0 45.5 68.0 76.7 64.0 74.8 53.0 45.0 52.8 51.0 50.6
-  + CoT 57.0 63.0 35.0 54.0 30.0 61.8 42.0 45.7 29.0 57.3 - 51.0 60.4 30.0 46.2 48.0 57.2 37.0 47.9 24.0 33.2 60.0 65.9 58.0 67.9 57.0 52.3 44.3 49.6 48.7
-${}^{\dagger}$ 48.0 70.0 32.0 35.0 3.3 51.8 13.5 62.6 11.0 55.0 61.6 50.0 57.2 24.0 26.3 30.0 35.4 38.0 41.5 22.0 22.5 67.0 76.4 35 55.9 58.0 46.3 37.0 42.6 41.6
-  + CoT 47.0 68.0 32.0 46.0 8.7 53.9 15.5 64.3 13.0 56.5 45.0 52.5 23.0 24.5 30.0 35.0 37.0 40.2 22.0 22.5 72.0 79.4 42 60.3 54.0 48.3 35.3 42.6 41.8
-Method BLEU-4 METEOR ROUGE-L CIDEr MATCH Overall Norm
-Human 39.9 40.4 56.3 397 98.1 274.4 100.0
-GPT4 8.23 31.27 28.84 38.45 90.41 162.59 59.25
-  + FS 28.64 38.99 55.69 298.64 90.11 243.29 88.66
-GPT3.5 13.38 30.12 35.91 125.41 78.76 170.70 62.21
-  + FS 27.24 33.77 51.18 282.75 76.54 217.01 79.08
-Baichuan2 8.33 25.86 30.07 82.63 70.63 143.15 52.17
-  + FS 15.79 30.23 40.96 169.14 71.01 174.91 63.74
-LLaMA2 4.66 21.43 20.80 17.72 61.62 110.28 40.19
-  + FS 15.15 27.49 37.55 138.13 64.94 158.93 57.92
-Vicuna1.5 7.73 26.35 29.15 69.16 71.91 142.06 51.77
-  + FS 6.85 18.66 25.99 92.96 46.19 106.99 38.99
-FLANT5 16.20 24.43 29.38 95.17 56.38 135.91 49.53
-  + FS 12.88 30.38 36.27 92.20 76.44 165.19 60.20
-Mistral 5.82 22.89 24.19 44.03 63.74 121.03 44.11
-  + FS 18.96 29.02 43.15 185.61 63.24 172.93 63.02
-ChatGLM3 6.56 21.11 21.96 41.48 53.02 106.80 38.92
-  + FS 10.53 24.17 33.44 124.50 56.94 137.53 50.12
-Baichuan2-7b 5.17 21.99 23.73 44.80 59.85 115.22 41.99
-  + FS 15.06 23.45 32.29 137.94 52.04 136.64 49.79
-LLaMA2-7b 2.77 13.46 14.69 14.34 34.83 67.18 24.48
-  + FS 6.90 15.82 21.77 52.99 33.81 83.60 30.47
-Vicuna1.5-7b 6.29 24.34 26.91 46.90 68.84 131.07 47.77
-  + FS 20.71 30.19 45.20 203.20 67.58 184.00 67.05
-${}^{\dagger}$ 17.86 29.75 44.28 198.83 66.35 178.12 64.91
-${}^{\dagger}$ 17.54 29.44 45.21 200.14 65.64 177.84 64.81
-${}^{\dagger}$ 15.30 27.54 41.80 171.59 62.40 164.20 59.84
-${}^{\dagger}$ 17.49 28.33 45.24 202.08 59.98 171.25 62.41
-${}^{\dagger}$ 14.54 27.39 41.72 168.89 59.42 159.96 58.30
-${}^{\dagger}$ 17.11 29.35 40.74 156.49 66.18 169.02 61.60
-Table 8: Full results of SituatedGen.
-Aligned models are under zero-shot setting by default.
-The top-3 results are bold.
-Methods with † are base models without alignment, under few-shot setting, thus incomparable with other methods.
-We consider human performance as 100 points and normalize models’ results accordingly.
-DurationQA, MCTACO
-Answer the following question, select all the possible correct options, and each question has at least one correct option.
-Context: {}
-Question: {}
-Options: {}
-Answer:
-TimeDial
+TimeXNLI Date Arith DurationQA McTACO TimeDial SitGen TimeQA MenatQA TempReason TRACIE Sym. Comm. Event Avg.
+s1 s2 s3 Acc EM F1 EM F1 EM F1 Norm E-EM E-F1 H-EM H-F1 S-EM S-F1 O-EM O-F1 C-EM C-F1 L2-EM L2-F1 L3-EM L3-F1 Acc
+Human 98.0 96.0 92.0 100.0 64.0 80.8 75.8 87.1 97.8 97.8 100.0 89.0 93.3 87.0 91.1 82.0 85.6 84.0 87.3 76.0 79.9 96.0 97.1 94.0 95.3 82.5 96.5 91.4 89.0 91.5
+GPT-4 78.6 76.0 50.7 98.0 35.0 59.2 61.2 80.0 72.0 91.1 59.3 48.9 60.6 40.4 46.5 44.4 57.0 49.0 57.0 22.0 23.1 91.0 95.3 94.0 95.0 64.8 75.8 72.4 62.4 68.3
++ CoT 80.0 76.0 60.0 92.0 35.0 58.1 67.0 82.6 65.0 89.3 - 50.0 61.3 33.0 41.2 43.4 54.6 53.0 59.6 20.0 22.6 93.0 97.0 93.0 94.5 58.0 77.0 76.7 61.1 68.5
++ FS 85.3 73.3 53.3 100.0 51.0 64.8 77.6 88.3 85.0 94.6 88.6 59.2 73.7 40.0 51.0 59.6 72.4 48.0 54.8 25.3 28.7 86.0 92.4 94.8 95.9 62.8 78.0 84.1 66.5 73.7
++ FS CoT 92.0 84.0 64.0 100.0 42.0 55.1 68.0 72.3 79.0 93.4 - 48.0 66.9 44.4 52.8 48.5 65.3 44.0 52.6 22.0 25.9 91.0 96.9 93.0 94.6 66.4 85.0 73.6 65.2 72.1
+GPT-3.5 45.4 67.6 31.2 97.0 19.2 50.5 34.1 68.6 39.2 69.1 62.3 60.5 70.8 29.5 35.4 36.5 40.9 37.5 43.9 21.0 22.9 73.6 81.2 61.8 73.8 57.4 60.3 62.6 53.3 57.4
++ CoT 33.6 64.8 33.6 71.0 12.4 23.2 28.1 45.1 34.6 67.0 - 52.5 64.4 29.0 35.1 35.8 39.7 38.5 42.9 24.0 26.3 32.0 57.6 54.2 68.1 52.0 50.8 45.1 48.3 48.3
++ FS 52.0 68.4 31.6 63.6 42.8 67.7 43.5 71.2 47.8 76.4 79.1 53.8 66.1 37.9 48.4 37.8 43.2 43.5 51.6 16.0 17.9 77.7 84.7 70.0 78.0 55.0 53.9 73.6 55.6 59.7
++ FS CoT 51.6 71.8 36.6 84.4 20.8 41.2 21.4 38.1 48.3 71.1 - 56.5 68.0 37.5 47.0 38.1 42.5 37.5 41.7 33.0 37.8 86.2 89.9 68.0 76.6 50.2 61.1 50.1 56.7 56.6
+LLaMA270b 44.0 47.0 32.0 78.5 12.7 59.2 23.0 68.9 10.0 57.0 25.0 28.0 40.8 31.0 40.6 8.0 18.9 11.0 16.6 9.0 12.0 50.0 63.5 39.0 54.5 48.0 50.4 52.5 36.8 44.1
++ CoT 30.0 66.0 28.0 53.5 8.0 57.3 21.0 67.1 9.0 58.6 17.0 31.4 13.0 19.5 5.0 12.2 8.0 12.7 18.0 20.8 12.0 37.5 20.0 40.5 51.0 44.4 61.0 28.2 39.1
++ FS 49.0 42.0 38.0 62.0 1.3 61.2 13.0 66.5 6.0 56.6 62.8 41.0 51.1 16.0 20.0 8.0 16.4 17.0 19.9 18.0 18.7 34.0 52.2 31.0 41.1 51.0 47.8 61.8 33.8 44.3
++ FS CoT 54.0 63.0 40.0 69.5 8.0 55.2 21.5 62.1 6.0 56.4 36.6 50.9 34.0 42.4 28.0 38.6 19.0 29.3 18.0 21.9 77.0 83.1 65.0 74.7 57.0 56.6 57.9 49.7 53.2
+LLaMA213b 30.0 49.0 34.0 22.5 4.0 38.5 8.5 40.6 10.0 35.4 57.9 46.0 61.9 21.0 30.5 28.0 46.1 23.0 36.1 18.0 26.9 43.0 53.1 55.0 69.4 49.0 33.9 43.1 46.6 42.6
++ CoT 36.0 50.0 38.0 6.0 7.3 39.2 14.0 51.7 10.0 36.9 - 45.0 58.7 30.0 38.9 20.0 40.9 18.0 32.5 21.0 33.6 43.0 58.0 56.0 68.4 47.0 32.5 42.6 47.3 42.4
++ FS 43.0 57.0 60.0 20.5 9.0 46.8 8.0 66.6 15.0 62.3 40.2 24.0 34.2 17.0 18.4 11.0 25.9 5.0 14.6 22.0 33.3 54.0 68.1 50.0 64.8 47.0 45.1 54.0 38.3 43.9
++ FS CoT 37.0 55.0 50.0 33.0 12.0 49.5 11.0 45.6 8.0 44.5 - 35.0 46.0 21.0 25.4 34.0 46.7 23.0 36.5 7.0 16.5 72.0 80.8 54.0 66.2 50.0 43.8 46.5 46.0 45.5
+LLaMA27b 39.0 53.0 30.0 13.0 2.7 39.3 4.0 41.0 1.0 6.3 24.5 37.0 49.0 14.0 29.0 7.0 26.8 8.0 21.1 9.0 16.0 48.0 63.9 32.0 47.9 49.0 33.8 27.8 37.8 34.3
++ CoT 44.0 50.0 33.0 5.0 2.7 35.0 4.5 40.0 1.0 1.7 - 27.0 49.9 17.0 31.6 11.0 31.4 10.0 24.5 7.0 17.8 44.0 56.9 32.0 48.1 46.0 33.0 25.6 38.3 34.3
++ FS 44.0 60.0 34.0 11.0 4.0 62.8 8.0 64.7 8.0 40.0 30.5 36.0 50.8 20.0 29.4 5.0 22.3 6.0 18.0 6.0 17.9 12.0 36.3 23.0 44.3 53.0 37.3 49.5 34.0 38.7
++ FS CoT 38.0 51.0 36.0 14.5 11.0 42.8 25.0 65.6 13.0 53.4 - 36.0 53.5 21.0 34.1 1.0 13.6 3.0 11.2 5.0 14.0 22.0 46.7 21.0 42.3 51.0 34.9 53.9 33.3 37.8
+Baichuan213b 41.0 61.0 37.0 12.5 4.0 52.0 18.5 63.4 15.0 57.7 52.2 45.0 55.4 29.0 34.6 31.0 48.8 34.0 44.3 30.0 39.5 40.0 57.4 45.0 61.4 49.0 37.9 56.3 48.8 48.0
++ CoT 40.0 57.0 31.0 10.0 3.3 44.6 20.0 61.9 13.0 58.1 - 36.0 41.5 36.0 40.9 39.0 52.0 27.0 38.5 29.0 43.2 46.0 62.8 46.0 64.3 55.0 34.5 54.9 49.8 46.7
++ FS 43.0 59.0 40.0 42.5 24.7 62.1 27.5 70.2 18.0 58.9 63.7 47.0 60.7 35.0 45.7 37.0 51.9 31.0 41.5 19.0 31.8 73.0 81.1 48.0 59.4 48.0 46.1 63.7 52.5 53.7
++ FS CoT 45.0 54.0 48.0 47.0 10.7 44.4 27.0 68.8 15.0 55.0 - 43.0 57.8 27.0 36.7 38.0 49.8 34.0 40.7 33.0 43.0 72.8 80.4 43.0 60.2 44.0 48.5 56.1 51.6 51.7
+Baichuan27b 35.0 50.0 37.0 4.5 4.0 47.9 10.5 55.3 15.0 54.3 42.0 26.0 41.5 20.0 34.7 20.0 35.2 19.0 31.2 6.0 20.4 22.0 43.4 29.0 47.7 55.0 31.6 49.9 38.6 39.7
++ CoT 38.0 43.0 32.0 1.0 5.3 37.9 13.0 58.0 15.0 44.2 - 41.0 53.5 28.0 38.8 29.0 39.9 23.0 33.2 18.0 29.3 21.0 41.2 29.0 47.2 54.0 28.5 46.7 42.1 39.4
++ FS 40.0 50.0 36.0 20.0 28.7 59.4 26.5 66.9 17.0 53.0 49.8 45.0 60.7 30.0 42.1 27.0 37.8 23.0 35.7 10.0 20.4 40.0 57.4 37.0 53.0 51.0 36.5 57.3 44.8 45.8
++ FS CoT 41.0 50.0 36.0 23.5 13.0 45.7 17.5 58.1 7.0 39.2 - 36.0 51.2 29.0 43.0 42.0 52.5 25.0 39.3 20.0 31.0 57.0 70.1 39.0 60.2 49.0 37.6 47.7 49.5 46.0
+Vicuna1.513b 35.0 50.0 36.0 15.0 8.0 39.2 21.5 59.1 7.0 34.2 51.8 43.0 60.4 29.0 37.0 38.0 46.8 22.0 37.4 17.0 23.2 14.0 42.1 13.0 43.6 46.0 34.0 46.1 42.1 41.1
++ CoT 42.0 51.0 37.0 3.0 1.3 29.8 11.5 50.0 7.0 33.7 - 44.0 56.9 31.0 36.4 16.0 38.2 25.0 37.7 13.0 20.4 31.0 49.0 29.0 49.1 51.0 33.3 37.8 42.3 39.0
++ FS 48.0 57.0 38.0 30.5 7.3 33.6 27.5 57.0 13.0 40.3 39.0 45.0 58.3 23.0 25.9 38.0 42.6 26.0 41.4 18.0 20.1 51.0 61.8 28.0 42.6 56.0 43.4 42.5 43.6 43.3
++ FS CoT 38.0 59.0 39.0 39.5 10.7 37.4 14.0 45.8 12.0 41.6 - 47.0 59.5 27.0 30.7 39.0 48.1 31.0 35.9 26.0 31.2 71.0 77.5 53.0 65.5 52.0 43.9 41.6 50.1 46.7
+Vicuna1.57b 37.0 58.0 43.0 5.0 1.3 40.4 9.5 52.5 6.0 32.0 47.8 35.0 47.1 11.0 18.5 20.0 35.7 15.0 25.7 12.0 17.3 14.0 33.0 14.0 46.8 54.0 35.8 43.2 34.8 37.1
++ CoT 36.0 50.0 36.0 1.5 1.3 39.4 8.5 49.2 9.0 36.2 - 30.0 40.9 14.0 24.6 16.0 26.2 14.0 28.5 12.0 25.0 9.0 27.7 7.0 40.3 54.0 30.9 41.6 33.4 34.4
++ FS 43.0 57.0 37.0 8.5 3.3 44.6 5.5 42.1 7.0 36.8 67.1 24.0 31.9 12.0 14.9 16.0 21.8 20.0 27.5 17.0 22.2 13.0 34.3 6.0 32.2 54.0 36.4 47.7 29.9 35.9
++ FS CoT 35.0 54.0 35.0 8.0 2.7 37.2 10.0 47.5 10.0 41.3 - 31.0 39.9 13.0 16.6 15.0 26.7 15.0 23.8 16.0 23.1 55.0 66.3 32.0 48.1 43.0 33.0 42.0 35.9 36.4
+FLANT511b 53.0 63.0 43.0 0.0 4.0 52.0 14.0 65.0 13.0 47.7 49.5 56.0 61.7 24.0 26.8 31.0 33.6 48.0 52.2 20.0 21.8 84.0 87.9 78.0 83.9 64.0 39.8 53.6 54.0 50.3
++ CoT 56.0 66.0 45.0 0.0 4.7 49.7 14.5 63.4 13.0 42.7 - 57.6 64.4 23.9 28.2 39.0 41.6 46.0 50.2 28.0 30.6 73.0 79.5 57.0 68.9 55.0 41.8 51.9 52.3 49.4
++ FS 53.0 65.0 43.0 3.5 4.0 50.2 15.5 65.8 9.0 35.0 60.2 55.0 64.3 23.0 25.0 31.0 33.6 47.0 50.6 21.0 22.5 82.0 87.0 78.0 84.5 65.0 41.1 52.8 54.1 50.5
++ FS CoT 54.0 68.0 46.0 3.5 4.0 50.7 13.0 64.0 11.0 43.7 - 54.0 59.4 19.0 21.2 34.0 36.6 45.0 49.2 20.0 21.7 89.0 93.8 72.0 79.7 66.0 42.9 52.8 53.5 50.5
+Mistral7b 47.0 50.0 43.0 26.5 11.3 49.8 15.0 58.8 6.0 23.2 58.3 7.0 28.2 5.0 21.4 4.0 24.3 7.0 22.3 4.0 21.7 2.0 39.6 1.0 31.6 51.0 41.6 47.5 30.0 37.3
++ CoT 38.0 56.0 35.0 16.5 13.3 36.6 14.5 49.3 8.0 19.3 - 13.0 31.3 11.0 22.4 8.0 21.1 14.0 24.9 12.0 25.6 5.0 34.0 4.0 31.2 61.0 36.4 35.1 31.4 33.5
++ FS 51.0 57.0 35.0 18.0 12.0 55.8 25.5 71.0 13.0 52.9 63.0 24.0 43.5 12.0 23.9 4.0 21.3 7.0 21.2 7.0 23.0 23.0 48.9 23.0 44.9 57.0 40.3 60.7 35.5 43.0
++ FS CoT 29.0 62.0 32.0 28.5 9.3 46.6 14.5 49.2 8.0 34.4 - 13.0 33.7 11.0 25.3 9.0 26.4 9.0 24.3 9.0 20.3 68.0 78.6 42.0 57.4 50.0 37.9 43.4 39.5 39.8
+ChatGLM36b 38.0 50.0 34.0 2.0 3.0 34.1 7.0 43.6 14.0 56.7 38.9 20.0 41.2 14.0 31.7 25.0 33.8 17.0 26.0 24.0 32.2 42.0 57.0 30.0 54.0 50.0 31.0 43.3 40.7 39.0
++ CoT 27.0 49.0 37.0 0.0 1.0 24.8 3.0 37.1 10.0 44.8 - 24.0 41.7 10.0 25.4 27.0 34.6 22.0 28.1 35.0 41.2 28.0 44.5 27.0 52.0 48.0 28.3 35.6 39.4 35.7
++ FS 37.0 52.0 30.0 0.0 10.0 53.0 9.0 52.9 19.0 54.7 50.1 11.0 19.6 13.0 17.0 25.0 35.0 24.0 30.4 20.0 25.6 27.0 46.5 33.0 53.3 54.0 29.8 52.7 35.2 38.2
++ FS CoT 32.0 66.0 31.0 0.0 4.0 34.8 8.0 43.6 11.0 44.0 - 25.0 43.3 19.0 27.5 23.0 30.2 19.0 25.7 36.0 43.0 50.0 62.8 34.0 56.0 48.0 32.3 40.8 42.1 39.2
+LLaMA2-Base70b 55.0 61.0 37.0 82.0 40.0 67.4 59.0 85.3 63.0 82.7 74.9 57.0 66.7 36.0 48.3 52.0 61.4 35.0 42.5 25.0 33.8 78.0 85.2 80.0 85.4 61.0 58.8 77.6 60.5 64.4
++ CoT 52.0 73.0 39.0 79.5 32.6 62.3 47.0 79.1 25.0 61.1 55.0 64.3 34.0 43.0 50.0 57.7 37.0 45.2 44.0 53.1 82.0 87.5 76.0 81.6 67.0 60.9 67.5 62.4 63.0
+LLaMA2-Base13b 50.0 54.0 30.0 29.5 19.3 53.3 26.5 66.0 20.0 55.6 64.8 48.0 59.3 34.0 48.6 41.0 49.6 38.0 43.4 34.0 37.5 68.0 78.7 49.0 62.7 58.0 40.9 59.9 54.7 52.6
++ CoT 40.0 61.0 37.0 52.0 25.3 59.3 26.0 68.8 11.0 40.8 - 46.0 59.4 37.0 49.1 49.0 58.4 34.0 43.8 38.0 44.1 70.0 78.0 55.0 68.2 58.0 47.5 56.3 57.4 54.5
+LLaMA2-Base7b 26.0 50.0 30.0 20.0 19.3 54.5 20.0 59.6 15.0 45.2 62.4 44.0 54.4 30.0 45.3 42.0 49.8 34.0 41.9 30.0 35.8 50.0 64.0 36.0 53.3 49.0 31.5 55.4 49.2 46.3
++ CoT 37.0 52.0 36.0 25.5 21.3 56.9 26.5 67.0 16.0 41.9 - 32.0 45.6 27.0 36.1 41.0 50.9 30.0 38.0 51.0 57.3 45.0 59.7 37.0 57.7 50.0 37.6 55.3 49.4 47.4
+Baichuan2-Base13b 38.0 48.0 33.0 42.5 20.7 54.8 42.5 73.0 11.0 45.7 64.9 50.0 59.4 40.0 54.2 42.0 52.7 31.0 38.0 13.0 21.4 68.0 77.3 50.0 63.5 54.0 40.4 59.6 52.6 51.3
++ CoT 50.0 56.0 34.0 47.0 29.3 62.0 22.5 69.3 12.0 43.8 - 46.0 58.2 39.0 49.6 39.0 49.8 34.0 40.1 37.0 45.6 73.0 81.3 46.0 65.6 60.0 46.8 58.4 56.3 54.2
+Baichuan2-Base7b 27.0 66.0 41.0 32.5 28.0 59.8 34.5 69.4 5.0 34.3 59.8 40.0 53.8 35.0 50.2 41.0 49.6 33.0 38.5 18.0 22.9 49.0 65.9 34.0 51.0 55.0 41.6 55.8 48.4 48.5
++ CoT 30.0 56.0 34.0 34.0 23.3 57.0 33.0 69.5 12.0 44.5 - 41.0 51.2 31.0 40.7 38.0 46.4 26.0 32.6 41.7 46.3 46.0 61.5 43.8 64.1 53.0 38.5 57.0 49.5 48.1
+Mistral-Base7b 48.0 53.0 38.0 41.0 34.0 61.8 42.5 76.2 35.0 61.8 58.3 43.0 55.9 30.0 45.3 37.0 49.4 38.0 47.8 37.0 45.5 68.0 76.7 64.0 74.8 53.0 45.0 64.5 56.1 55.4
++ CoT 57.0 63.0 35.0 54.0 30.0 61.8 42.0 45.7 29.0 57.3 - 51.0 60.4 30.0 46.2 48.0 57.2 37.0 47.9 24.0 33.2 60.0 65.9 58.0 67.9 57.0 52.3 54.9 54.5 54.0
+ChatGLM3-Base6b 48.0 70.0 32.0 35.0 3.3 51.8 13.5 62.6 11.0 55.0 61.6 50.0 57.2 24.0 26.3 30.0 35.4 38.0 41.5 22.0 22.5 67.0 76.4 35.0 55.9 58.0 46.3 57.8 46.7 49.3
++ CoT 47.0 68.0 32.0 46.0 8.7 53.9 15.5 64.3 13.0 56.5 45.0 52.5 23.0 24.5 30.0 35.0 37.0 40.2 22.0 22.5 72.0 79.4 42.0 60.3 54.0 48.3 58.2 46.1 49.1
+Answer the following question, select all the possible correct options, and each question has at least one correct option. Context: {} Question: {} Options: {} Answer: DurationQA, MCTACO
 There is a two-person dialogue with several options.
 Choose all appropriate options to substitute the <mask> in the dialogue, and each question has at least one correct option.
-Dialogue: {}
-Options: {}
-Answer:
-TRACIE
+Dialogue: {} Options: {} Answer: TimeDial
 Read the following story and hypothesis, determine whether the hypothesis can be inferred from the story.
 You need to understand the implicit temporal relationships between events to make judgments.
-Story: {}
-Hypothesis: {}
-Options: A. Entailment B. Contradiction
-Answer:
-SituatedGen
-Generate a pair of contrastive sentences with the given set of keywords.
-Keywords: {}
-Date Arithmetic
-Question: {}? Answer:
-TimeQA
+Story: {} Hypothesis: {} Options: A.
+Entailment B.
+Contradiction Answer: TRACIE
+Generate a pair of contrastive sentences with the given set of keywords. Keywords: {} SituatedGen
+Question: {}? Answer: Date Arithmetic
 I will give you a question with context.
 You need to answer my question based on the context.
-If you can infer the answer from the context, then output your answer. Otherwise, if there is no answer, output [unanswerable].
-Context: {}
-Question: {}
-Answer:
-TempReason
-I will give you a question with context.
-You need to answer my question based on the context.
-Context: {}
-Question: {}
-Answer:
-MenatQA
-Get answers for the question based on the contxt, where answers derived from substrings in the context or categorized as [unanswerable].
-Context: {}
-Question: {}
-Answer:
-TimeX-NLI
+If you can infer the answer from the context, then output your answer.
+Otherwise, if there is no answer, output [unanswerable].
+Context: {} Question: {} Answer: TimeQA
+I will give you a question with context. You need to answer my question based on the context. Context: {} Question: {} Answer: TempReason
+Get answers for the question based on the contxt, where answers derived from substrings in the context or categorized as [unanswerable]. Context: {} Question: {} Answer: MenatQA
 Read the following statements about time and determine if the hypothesis can be inferred from the premise.
-Premise: {}
-Hypothesis: {}
-Options: A. Entailment B. Contradiction C. Neutral
-Answer:
+Premise: {} Hypothesis: {} Options: A.
+Entailment B.
+Contradiction C.
+Neutral Answer: TimeX-NLI
 Figure 9: Zeroshot instructions and input formats.
-CoT Demonstration of TimeX-NLI (3-shot, order)
 Answer the following question, select all the possible correct options, and each question has at least one correct option.
@@
 Hypothesis: Before Friday, they got married.
-Options: A. Entailment B. Contradiction C. Neutral
-Answer: Wednesday is before Friday. As a result, we can infer that if something happens on Wednesday, it definitely happens before Friday. Therefore, the answer is A. Entailment.
+Options: A.
+Entailment B.
+Contradiction C.
+Neutral Answer: Wednesday is before Friday.
+As a result, we can infer that if something happens on Wednesday, it definitely happens before Friday.
+Therefore, the answer is A.
+Entailment.
 Premise: We went to Disneyland on Monday.
 Hypothesis: We went to Disneyland after Wednesday.
-Options: A. Entailment B. Contradiction C. Neutral
-Answer: Monday is before Wednesday. As a result, We can infer that if something happens on Monday, it definitely can not happen after Wednesday. Therefore, the answer is B. Contradiction.
+Options: A.
+Entailment B.
+Contradiction C.
+Neutral Answer: Monday is before Wednesday.
+As a result, We can infer that if something happens on Monday, it definitely can not happen after Wednesday.
+Therefore, the answer is B.
+Contradiction.
 Premise: The failing company issued major layoffs after Tuesday.
 Hypothesis: The failing company issued major layoffs after Thursday.
-Options: A. Entailment B. Contradiction C. Neutral
-Answer: Tuesday is before Thursday. If something happened after Tuesday, we cannot be certain whether it occurred after Thursday. Therefore, the answer is C. Neutral.
+Options: A.
+Entailment B.
+Contradiction C.
+Neutral Answer: Tuesday is before Thursday.
+If something happened after Tuesday, we cannot be certain whether it occurred after Thursday.
+Therefore, the answer is C.
+Neutral.
+CoT Demonstration of TimeX-NLI (3-shot, order)
 Figure 10: Chain-of-Thought demonstrations of TimeX-NLI (s1-order).
+Question: What is the time 4 year and 1 month after Apr, 2000? Answer: First, 4 years after 2000 is 2004. Next, 1 month after April is May. Therefore, 4 year and 1 month after Apr, 2000 is May, 2004.
+Question: What is the time 3 year and 4 month before Jun, 1840?
+Answer: First, subtracting 3 years from 1840 gives 1837.
+Next, subtracting 4 months from June gives February.
+Therefore, 3 year and 4 month before Jun, 1840 is Feb, 1837.
+Question: What is the time 7 year and 11 month after Feb, 1819?
+Answer: First, 7 years after 1819 is 1826.
+Next, 11 months after February is January of the next year.
+Therefore, 7 years and 11 months after Feb, 1819 is Jan, 1827.
+Question: What is the time 6 year and 9 month before Jan, 1234?
+Answer: First, subtracting 6 years from 1234 gives 1228.
+Next, subtracting 9 months from January gives April of the previous year.
+Therefore, 6 year and 9 month before Jan, 1234 is Apr, 1227.
 CoT Demonstration of Date Arithmetic (4-shot)
-Question: What is the time 4 year and 1 month after Apr, 2000?
-Answer: First, 4 years after 2000 is 2004. Next, 1 month after April is May. Therefore, 4 year and 1 month after Apr, 2000 is May, 2004.
-Question: What is the time 3 year and 4 month before Jun, 1840?
-Answer: First, subtracting 3 years from 1840 gives 1837. Next, subtracting 4 months from June gives February. Therefore, 3 year and 4 month before Jun, 1840 is Feb, 1837.
-Question: What is the time 7 year and 11 month after Feb, 1819?
-Answer: First, 7 years after 1819 is 1826. Next, 11 months after February is January of the next year. Therefore, 7 years and 11 months after Feb, 1819 is Jan, 1827.
-Question: What is the time 6 year and 9 month before Jan, 1234?
-Answer: First, subtracting 6 years from 1234 gives 1228. Next, subtracting 9 months from January gives April of the previous year. Therefore, 6 year and 9 month before Jan, 1234 is Apr, 1227.
 Figure 11: Chain-of-Thought demonstrations of Date Arithmetic.
-CoT Demonstration of TRACIE (4-shot)
-Read the following story and hypothesis, determine whether the hypothesis can be inferred from the story.
-You need to understand the implicit temporal relationships between events to make judgments
+Read the following story and hypothesis, determine whether the hypothesis can be inferred from the story. You need to understand the implicit temporal relationships between events to make judgments
 Story: Joe was a police officer.
@@
 Joe arrested the criminal and was promoted.
-Hypothesis: Joe put on his police uniform. starts after Joe arrest the criminal
-Options: A. Entailment B. Contradiction
-Answer: From the story we know Joe was patrolling.
+Hypothesis: Joe put on his police uniform.
+starts after Joe arrest the criminal Options: A.
+Entailment B.
+Contradiction Answer: From the story we know Joe was patrolling.
 In the work state, Joe has already put on the police uniform.
@@
 Contradiction.
+CoT Demonstration of TRACIE (4-shot)
 Figure 12: Chain-of-Thought demonstrations of TRACIE.
-CoT Demonstration of DurationQA (4-shot)
 Answer the following question, select all the possible correct options, and each question has at least one correct option.
-Context: actually i have an project on it so please give me as much as you have information about migratory birds in punjab
-Question: How long did it take for them to have information about migratory birds in punjab?
-Options: A. several months B. 12 weeks C. a few minutes D. almost instantly
-Answer: This is a conversation scenario.
+Context: actually i have an project on it so please give me as much as you have information about migratory birds in punjab Question: How long did it take for them to have information about migratory birds in punjab?
+Options: A.
+several months B.
+12 weeks C.
+a few minutes D.
+almost instantly Answer: This is a conversation scenario.
 In the conversation, providing relevant information about migratory birds in punjab to him is in real-time and takes very little time.
@@
 Question: How long did it take for her to lay eggs?
-Options: A. 1 week B. 22 hours C. 2 years D. 4 years
-Answer: According to commonsense knowledge, the time it takes for birds to lay eggs typically varies from one day to several days. Therefore, the answer is A. 1 week, B. 22 hours.
+Options: A.
+1 week B.
+22 hours C.
+2 years D.
+4 years Answer: According to commonsense knowledge, the time it takes for birds to lay eggs typically varies from one day to several days.
+Therefore, the answer is A.
+1 week, B.
+22 hours.
+CoT Demonstration of DurationQA (4-shot)
 Figure 13: Chain-of-Thought demonstrations of DurationQA.
-CoT Demonstration of MCTACO (4-shot)
 Answer the following question, select all the possible correct options, and each question has at least one correct option.
@@
 Question: How often does she order vegetables and flowers?
-Options: A. once a second B. three days a week C. every 10 centuries D. once a week
-Answer: According to commonsense knowledge, ordering vegetables and flowers typically happens on a regular basis, usually every few days. Therefore, the answer is B. three days a week, D. once a week.
+Options: A.
+once a second B.
+three days a week C.
+every 10 centuries D.
+once a week Answer: According to commonsense knowledge, ordering vegetables and flowers typically happens on a regular basis, usually every few days.
+Therefore, the answer is B.
+three days a week, D.
+once a week.
 Context: Wallace, 38, called Gastonia home from the age of 8 until she graduated from Hunter Huss High School in 1983.
 Question: When did Wallace wake up for high school?
-Options: A. at 6 am B. at 1 am C. 7:00 AM D. at 6 pm
-Answer: According to commonsense knowledge, waking up for high school typically happens in the morning, usually between 6 AM and 8 AM. Therefore, the answer is A. at 6 am, C. 7:00 AM.
+Options: A.
+at 6 am B.
+at 1 am C.
+7:00 AM D.
+at 6 pm Answer: According to commonsense knowledge, waking up for high school typically happens in the morning, usually between 6 AM and 8 AM.
+Therefore, the answer is A.
+at 6 am, C.
+7:00 AM.
+CoT Demonstration of MCTACO (4-shot)
 Figure 14: Chain-of-Thought demonstrations of MCTACO.
-CoT Demonstration of TimeDial (4-shot)
-There is a two-person dialogue with several options.
-Choose all appropriate options to substitute the <mask> in the dialogue, and each question has at least one correct option.
-Dialogue:
-A:What schools have you attended ?
+There is a two-person dialogue with several options. Choose all appropriate options to substitute the <mask> in the dialogue, and each question has at least one correct option.
+Dialogue: A:What schools have you attended ?
 B: I finished Young Primary School in 1998 , and entered Xi ’ an Middle School that same September .
@@
 B: I think I have a good understanding of fundamentals in the areas your company deals with , and I can go on from here to build up the specific skills and knowledge I need to do my job well .
-A: Your graduation thesis was on Medical Application of Laser , right ? What were your conclusions ?
+A: Your graduation thesis was on Medical Application of Laser , right ?
+What were your conclusions ?
 B: Yes .
@@
 I think the most important thing , I learnt though , was the importance of keeping good records .
-Options: A. 1998 B. July of 2004 C. March of 2003 D. twenty years ago
-Answer: Based on the dialogue, B entered middle school in Sep 1998.
+Options: A.
+1998 B.
+July of 2004 C.
+March of 2003 D.
+twenty years ago Answer: Based on the dialogue, B entered middle school in Sep 1998.
 According to commonsense knowledge, it usually takes around 6 years from entering middle school to graduating from high school (and entering university).
@@
 March of 2003.
+CoT Demonstration of TimeDial (4-shot)
 Figure 15: Chain-of-Thought demonstrations of TimeDial.
-CoT Demonstration of TimeQA, MenatQA (2-shot, implicit)
 I will give you a question with context.
 You need to answer my question based on the context.
-If you can infer the answer from the context, then output your answer. Otherwise, if there is no answer, output [unanswerable]
+If you can infer the answer from the context, then output your answer.
+Otherwise, if there is no answer, output [unanswerable]
 Context: Theo-Ben Gurirab Theo-Ben Gurirab ( 23 January 1938 2̆013 14 July 2018 ) was a Namibian politician who served in various senior government positions .
@@
 Therefore, the answer is Speaker of the National Assembly of Namibia.
+CoT Demonstration of TimeQA, MenatQA (2-shot, implicit)
 Figure 16: Chain-of-Thought demonstrations of TimeQA, MenatQA, implicit reasoning.
-CoT Demonstration of TempReason (4-shot, event-time)
-I will give you a question with context.
-You need to answer my question based on the context.
-Context (facts): Gian Piero Gasperini is the head coach of Atalanta B.C. from Jun, 2016 to Dec, 2022.
-Edoardo Reja is the head coach of Atalanta B.C. from Mar, 2015 to Jun, 2016.
-Stefano Colantuono is the head coach of Atalanta B.C. from Jun, 2010 to Mar, 2015.
-Bortolo Mutti is the head coach of Atalanta B.C. from Jan, 2010 to Jun, 2010.
-Emiliano Mondonico is the head coach of Atalanta B.C. from Jul, 1987 to Jun, 1990.
-Marcello Lippi is the head coach of Atalanta B.C. from Jul, 1992 to Jun, 1993.
-Angelo Gregucci is the head coach of Atalanta B.C. from Jul, 2009 to Sep, 2009.
-Luigi Delneri is the head coach of Atalanta B.C. from Jul, 2007 to Jun, 2009.
-Ottavio Bianchi is the head coach of Atalanta B.C. from Jul, 1981 to Jun, 1983.
-Antonio Conte is the head coach of Atalanta B.C. from Sep, 2009 to Jan, 2010.
-Nedo Sonetti is the head coach of Atalanta B.C. from Jul, 1983 to Jun, 1987.
-Valter Bonacina is the head coach of Atalanta B.C. from Jan, 2010 to Jan, 2010.
-Question: Who was the head coach of the team Atalanta B.C. in Feb, 2016?
+I will give you a question with context. You need to answer my question based on the context.
+Context (facts): Gian Piero Gasperini is the head coach of Atalanta B.C.
+from Jun, 2016 to Dec, 2022.
+Edoardo Reja is the head coach of Atalanta B.C.
+from Mar, 2015 to Jun, 2016.
+Stefano Colantuono is the head coach of Atalanta B.C.
+from Jun, 2010 to Mar, 2015.
+Bortolo Mutti is the head coach of Atalanta B.C.
+from Jan, 2010 to Jun, 2010.
+Emiliano Mondonico is the head coach of Atalanta B.C.
+from Jul, 1987 to Jun, 1990.
+Marcello Lippi is the head coach of Atalanta B.C.
+from Jul, 1992 to Jun, 1993.
+Angelo Gregucci is the head coach of Atalanta B.C.
+from Jul, 2009 to Sep, 2009.
+Luigi Delneri is the head coach of Atalanta B.C.
+from Jul, 2007 to Jun, 2009.
+Ottavio Bianchi is the head coach of Atalanta B.C.
+from Jul, 1981 to Jun, 1983.
+Antonio Conte is the head coach of Atalanta B.C.
+from Sep, 2009 to Jan, 2010.
+Nedo Sonetti is the head coach of Atalanta B.C.
+from Jul, 1983 to Jun, 1987.
+Valter Bonacina is the head coach of Atalanta B.C.
+from Jan, 2010 to Jan, 2010.
+Question: Who was the head coach of the team Atalanta B.C.
+in Feb, 2016?
 Answer: According to the context, Edoardo Reja was the head coach of Atalanta B.C.
@@
 Therefore, the answer is Edoardo Reja.
+CoT Demonstration of TempReason (4-shot, event-time)
 Figure 17: Chain-of-Thought demonstrations of TempReason, event-time reasoning.
-CoT Demonstration of TempReason (4-shot, event-event)
-I will give you a question with context.
-You need to answer my question based on the context.
+I will give you a question with context. You need to answer my question based on the context.
 Context (facts): Nicholas Macpherson holds the position of Member of the House of Lords from Oct, 2016 to Dec, 2022.
@@
 Therefore, the answer is Permanent Secretary to the Treasury."
+CoT Demonstration of TempReason (4-shot, event-event)
 Figure 18: Chain-of-Thought demonstrations of TempReason, event-event reasoning.